### v2

# Accurate BGV Parameters Selection: Accounting for Secret and Public Key Dependencies in Average-Case Analysis

Abstract
The Brakerski-Gentry-Vaikuntanathan (BGV) scheme is one of the most significant fully homomorphic encryption (FHE) schemes.
It belongs to a class of FHE schemes whose security is based on the presumed intractability of the Learning with Errors (LWE) problem and its ring variant (RLWE). Such schemes deal with a quantity, called noise, which increases each time a homomorphic operation is performed. Specifically, in order for the scheme to work properly, it is essential that the noise remains below a certain threshold throughout the process. For BGV, this threshold strictly depends on the ciphertext modulus, which is one of the initial parameters whose selection heavily affects both the efficiency and security of the scheme.
In this paper, we provide a new method to estimate noise growth, closely aligning with experimental results and forming the basis for parameter selection that ensures correctness and improves efficiency.
1 Introduction
The first Fully Homomorphic Encryption (FHE) scheme was introduced in 2009 by Gentry [18]. Since then, several FHE constructions have been proposed, such as BGV [4], BFV [3, 17], FHEW [16], TFHE [9, 10], and CKKS [8, 7].
The homomorphic encryption schemes currently in use base their security on the presumed intractability of the Learning with Errors (LWE) problem, [26], and its ring variant (RLWE) [23]. Informally, the decisional version of RLWE consists of distinguishing polynomial equations $(a,b=s\cdot a+e)\in\mathcal{R}_{q}\times\mathcal{R}_{q}$, perturbed by small noise $e$ (also called error), from uniform random tuples from $\mathcal{R}_{q}\times\mathcal{R}_{q}$, where $\mathcal{R}_{q}=\mathbb{Z}_{q}[x]/\langle x^{n}+1\rangle$ and $q$ is a positive integer.
Schemes based on the (R)LWE problem face a critical challenge related to the growth of noise during homomorphic operations, which must be carefully controlled to ensure the correct functioning of the encryption scheme. Specifically, the noise must be kept below a certain threshold, which, in the case of BGV, is directly related to the ciphertext modulus parameter $q$.
As homomorphic operations are performed, the noise increases, and therefore, to maintain the integrity of the scheme, the parameter $q$ must be chosen sufficiently large. However, although increasing $q$ allows for a greater number of operations, it simultaneously compromises both the security and efficiency of the scheme. Therefore, selecting an appropriate value for $q$ and, in general, determining an optimal set of parameters, is critical. This process requires a balance between security and efficiency while ensuring the correctness of the scheme.
One of the key factors in achieving this balance and determining suitable parameters is providing accurate estimates of the error and its growth during the homomorphic operations in the circuit.
This issue is central to research in FHE, and over the years, various approaches have been proposed to address it. As for example, employing the Euclidean norm [4], the infinity norm [17, 22], and the canonical norm, also called worst-case analysis [11, 13, 19, 21, 24]. The prevailing trend in the current literature adopts the average-case analysis, which involves treating the noise coefficients as random variables distributed according to a Gaussian distribution and studying their expected value and variance.
Interest in this method, initially applied in the TFHE scheme [9], and subsequently in the CKKS [12], BGV [14, 25] and BFV [2] schemes, grew due to a recognized discrepancy between the estimates based on worst-case technique and experimental data, as highlighted in [13]. The introduction of the average-case approach, as seen in [2, 14], offers a potential resolution to these disparities, indeed, with this method, it is possible to compute a tight probabilistic upper bound.
However, the heuristics used for the BGV [25] and CKKS [12] schemes often underestimate the noise growth due to the assumption of the noises independence, leading to imprecise bounds, as also pointed out in [1, 12, 25]. Such underestimates lead to two potential issues: first, the ciphertext is not correctly decrypted with non-negligible probability due to excessive noise and, second, the scheme is exposed to security vulnerabilities, as shown in recent papers [5, 6].
In light of this, it becomes evident that accounting for the dependencies between the error coefficients is crucial in order to derive increasingly tighter and correct bounds. This, in turn, enables the definition of more accurate operational parameters, making the scheme both more secure and efficient, which is essential for the widespread adoption of FHE.
In this paper, we propose the first average-case noise analysis for BGV that does not provide underestimates, taking in account the dependencies introduced by the common secret and public key. We extend the approach of BFV [2], where the authors consider the fact that the errors are not independent among each other, and introduce a function $F$ to “correct” the product of the variances for homomorphic multiplication. We notice the additional dependencies given by the public key and employ a second correcting function to address its contribution. The results obtained in this study suggest that this approach leads to significant improvements in noise analysis.
A related average-case analysis for the BGV scheme is presented in [14], where the authors develop a noise estimation method tailored to the specific implementation of BGV in HElib [20]. In contrast, our work proposes a general analysis that does not depend on the specific library and instead focuses on capturing the structural dependencies among the errors. We show that considering these dependencies is essential to derive correct, accurate and tighter bounds, independently of specific implementations.
In the BGV scheme, each ciphertext is associated with a critical quantity $\nu$ which is a polynomial in $\mathcal{R}$. The critical quantity of a ciphertext $\mathbf{c}$ defines whether $\mathbf{c}$ can be correctly decrypted. Specifically, if the size of $\nu$ is below a given bound (depending on $q$) the decryption algorithm works. Otherwise, the plaintext cannot be recovered due to excessive noise growth. Therefore, as previously mentioned, tracking the size of this critical quantity is essential to ensure correct decryption.
To provide a clearer picture of what happens to the coefficients of $\nu$, we focus on multiplication, which is the homomorphic operation that highlights most clearly and significantly the dependencies among the critical quantities.
The BGV public key $\mathsf{pk}\in\mathcal{R}_{q}\times\mathcal{R}_{q}$ consists of two polynomials $(-a\cdot s+te,a)$, where $s$ is the secret key, $t$ is the plaintext modulus, $a\in\mathcal{R}_{q}$ is randomly chosen and $e\in\mathcal{R}_{q}$ is the error sampled from a discrete Gaussian distribution $\chi_{e}$. Roughly speaking, when two ciphertexts are multiplied — even if they were independently computed — their noises share some common terms which affect the resulting critical quantity $\nu_{\mathsf{mult}}$. More specifically, we observed that the noise in the ciphertexts contains terms that include powers of the secret key $s$ and powers of the term $e$. Note that these terms are common to all ciphertexts calculated using the same public key and are responsible for the dependence of the noise.
To account for these dependencies, we applied the correction function $F$ introduced in [2]. It is important to note that the case of BGV is more complex than that of BFV. In BFV, the dependency due to $e$ appears only in a negligible term. In contrast, for BGV, we must take into account the dependencies on both $s$ and $e$. This requires the use of two distinct correction functions to accurately model and mitigate these dependencies.
The paper is structured as follows. Section 2 introduces essential definitions and fundamental properties that are instrumental for understanding both our contribution and the scheme more broadly. Section 3 provides a concise overview of the main features and structure of the BGV scheme. In Section 4, we present our key results concerning the behavior of the error term and its growth under homomorphic operations. Section 5 then demonstrates how the findings from Section 4 can be leveraged to estimate error growth in fixed-operation circuits, laying the groundwork for novel approaches to selecting the scheme’s initial parameters. Finally, Section 6 concludes the paper and outlines possible directions for future research inspired by our results.
2 Preliminaries
In this section, we define the general notations that we will use in the remainder of the work.
In our analysis we will describe the leveled version of the BGV scheme, noting that this implies that the parameters of the scheme depend (polynomially) on the depth of the circuits that the scheme is capable of evaluating [4]. It should be noted that this version is also the most widely adopted because it does not require the usage of the bootstrapping technique, which is highly complex from a computational perspective. Before recalling the BGV scheme, some preliminary parameters must be fixed.
Let $L$ denote the number of levels of the arithmetic circuit that the scheme must be able to evaluate. Moreover, assume that $m$ is a suitably chosen integer. The notation $\mathcal{R}_{d}$, for a fixed $d\in\mathbb{Z}$, will denote the ring $\mathcal{R}_{d}=\mathbb{Z}_{d}[x]/\langle\Phi_{m}(x)\rangle$, where $\Phi_{m}(x)$ is the $m$-th cyclotomic polynomial, a definition of which is briefly recalled below.
Definition 1
Given a positive integer $m$, the $m$-th cyclotomic polynomial is defined as
$\Phi_{m}(x)=\prod_{\begin{subarray}{c}1\leq j<m\\ \gcd(j,m)=1\end{subarray}}(x-\zeta^{j}),$
where $\zeta$ is a primitive $m$-th root of unity. This polynomial has degree $\phi(m)$, where $\phi$ denotes the Euler function.
Eventually, we will denote with $\mathcal{R}$ the ring $\mathcal{R}=\mathbb{Z}[x]/\langle\Phi_{m}(x)\rangle$. To define the plaintext and ciphertext spaces, $L+1$ moduli are selected as follows:
  • An integer $t$, denoted as the plaintext modulus, is chosen such that $t\equiv 1\mod m$.
  • $L$ integers $q_{l}$, for $l\in\{0,\dots,L-1\}$, are defined as:
    $q_{l}=\prod_{j=0}^{l}p_{j},$
    where the $p_{j}$ are primes satisfying $p_{j}\equiv 1\mod m$ for $j=0,\dots,L-1$ and $p_{1},...,p_{L-2}$ have approximately the same size. Each $q_{L-1-l}$ is referred to as the ciphertext modulus for the $l$-th level.
According to these notations, we can define the plaintext and ciphertext spaces, respectively, as $P=\mathcal{R}_{t}$, $C=\mathcal{R}_{q_{l}}\times R_{q_{l}}$.
Due to the ciphertext coefficients being reduced according to different moduli at each level, BGV is commonly referred to as scale-dependent.
We use lowercase letters such as $x$ to denote polynomials, and bold symbols like $\bm{x}$ to represent vectors of polynomials. Given a polynomial $x\in\mathcal{R}_{n}$, we use the notation $[x]_{n}$ to refer to the centered representation of $x$ modulo $n$ i.e., the representative of $x$ with coefficients in $[-\frac{n}{2},\frac{n}{2})$. In general, unless otherwise specified, we always assume the coefficients of polynomials in $\mathcal{R}_{d}$ to be centered with respect to the modulus $d$.
Moreover, given a random polynomial $p\in\mathcal{R}$ and a probabilistic distribution $X$, the notation $p\leftarrow X$ is used to indicate that each coefficient of $p$ is chosen independently according to $X$.
Some distributions that will be frequently considered are the following:
  • $\mathcal{U}_{q}$ as the uniform distribution over $\nicefrac{{\mathbb{Z}}}{{q\mathbb{Z}}}$;
  • $\mathcal{N}(0,\sigma^{2})$ as the normal distribution, also referred to as the Gaussian distribution, over $\mathbb{R}$, with mean 00 and variance $\sigma^{2}$;
  • $\mathcal{DG}_{q}(\sigma^{2})$ as the discrete Gaussian distribution, which involves sampling a value according to $\mathcal{N}(0,\sigma^{2})$, rounding it to the nearest integer and then reducing it modulo $q$.
    Moreover, the representative modulo $q$ is taken in the interval $\left[-\frac{q}{2},\frac{q}{2}\right)$;
For the BGV scheme the notation $\chi_{e},\chi_{s}$ will be adopted in order to indicate the distribution of the error for a RLWE instance and the secret key coefficients, respectively. Typically, $\chi_{e}$ is a discrete Gaussian distribution with a suitable standard deviation, while for the secret key, the ternary uniform distribution $\mathcal{U}_{3}$ is usually considered. However, in some special cases, where bootstrapping is needed, the choice for the secret key distribution falls on the Hamming Weight distribution, a definition of which is recalled below [2].
Definition 2
The Hamming Weight distribution $\mathcal{HWT}_{n,h}$ consists of selecting uniformly at random a vector of $n$ entries in $\{-1,0,1\}$, where exactly $h<n$ entries are non-zero.
To ensure the correctness of the scheme, it is important to underline that the error added to the plaintext during encryption must be sufficiently small [4].
This will become more apparent through the presentation of the BGV functions.
For the purposes of our subsequent analysis, it is helpful to recall that for two polynomials $r,s\in\mathcal{R}$, and denoting by $r|_{i}$ and $s|_{i}$ their coefficients with respect to $x^{i}$, the $i$-th coefficient of their product is given by [2]
$rs|_{i}=\sum_{j=0}^{\phi(m)-1}\xi(i,j)r|_{j}s|_{i-j},$ (1)
where
$\xi(i,j)=\left\{\begin{matrix}1&\text{\hskip 51.21504ptfor }i-j\in[0,\phi(m))\\[5.0pt] -1&\text{otherwise}\end{matrix}\right.$
Furthermore, the following notation will be adopted.
Notation 1
Given a polynomial $p\in\mathcal{R}_{q}$, for a fixed integer $q$, the variance of its coefficients, namely $\mathsf{Var}(p|_{i})$, is denoted as $V_{p}$.
Proposition 1
Let $p,q$ be two independent polynomials in $\mathcal{R}_{q}$, where $q$ is a fixed integer, and let $k$ be an integer in $\nicefrac{{\mathbb{Z}}}{{q\mathbb{Z}}}$. Moreover, assume that the coefficients of each polynomial are independent and identically distributed, with 00-mean. Then,
  • a.
    $V_{p+q}=V_{p}+V_{q}$
  • b.
    $V_{kp}=k^{2}V_{p}$
  • c.
    $V_{p\cdot q}=\phi(m)V_{p}\cdot V_{q}$
Finally, the variance values for some common distributions, which will often be employed in the BGV scheme, are:
  • $V_{\mathcal{DG}_{q}(\sigma^{2})}=\sigma^{2}$ for the discrete Gaussian distribution centered at 0 with standard deviation $\sigma$;
  • $V_{3}=\frac{2}{3}$ for the ternary distribution $\mathcal{U}_{3}$;
  • $V_{q}=\frac{q^{2}-1}{12}\approx\frac{q^{2}}{12}$ for the uniform distribution over integer values in $\left[-\frac{q}{2},\frac{q}{2}\right)$;
3 The BGV scheme
Once the parameters have been specified, the framework of the BGV scheme can be defined according to three main functions: Key Generation, Encryption and Decryption.
3.1 Key Generation
The key generation function is responsible for producing:
  • the secret key, represented by a polynomial in $\mathcal{R}_{q_{L-1}}$;
  • the public key, which consists of a pair of polynomials, each belonging to $\mathcal{R}_{q_{L-1}}$;
To achieve this, three polynomials in $\mathcal{R}_{q_{L-1}}$ are generated according to the following distributions: $s\leftarrow\chi_{s}$ , $a\leftarrow\mathcal{U}_{q_{L-1}}$ and $e\leftarrow\chi_{e}$. The resulting keys are given by
$\begin{cases}sk&=s\\ pk&=(b,a)\equiv(-a\cdot s+te,a)\mod q_{L-1}\end{cases}$
3.2 Encryption
Given the plaintext $m\in\mathcal{R}_{t}$ and the public key $pk=(b,a)$, the encryption function returns as output $(\bm{c},q_{l},\nu_{clean})$ where:
  • $\bm{c}\in\mathcal{R}_{q_{l}}\times\mathcal{R}_{q_{l}}$ is the actual ciphertext defined as
    $\bm{c}=(c_{0},c_{1})\equiv(b\cdot u+te_{0}+m,a\cdot u+te_{1})\mod q_{l},$
    where $u,e_{0},e_{1}\in\mathcal{R}_{q_{l}}$, with coefficients distributed as $u\leftarrow\chi_{s}$ and $e_{0},e_{1}\leftarrow\chi_{e}$.
  • $q_{l}$ represents the ciphertext modulus for the level at which computations are performed. Specifically, as the encryption function is computed in the first level 00, the corresponding $q_{l}$ is $q_{L-1}$.
  • $\nu$ is the critical quantity associated to the ciphertext and is defined as
    $\nu=[c_{0}+c_{1}\cdot s]_{q_{l}}.$
We refer to the triple $\tilde{c}=(\bm{c},q_{l},\nu_{clean})$ with the term extended ciphertext.
Broadly speaking, the encryption of the plaintext message results in a pair of ciphertext values: the first encodes the masked plaintext using the public key, while the second supplies auxiliary information necessary for the receiver to recover the original message, provided they possess the secret key.
Regarding the critical quantity, it is important to emphasize that this value is strictly related to the error introduced by the encryption operator. Thus, it essentially determines whether the ciphertext can be correctly decrypted or not. For this reason, its behavior is analyzed to gain a clear understanding of the functioning of the scheme.
3.3 Decryption
At any level, it should be possible to recover the original plaintext using the decryption function.
Given the extended ciphertext $(\bm{c},q_{l},\nu)$ and the secret key $sk=s$, the plaintext is recovered performing the following computations
$m=\left[\left[c_{0}+c_{1}\cdot s\right]_{q_{l}}\right]_{t}.$
Using the previously established notation, the critical quantity can be rewritten as
$[c_{0}+c_{1}\cdot s]_{q_{l}}=\left[m+t(e\cdot u+e_{1}\cdot s+e_{0})\right]_{q_{l}}=\left[m+t\epsilon\right]_{q_{l}},$
where $t\epsilon$ denotes the error introduced during encryption.
By considering the reduction modulo $t$ of the critical quantity, it is possible to verify that the plaintext is successfully recovered. However, if the error is too large, the value $m+t(e\cdot u+e_{1}\cdot s+e_{0})$ could wrap around the modulus, resulting in an incorrect decryption.
So, decryption is correct only if the error magnitude remains below a certain threshold, as previously emphasized. In addition, the error associated to the ciphertext $\bm{c}$ at level $l$ increases each time the homomorphic circuit is evaluated. Therefore, for constructing a leveled homomorphic encryption scheme, it is crucial to estimate and manage this error properly. To deal with the error effectively, the magnitude of the critical quantity, often referred to as noise, is typically analyzed using its norm. In the context of homomorphic encryption, the infinity and the canonical norm, a definition of which are briefly recalled below [19], assume a central role.
Definition 3
The infinity norm of a polynomial $p\in\mathcal{R}$, of the form $p=p|_{0}+p|_{1}x+...+p|_{\phi(m)-1}x^{\phi(m)-1}$ is defined as
$\|p\|_{\infty}=\max_{0\leq i<\phi(m)}|\hskip 2.84544ptp|_{i}\hskip 2.84544pt|.$
Definition 4
The canonical embedding norm of a polynomial $p\in\mathcal{R}$ is defined as
$\|p\|_{can}=\max_{\begin{subarray}{c}1\leq j<m\\ \gcd(j,m)=1\end{subarray}}|p(\zeta^{j})|,$
where $\zeta$ is a primitive $m$-th root of unity.
Essentially, it corresponds to the infinity norm of the canonical embedding of $p$, denoted as $\sigma(p)$, which is the $\phi(m)$-vector defined as
$\sigma(p)=(p(\zeta^{i}))_{i},$
where $1\leq i<m$ and $\gcd(j,m)=1$.
A relationship between these two norms is given by the following property
$||p||_{\infty}\leq\alpha_{m}||p||_{can}\hskip 28.45274pt\forall p\in\mathcal{R},$
where $\alpha_{m}$ is a constant related to the ring $\mathcal{R}$, which depends exclusively on $m$. Another property, which will be useful in the context of the noise growth analysis, is recalled below [15].
Proposition 2
Let $p,q$ be two polynomials in $\mathcal{R}$. Then,
$||p\cdot q||_{\infty}\hskip 3.41418pt\leq\phi(m)||p||_{\infty}\cdot||q||_{\infty}$
$||p\cdot q||_{can}\leq||p||_{can}\cdot||q||_{can}$
In light of this, to guarantee the correctness of the decryption, the condition on the critical quantity can be expressed as follow
$||\nu||_{\infty}\leq\alpha_{m}||\nu||_{can}<\frac{q_{l}}{2},$
which allows to avoid the error wrapping around the modulus [24]. Naturally, in order to bound the noise, any type of norm could be used. Nevertheless, in the state-of-the-art analysis of noise growth for homomorphic schemes, the canonical norm is typically preferred, due to its properties, as it permits obtaining more accurate estimates. Finally, another concept that is often introduced for the error growth estimation is the noise budget.
Definition 5
Let $(\bm{c},q_{l},\nu)$ be an extended ciphertext. The noise budget associated to $\bm{c}$ is the quantity
$\log_{2}(q_{l})-\log_{2}(||\nu||)-1,$
where |||||$\cdot$||| | ⋅ | | refers to a fixed norm.
The key operations involved in the BGV scheme are briefly summarized in the following.
3.4 Ciphertext Addition and Constant Multiplication
Assume that $\bm{c}=(c_{0},c_{1})$ and $\bm{c^{\prime}}=(c_{0}^{\prime},c_{1}^{\prime})$ are two ciphertexts defined with respect to the same modulus $q_{l}$ and computed encrypting two plaintexts $m$ and $m’$ ’, respectively, using the same key. Their homomorphic sum is defined as
$\mathsf{Add}(\bm{c},\bm{c^{\prime}}):=(c_{0}+c_{0}^{\prime},\;c_{1}+c_{1}^{\prime})\mod q_{l}.$
The resulting ciphertext encrypts the sum $m+m^{\prime}$, and the corresponding noise grows additively. In fact, the resulting critical quantity is defined as
$\nu_{\mathsf{add}}=\nu+\nu^{\prime},$
where $\nu=c_{0}+c_{1}s$ and $\nu^{\prime}=c_{0}^{\prime}+c_{1}^{\prime}s$ denote the critical quantities of $\bm{c}$ and $\bm{c^{\prime}}$, respectively.
Similarly, let $\bm{c}=(c_{0},c_{1})$ be a ciphertext and $k\in R_{t}$ a constant polynomial. The homomorphic multiplication by $k$ is performed as follows
$\mathsf{Mul}_{k}(\bm{c}):=(kc_{0},\;kc_{1})\mod q_{l},$
and the corresponding critical quantity scales by $k$, yielding
$\nu_{\mathsf{const}}=k\cdot\nu.$
In both cases, decryption correctness is preserved as long as the resulting noise does not cause the message to wrap around the modulus.
3.5 Homomorphic multiplication and Key switching
Homomorphic multiplication is one of the fundamental operations supported by the BGV scheme. Given two ciphertexts $\bm{c}=(c_{0},c_{1})$ and $\bm{c^{\prime}}=(c_{0}^{\prime},c_{1}^{\prime})$, and assuming that both are defined with respect to the same modulus $q_{l}$, their product is given by
$\mathsf{Mul}(\bm{c},\bm{c^{\prime}}):=(c_{0}^{\mathsf{mul}},c_{1}^{\mathsf{mul}},c_{2}^{\mathsf{mul}})=(c_{0}\cdot c_{0}^{\prime},c_{0}\cdot c_{1}^{\prime}+c_{1}\cdot c_{0}^{\prime},c_{1}\cdot c_{1}^{\prime})\mod q_{l}.$
As a result, the ciphertext expands from two to three polynomials, which violates the compactness property and makes subsequent operations more costly. This phenomenon becomes even more problematic when performing multiple multiplications. Recovering the message from $\mathsf{Mul}(\bm{c},\bm{c^{\prime}})$ requires computing the reduction modulo $t$ of the resulting critical quantity given by $\nu_{\mathsf{mul}}=c_{0}^{\mathsf{mul}}+c_{1}^{\mathsf{mul}}s+c_{2}^{\mathsf{mul}}s^{2}$.
It is significant to note that, unlike other operations, multiplication causes the noise to grow multiplicatively, making it the most critical case to handle among the basic homomorphic operations in terms of error growth.
To address this, a technique known as relinearization, or key switching, is employed. The key idea is to reduce the ciphertext back to its original size while preserving its correctness. In practice, this means converting a ciphertext of the form $(c_{0},c_{1},c_{2})$ into a new, equivalent one of the form $(\hat{c}_{0},\hat{c}_{1})$, using auxiliary data called the key switching key. More intuitively, key switching allows the expression $c_{0}+c_{1}s+c_{2}s^{2}$ to be re-expressed as a linear polynomial in $s$, restoring the standard noise form.
Although the relinearization step introduces additional noise, it is essential for maintaining the practicality of the scheme. Several key switching techniques have been proposed in the literature, each aiming to optimize this trade-off between correctness and noise growth. The most commonly used are the Brakerski Vaikuntanathan (BV) variant, the Gentry Halevi Smart (GHS) variant and the Hybrid variant, which can be considered as a combination of the previous ones.
3.6 Modulus switching
The primary aim of the modulus switching technique is to reduce the noise resulting from homomorphic evaluations. In practical circuits, modulus switching is typically applied only after homomorphic multiplications, due to their higher cost in terms of error growth. Similarly to the bootstrapping technique, the main purpose is to transition from an extended ciphertext $(\bm{c},q_{l},\nu)$ to another $(\bm{c}^{\prime},q_{l^{\prime}},\nu^{\prime})$ such that the error associated to the latter is smaller, while preserving the fact that, using the same key, both can be decrypted to the same original message. The key difference lies in how this is achieved, and it is in this aspect that modulus switching demonstrates greater efficiency compared to bootstrapping.
Let $(\bm{c},q_{l},\nu)$ be the extended ciphertext, whose error we want to reduce, and let $l^{\prime}$ be an integer, such that $q_{l^{\prime}}<q_{l}$. The new ciphertext produced by the modulus switch is given by
$\bm{c}^{\prime}=\frac{q_{l^{\prime}}}{q_{l}}(\bm{c}+\bm{\delta})\mod q_{l^{\prime}},$
where $\bm{\delta}=t[-\bm{c}t^{-1}]_{\frac{q_{l}}{q_{l^{\prime}}}}$. The $\bm{\delta}$ value can be interpreted as a correction required to ensure that the ciphertext is divisible by $\frac{q_{l}}{q_{l^{\prime}}}$ and does not affect the original message. In fact, it only influences the error since $\bm{\delta}\equiv 0\mod t$. Therefore, the new ciphertext $\bm{c}^{\prime}$ will still decrypt to the original plaintext (scaled by a factor of $\frac{q_{l}}{q_{l^{\prime}}}$), provided that the necessary conditions are met.
The critical quantity associated to the new ciphertext $\bm{c^{\prime}}$ can be expressed in terms of that of $\bm{c}$, namely $\nu$, as
$\nu_{\mathsf{modswitch}}$ $=[c_{0}^{\prime}+c_{1}^{\prime}\cdot s]_{q_{l^{\prime}}}=\frac{q_{l^{\prime}}}{q_{l}}([c_{0}+c_{1}\cdot s]_{q_{l}}+\delta_{0}+\delta_{1}\cdot s)=\frac{q_{l^{\prime}}}{q_{l}}(\nu+\delta_{0}+\delta_{1}\cdot s).$
To conclude, it is worth noting that this correctness verification implicitly relies on the common value $h$ in the two equalities
$\begin{split}[c_{0}+c_{1}\cdot s]_{q_{l}}\hskip 2.84544pt&=c_{0}+c_{1}\cdot s-hq_{l}\\ [c_{0}^{\prime}+c_{1}^{\prime}\cdot s]_{q_{l^{\prime}}}&=c_{0}^{\prime}+c_{1}^{\prime}\cdot s-hq_{l^{\prime}}\end{split}$
The reason why the value $h$ is the same for both the expressions can be found in [4, Lemma 1].
4 Average-Case Noise Analysis for BGV
The aim of this section is to investigate the behavior of the noise resulting from the main homomorphic operations supported by the BGV scheme. Throughout this analysis, we consider ciphertexts that are mutually independent, obtained by encrypting independently generated random messages under the same public key.
As previously mentioned, the novel approach introduced in this paper seeks to analyze noise growth by accounting for dependencies among the coefficients of the critical quantities involved. Before delving into the details, it is important to highlight that in BGV, the critical quantity resulting from homomorphic operations can be affected by such dependencies, even when the ciphertexts involved are independent. These dependencies stem from the fact that the noise in the ciphertexts includes terms involving powers of the secret key $s$ and powers of the error term $e$, which makes it necessary to explicitly consider these contributions when analyzing the variance of the noise.
To study the impact of $s$ and $e$, we isolate their contribution in the expression of the critical quantity $\nu$, using the following notation:
$\nu=\sum_{\iota}a_{\iota}s^{\iota}=\sum_{\iota}\sum_{\mu}b_{\mu}(\iota)e^{\mu}s^{\iota},$
where $a_{\iota}=\sum_{\mu}b_{\mu}(\iota)e^{\mu}$, and $b_{\mu}(\iota)$ contains no powers of $s$ or $e$.
To enhance clarity, for the critical quantity $\nu$ of a fresh ciphertext $\bm{c}$, this notation yields
$\nu=a_{0}+a_{1}s=b_{0}(0)+b_{1}(0)e+b_{0}(1)s,$
where
$\begin{cases}a_{0}=b_{0}(0)+b_{1}(0)\cdot e=(m+te_{0})+tu\cdot e\\ a_{1}=b_{0}(1)=te_{1}\end{cases}$
Before introducing our method for studying the growth of error through its variance in fixed circuits, we begin by presenting some considerations and results we have derived regarding the distribution of the coefficients of a generic error term, along with certain properties related to their variances. We then proceed to describe how these properties are used to estimate the variance of the error coefficients after homomorphic multiplications, and finally how such estimates can be applied to circuits consisting of fixed sequences of operations.
From this point onward, we adopt the notation $n$ to denote $\phi(m)$, in order to simplify the presentation. It is interesting to point out that when $n$ is a power of two, $m=2n$.
4.1 Distribution Analysis
As previously mentioned, the noise coefficients are treated as random variables. We observed that their distributions are well-approximated by identically distributed, dependent Gaussian variables centered at zero. This empirical observation was validated through experiments using the Python fitter package, which confirmed the Gaussian nature of the noise coefficients. This property is particularly advantageous, as it allows us to bound the maximum absolute value of the noise coefficients with high probability by simply controlling their variance $V$. In particular, we recall the following fundamental property of Gaussian random variables.
Proposition 3
Let $Z$ be a real-valued, zero-mean Gaussian random variable with variance $V$. Then, for any fixed value $z>0$, the probability that $Z$ lies within the interval $(-z,z)$ is given by
$\mathbb{P}(|Z|<z)=\text{erf}\left(\frac{z}{\sqrt{2V}}\right),$
where $\text{erf}(z)$ is the error function, defined as
[…]
Extending this result to a vector $\bm{Z}\in\mathbb{R}^{n}$, whose entries are distributed as $Z$, we obtain the following bound on the tail probability of the infinity norm
$\mathbb{P}(\|\bm{Z}\|_{\infty}>z)\leq n\left(1-\text{erf}\left(\frac{z}{\sqrt{2V}}\right)\right).$ (2)
This implies that, in order to ensure that a noise vector $\nu$ satisfies the correctness condition $\|\nu\|_{\infty}<{q}/{2},$ we can use Proposition 3 to bound the failure probability as
$\mathbb{P}\left(\|\nu\|_{\infty}>\frac{q}{2}\right)\leq n\left(1-\text{erf}\left(\frac{q}{2\sqrt{2V}}\right)\right),$
where $V$ denotes the estimated variance of the noise coefficients. To express this bound more conveniently, we introduce a security parameter $D$ such that
$D\leq\frac{q}{2\sqrt{2V}},$
from which it follows, using the monotonicity of the error function, that
$\mathbb{P}\left(\|\nu\|_{\infty}>\frac{q}{2}\right)\leq n\left(1-\text{erf}(D)\right).$
Thus, by appropriately choosing the security parameter $D$, we can ensure that the probability of decryption failure remains negligibly small. For instance, setting $D=6$ and $n=2^{13}$ yields a failure probability of approximately $2^{-42}$. For practical applications, $D=8$ should be preferred, resulting in a probability of approximately $2^{-83}$, when the ring dimension is $n=2^{13}$.
Consequently, the ciphertext modulus $q$ can be selected according to the inequality
$q\geq 2D\sqrt{2V},$ (3)
which highlights the importance of accurately estimating $V$.
It is important to emphasize that the bounds derived in our analysis provide insight into the minimum ciphertext modulus $q$ required to guarantee the correctness of the scheme. Although larger values of $q$ may be used in practice, identifying the minimal admissible value is essential for optimizing performance and efficiency. This is particularly relevant in the context of the BGV scheme, where our results can serve as a guideline for carefully selecting the setting parameters. By determining tight lower bounds on $q$, it is possible to lay the foundation for sound and efficient parameter selection that ensures correctness without over-provisioning resources.
4.2 Mean and Variance Analysis
As previously discussed, we have demonstrated that the coefficients of the error term are centered at zero. Furthermore, we have shown that the coefficients of the $b_{\mu}$ terms are uncorrelated, meaning that their pairwise covariance is zero.
Lemma 1
Let $\nu=\sum_{\iota}\sum_{\mu}b_{\mu}(\iota)e^{\mu}s^{\iota}$ be the critical quantity associated with a given ciphertext. Then, the following properties hold
  • a)
    $\mathsf{Cov}(b_{\mu_{1}}(\iota_{1})|_{j_{1}},b_{\mu_{2}}(\iota_{2})|_{j_{2}})=0$ for $\mu_{1}\neq\mu_{2}\text{ or }j_{1}\neq j_{2}$, $\forall\iota_{1},\iota_{2}$;
  • b)
    $\mathbb{E}[b_{\mu}(\iota)|_{i}]=0$, $\forall\iota,\mu,i$;
A proof of Lemma 1 can be found in Appendix 0.A
Lemma 2
Let $\nu=\sum_{\iota}a_{\iota}s^{\iota}$ represent the critical quantity associated with a given ciphertext, where, for a fixed $\iota$, $a_{\iota}=\sum_{\mu}b_{\mu}(\iota)e^{\mu}$. Then, the following equivalence holds
$\mathsf{Var}(a_{\iota}s^{\iota}|_{i})=\sum_{\mu\geq 0}\mathsf{Var}(b_{\mu}(\iota)|_{i})\sum_{k=0}^{n-1}\mathbb{E}[e^{\mu}|_{k}^{2}]\sum_{j=0}^{n-1}\mathbb{E}[s^{\iota}|_{j}^{2}]$
A proof of Lemma 2 can be found in Appendix 0.B.
4.3 Homomorphic multiplications
The main goal of our analysis is to provide estimates on the growth of the error resulting from homomorphic multiplications, which are the most complex operations and therefore the most interesting to study. It thus becomes crucial to account for the dependencies that arise among the noise coefficients as a consequence of these multiplications. To this end, the approach we adopted relies on the correction functions introduced in [2].
Therefore, before stating the theorem, some of the key properties of $F$ will be recalled. Unlike the BFV scheme, the BGV scheme requires additional care in the use of the correction function $F$, as not only the impact of the term $s$ must be taken into account, but also that of the term $e$, which is negligible in BFV. Consequently, our analysis requires the introduction of two distinct correction functions, which will be denoted as $F_{e}$ and $F_{s}$. To define these two functions, we need the following
Heuristic 1
Let $x$ be a polynomial whose coefficients are independent and identically distributed according to a fixed distribution. The function $f_{x}$ defined as
$f_{x}(\iota)=-e^{\alpha-\beta\iota-\gamma\iota^{2}}+\delta,$ (4)
where $\alpha,\beta,\gamma,\delta$ depend only on the coefficient distribution of $x$ and the ring dimension $n$, satisfies the following
$f_{x}(\iota)\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[x^{\iota}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[x^{\iota-1}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[x|_{i_{2}}^{2}]}.$
In Table 1, we report the values of $\alpha,\beta,\gamma,\delta$ required to define the function $f_{s}$, where $s$ denotes the secret key with coefficients sampled from the ternary distribution $\chi_{s}=\mathcal{U}_{3}$, as in [2].
        $n$         $\alpha$         $\beta$         $\gamma$         $\delta$
        $2^{12}$         2.8732         0.0160         0.0049         19.1895
        $2^{13}$         2.9644         0.0196         0.0046         20.4747
        $2^{14}$         2.9578         0.0386         0.0032         19.5755
        $2^{15}$         2.9765         0.0197         0.0043         20.7760
Table 1: Parameters for $\chi_{s}=\mathcal{U}_{3}$
Similarly, it is possible to define the function $f_{e}$ for the error term $e$, whose coefficients are distributed according to a discrete Gaussian distribution. For completeness, a table containing the parameters depending on the distribution $\chi_{e}$ can be found in Table 2 [2].
        $n$         $\alpha$         $\beta$         $\gamma$         $\delta$
        $2^{12}$         2.9000         0.0157         0.0051         19.5356
        $2^{13}$         2.9340         0.0042         0.0055         20.7063
        $2^{14}$         2.9138         0.0290         0.0039         19.2973
        $2^{15}$         2.9511         0.0129         0.0046         20.7263
Table 2: Parameters for $\chi_{e}=\mathcal{DG}_{q}(\sigma^{2})$, $\sigma=3.19$
Definition 6
Let $g_{x}(\iota)=\prod_{i=0}^{\iota}f_{x}(i)$ with $f_{x}$ defined as in (4) and $f_{x}(0)=f_{x}(1)=~{}1$. The correction function $F_{x}$ is
$F_{x}(\iota_{1},\iota_{2}):=\frac{g_{x}(\iota_{1}+\iota_{2})}{g_{x}(\iota_{1})\cdot g_{x}(\iota_{2})}$
Moreover, this function satisfies
$F_{x}(\iota_{1},\iota_{2})\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[x^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[x^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[x^{\iota_{2}}|_{i_{2}}^{2}]}$
A proof of this result can be found in [2]. For the purposes of our study, we denote by $F_{e}$ the correction function required to account for the dependencies induced by the term $e$, and by $F_{s}$ the correction function associated with the term $s$.
We can now introduce our main theorem.
Theorem 4.1
Let $\nu=\sum_{\iota}a_{\iota}s^{\iota},\nu^{\prime}=\sum_{\iota}a^{\prime}_{\iota}s^{\iota}$ be the critical quantities of two independently computed ciphertexts defined with respect to the same modulus $q$. Then
$\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})\leq n\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}})|_{i})\mathsf{Var}((a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})F_{s}(\iota_{1},\iota_{2})F_{e}(K_{1},K_{2}),$
where $K_{1},K_{2}$ arise from the expansions
$a_{\iota_{1}}=\sum_{\mu_{1}=0}^{K_{1}}b_{\mu_{1}}(\iota_{1})e^{\mu_{1}},\hskip 7.11317pta^{\prime}_{\iota_{2}}=\sum_{\mu_{2}=0}^{K_{2}}b^{\prime}_{\mu_{2}}(\iota_{2})e^{\mu_{2}}$
and represent the highest power of $e$ appearing in $a_{\iota_{1}},a^{\prime}_{\iota_{2}}$, respectively.
Proof
From lemma 2 it is possible to express the variance of two generic terms $a_{\iota_{1}}s^{\iota_{1}}|_{i}$ and $a_{\iota_{2}}^{\prime}s^{\iota_{2}}|_{i}$ as
$\begin{cases}\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})=\sum_{\mu_{1}=0}^{K_{1}}\mathsf{Var}(b_{\mu_{1}}(\iota_{1})|_{i})\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{2}}]\\ \mathsf{Var}(a_{\iota_{2}}^{\prime}s^{\iota_{2}}|_{i})=\sum_{\mu_{2}=0}^{K_{2}}\mathsf{Var}(b_{\mu_{2}}^{\prime}(\iota_{2})|_{i})\sum_{j_{3}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{3}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}]\end{cases}$
By observing that
$a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}}=\left(\sum_{\mu}\sum_{\mu_{1}+\mu_{2}=\mu}b_{\mu_{1}}(\iota_{1})b_{\mu_{2}}^{\prime}(\iota_{2})e^{\mu}\right)s^{\iota_{1}+\iota_{2}},$
and using lemma 2, it is possible to write the variance $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})$ as
$\sum_{\mu}\mathsf{Var}\left(\sum_{\mu_{1}+\mu_{2}=\mu}(b_{\mu_{1}}(\iota_{1})b_{\mu_{2}}^{\prime}(\iota_{2}))|_{i}\right)\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{1}+\iota_{2}}|^{2}_{j_{2}}]$
${=n\sum_{\mu}\sum_{\mu_{1}+\mu_{2}=\mu}\mathsf{Var}\left(b_{\mu_{1}}(\iota_{1})|_{i}\right)\mathsf{Var}\left(b_{\mu_{2}}^{\prime}(\iota_{2})|_{i}\right)\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{1}+\iota_{2}}|^{2}_{j_{2}}]},$
where the second equality follows from the independence of $b_{\mu_{1}}(\iota_{1}),b_{\mu_{2}}^{\prime}(\iota_{2})$ and from $\mathsf{Cov}(b_{\mu_{1}}(\iota)|_{j_{1}},b_{\mu_{2}}({\iota})|_{j_{2}})=0$ for $\mu_{1}\neq\mu_{2}\text{ or }j_{1}\neq j_{2}$.
Moreover, it can be noted that $n\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})\mathsf{Var}(a^{\prime}_{\iota_{2}}s^{\iota_{2}}|_{i})$ can be written as
$n\left(\sum_{\mu_{1}=0}^{K_{1}}\mathsf{Var}(b_{\mu_{1}}(\iota_{1})|_{i})\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{2}}]\right)\cdot$
$\cdot\left(\sum_{\mu_{2}=0}^{K_{2}}\mathsf{Var}(b^{\prime}_{\mu_{2}}(\iota_{2})|_{i})\sum_{j_{3}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{3}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}]\right),$
which can be reordered as
$n\sum_{\mu}\sum_{\mu_{1}+\mu_{2}=\mu}\mathsf{Var}(b_{\mu_{1}}(\iota_{1})|_{i})\mathsf{Var}(b^{\prime}_{\mu_{2}}(\iota_{2})|_{i})$
$\hskip 5.69046pt\cdot\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{3}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{3}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{2}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}],$
Recalling that the correction function $F$ approximates the following quantities
$\begin{cases}F_{e}(\iota_{1},\iota_{2})\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[e^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[e^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[e^{\iota_{2}}|_{i_{2}}^{2}]}\\ F_{s}(\iota_{1},\iota_{2})\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[s^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|_{i_{2}}^{2}]}\par\par\end{cases}$
It is possible to use these functions in order to express $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})$ in terms of $\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})\mathsf{Var}(a^{\prime}_{\iota_{2}}s^{\iota_{2}}|_{i})$. In fact, $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})$ can be written as
$n\sum_{\mu}\sum_{\mu_{1}+\mu_{2}=\mu}\mathsf{Var}\left(b_{\mu_{1}}(\iota_{1})|_{i}\right)\mathsf{Var}\left(b^{\prime}_{\mu_{2}}(\iota_{2})|_{i}\right)\sum_{j=0}^{n-1}\mathbb{E}[e^{\mu_{1}+\mu_{2}}|^{2}_{j}]\sum_{j^{\prime}=0}^{n-1}\mathbb{E}[s^{\iota_{1}+\iota_{2}}|^{2}_{j^{\prime}}].$
Thus, by leveraging the properties of the correction functions
$nF_{s}(\iota_{1},\iota_{2})$ $\sum_{\mu}\sum_{\mu_{1}+\mu_{2}=\mu}\mathsf{Var}\left(b_{\mu_{1}}(\iota_{1})|_{i}\right)\mathsf{Var}\left(b^{\prime}_{\mu_{2}}(\iota_{2})|_{i}\right)F_{e}(\mu_{1},\mu_{2})\cdot$
$\cdot$ $\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{2}}]\sum_{j_{3}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{3}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}]$
[…]
$\cdot$ $\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{2}}]\sum_{j_{3}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{3}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}].$
Then, by exploiting the monotonicity of $F_{e}$ and $F_{s}$ [2], it is possible to derive an upper bound given by
$nF_{s}(\iota_{1},\iota_{2})F_{e}(K_{1},K_{2})$ $\sum_{\mu}\sum_{\mu_{1}+\mu_{2}=\mu}\mathsf{Var}\left(b_{\mu_{1}}(\iota_{1})|_{i}\right)\mathsf{Var}\left(b^{\prime}_{\mu_{2}}(\iota_{2})|_{i}\right)\cdot$
[…]
where $K_{1},K_{2}$ represent the highest power of $e$ appearing in $a_{\iota_{1}},a^{\prime}_{\iota_{2}}$, respectively.
It is straightforward to verify that this concludes our proof, yielding
$\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})=nF_{s}(\iota_{1},\iota_{2})F_{e}(K_{1},K_{2})\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})\mathsf{Var}(a^{\prime}_{\iota_{2}}s^{\iota_{2}}|_{i}).$
Remark 1
Theorem 4.1 provides an upper bound on the variance of the product $a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}}$. However, it is significant to point out that this estimate can be regarded as a faithful approximation, i.e.,
$\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})\approx n\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}})|_{i})\mathsf{Var}((a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})F_{s}(\iota_{1},\iota_{2})F_{e}(K_{1},K_{2}).$
The reason behind this statement can be immediately found in the distributions of the $b_{\mu}$-terms in the noise of a fresh ciphertext. In fact, for a fresh ciphertext, the critical quantity can be expressed as
$\nu=a_{0}+a_{1}s=(b_{0}(0)+b_{1}(0)\cdot e)+b_{0}(1)s,$
where
$\begin{cases}a_{0}=b_{0}(0)+b_{1}(0)\cdot e=(m+te_{0})+tu\cdot e\\ a_{1}=b_{0}(1)=te_{1}\end{cases}$
Then, it can be observed that the variance of the $b_{\mu}$-values is given by
$\begin{cases}\mathsf{Var}(b_{0}(0)|_{i})=V_{m}+t^{2}V_{e}\\ \mathsf{Var}(b_{1}(0)|_{i})=nt^{2}V_{e}V_{u}\\ \mathsf{Var}(b_{0}(1)|_{i})=nt^{2}V_{e}V_{s}\end{cases}$
where $n$ typically assumes values in $\{2^{13},2^{14},2^{15},2^{16}\}$. Consequently,
$\mathsf{Var}(b_{0}(0)|_{i})\ll\mathsf{Var}(b_{1}(0)|_{i}),\mathsf{Var}(b_{0}(1)|_{i}).$
Therefore, $b_{0}(0)$ becomes negligible in our analysis, and we can assume $a_{0}\approx b_{1}(0)\cdot e$, obtaining the approximated version of the theorem.
However, in order to maintain our results as general as possible, our next analysis will refer to the upper bound version of the theorem, still knowing that it corresponds to a reliable approximation.
5 Application of Our Variance Estimation Method to Fixed Circuits
This section employs the results of Section 4 to propose an effective method for tracking the error growth in circuits with a fixed number of operation. As previously highlighted, obtaining bounds sufficiently tight with respect to the experimental values of the error coefficients’ variance is crucial for development and real use of homomorphic encryption.
These bounds, in fact, allow one to construct new techniques for efficiently finding initial scheme parameters that improve performance and security simultaneously. Naturally, the behavior of the error is very dependent on the specific circuit under consideration, that is, the sequence of operations that the scheme must accommodate.
We analyze a circuit in which pairs of ciphertexts are progressively multiplied. We focus on circuits that primarily involves homomorphic multiplication, since, as previously mentioned, it is the most complex and therefore the most significant operation to study.
For clarity, we divide the analysis into two distinct scenarios. In the first scenario, only theoretical, the modulus switch technique is not applied, resulting in a very rapid increase of the error. Instead, in the second case, the modulus switch is performed after each multiplication.
Typically, the first scenario is not taken into account, as the error magnitude, without modulus switching, quickly becomes unmanageable, making the scheme impractical for real-world applications. In contrast, with modulus switching, error growth is significantly diminished, and the noise remains within a manageable range. This is, in fact, one of the primary goals pursued in designing practical FHE schemes.
However, the application of modulus switch drastically reduces the differences in estimates derived from even significantly distant approaches. Instead, by studying circuits without the modulus switch, we believe that the necessity of considering the dependencies in the coefficients of the error polynomial becomes unequivocal.
For this reason, although the absence of modulus switching does not reflect a realistic scenario, we decided to start our analysis in this simplified context. This allows us to highlight the key structural aspects of error growth, which are then carried over and extended to the more realistic case involving modulus switching.
Again, we want to stress that circuits without modulo switch are not practicable and do not have real world applications. The choice to describe this scenario is purely to complete the explanation of our method, further emphasizing the differences between our approach and others considered.
The remainder of this section is structured as follows. In the first part, we present circuits without modulus switching. Subsequently, we focus on practical circuits that employ modulus switching, showing how the results from Section 4 can be applied in this setting.
Based on the results, we are confident that our estimates represent a tangible improvement over the current state of the art. Our bounds closely match the experimental results, while consistently avoiding underestimation, which affects instead the existing approaches and must be avoided for the proper functioning of the scheme together with its security.
5.1 Circuits without Modulo Switch
Figure 1: Reference circuit
Figure 1 provides a schematic representation of the circuit used as a reference for our analysis. Let $M$ denote the multiplicative depth of the circuit, that is, the number of multiplications to be supported. To facilitate the analysis, we divide the circuit into $L=M+1$ levels.
For the considered circuit, we assume the independent and random generation of $2^{L-1}$ plaintexts, which are then encrypted using the same public key, resulting in $2^{L-1}$ fresh ciphertexts, denoted by $\bm{c}_{1},\dots,\bm{c}_{2^{L-1}}$.
At each level $l\in[0,\dots,L-1]$, pairs of ciphertexts are multiplied, yielding the corresponding ciphertexts of the $(l+1)$-th level, and then the process is repeated until the final ciphertext is computed.
Our study aims to estimate the variance of the coefficients of the critical quantity $\nu^{(l)}|_{i}$, denoted as $V_{l}$, which corresponds to the ciphertexts of the circuit at level $l$.
According to the current average-case approach, the variance at level $l$ can be approximated as
$V_{l}\leq nV_{l-1}^{2}.$
However, as pointed out by the authors of [25], this estimation tends to underestimate the variance observed in experimental evaluations. Such underestimation may lead to security vulnerabilities, as previously discussed, and, therefore, must be avoided.
In contrast, in our investigation, we proved that
$V_{l}\leq nh(l)V_{l-1}^{2},$ (5)
where $h$ is a function specifically designed to account for the dependencies among the coefficients of the error polynomial. This construction is based on the correction function $F$ introduced in [2] and represents the crucial difference compared to the current approach [25].
Section 4 has laid the groundwork for deriving this bound. Thus, the construction of our bound can now be presented, beginning with the initial levels and subsequently extending the reasoning to the general case for a level $l$.
First Multiplication
Let $\bm{c},\bm{c^{\prime}}$ be two fresh ciphertexts with critical quantities $\nu=a_{0}+a_{1}s$, $\nu^{\prime}=a^{\prime}_{0}+a^{\prime}_{1}s$, respectively, where
$\begin{cases}a_{0}=b_{0}(0)+b_{1}(0)\cdot e=(m+te_{0})+tu\cdot e\\ a_{1}=b_{0}(1)=te_{1}\end{cases}$
$\begin{cases}a^{\prime}_{0}=b^{\prime}_{0}(0)+b^{\prime}_{1}(0)\cdot e=(m^{\prime}+te^{\prime}_{0})+tu^{\prime}\cdot e\\ a^{\prime}_{1}=b^{\prime}_{0}(1)=te^{\prime}_{1}\end{cases}$
Let $V_{0}$ be the variance of the critical quantity associated with the fresh ciphertexts. Then, it can be observed that
$\begin{cases}V(a_{0}|_{i})=V(a^{\prime}_{0}|_{i})=t^{2}(\frac{1}{12}+V_{e}+nV_{e}V_{u})\\ V(a_{1}s|_{i})=V(a^{\prime}_{1}s|_{i})=t^{2}nV_{e}V_{s}\end{cases}$
Therefore, it is possible to assume that
$V(a_{0}|_{i})\approx V(a_{1}s|_{i})\approx\frac{V_{0}}{2},$ (6)
since $V_{u}=V_{s}$, due to the choice of the distributions of $s$ and $u$, and
$\frac{1}{12}+V_{e}\ll nV_{e}V_{u},$
which is then negligible, as highlighted in 1.
Clearly, the same reasoning can be applied to $V(a^{\prime}_{0}|_{i}),V(a^{\prime}_{1}s|_{i})$, obtaining
$V(a^{\prime}_{0}|_{i})\approx V(a^{\prime}_{1}s|_{i})\approx\frac{V_{0}}{2}.$ (7)
Our goal is to estimate the coefficients variance of $\nu_{\mathsf{mul}}=\nu\cdot\nu^{\prime}$, which will be referred to as $V_{1}$, following the level notation introduced at the beginning. Moreover, according to the previously introduced notation, it is possible to write $\nu_{\mathsf{mul}}$ as
$\nu_{\mathsf{mul}}=a_{0}a^{\prime}_{0}+(a_{0}a^{\prime}_{1}+a_{1}a^{\prime}_{0})\cdot s+a_{1}a^{\prime}_{1}\cdot s^{2}.$
Consequently, when the variance is considered, applying Theorem 4.1 and using that the covariance of this addends is zero, it follows that
$V_{1}:=\mathsf{Var}(\nu_{\mathsf{mul}}|_{i})$ $\leq n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i}){F_{s}(0,0)}{F_{e}(1,1)}$
$+n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}((a^{\prime}_{1}s)|_{i}){F_{s}(0,1)}{F_{e}(1,0)}$
$+n\mathsf{Var}((a_{1}s)|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i}){F_{s}(1,0)}{F_{e}(0,1)}$
$+n\mathsf{Var}((a_{1}s)|_{i}))\mathsf{Var}((a^{\prime}_{1}s)|_{i}){F_{s}(1,1)}{F_{e}(0,0)}.$
Moreover, this bound can be simplified, observing that $F_{s}(0,0)=F_{s}(1,0)=F_{s}(0,1)=1$ and $F_{e}(0,0)=F_{e}(1,0)=F_{e}(0,1)=1$. Therefore, it follows that
$V_{1}\leq$ $\,n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i}){F_{e}(1,1)}+n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}((a^{\prime}_{1}s)|_{i})$
$+n\mathsf{Var}((a_{1}s)|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i})+n\mathsf{Var}((a_{1}s)|_{i}))\mathsf{Var}((a^{\prime}_{1}s)|_{i}){F_{s}(1,1)}.$
Again, this can be reformulated using (6) and (7) as
$V_{1}$ $\leq\frac{nV_{0}^{2}}{4}\left(F_{e}(1,1)+{F_{s}(1,1)}+2\right).$
Hence, the first value of the function $h$ we aim to construct, can be fixed as
$h(1)=\frac{F_{s}(1,1)+{F_{e}(1,1)}+2}{4}.$
Thus, for the first level, the following inequality has been proven
$V_{1}\leq nh(1)V_{0}^{2}.$
Second Multiplication
In the second multiplication, the ciphertexts involved will have an associated critical quantity of the form
$\nu=a_{0}+a_{1}\cdot s+a_{2}\cdot s^{2},$
where, according to our study, each coefficient has variance
$\begin{cases}\mathsf{Var}(a_{0}|_{i})\approx\frac{F_{e}(1,1)V_{1}}{F_{s}(1,1)+{F_{e}(1,1)}+2}\\ \mathsf{Var}(a_{1}s|_{i})\approx\frac{2V_{1}}{F_{s}(1,1)+{F_{e}(1,1)}+2}\\ \mathsf{Var}(a_{2}s^{2}|_{i})\approx\frac{F_{s}(1,1)V_{1}}{F_{s}(1,1)+{F_{e}(1,1)}+2}\end{cases}$
Therefore, by repeating the same procedure depicted above, the critical quantity after the second multiplication will be of the form
$\nu\cdot\nu^{\prime}=a_{0}a^{\prime}_{0}+(a_{0}a^{\prime}_{1}+a^{\prime}_{0}a_{1})s+(a_{1}a^{\prime}_{1}+a_{0}a^{\prime}_{2}+a^{\prime}_{0}a_{2})s^{2}+(a_{1}a^{\prime}_{2}+a^{\prime}_{1}a_{2})s^{3}+a_{2}a^{\prime}_{2}s^{4}$
Thus, its coefficients’ variance can be estimated as
$V_{2}$ $\leq\frac{nV_{1}^{2}}{(2+F_{s}(1,1)+{F_{e}(1,1)})^{2}}\bigl{[}F_{e}(1,1)^{2}\cdot{F_{s}(0,0)}{F_{e}(2,2)}$
$\quad+4F_{e}(1,1)\cdot{F_{s}(0,1)}{F_{e}(1,2)}+4\cdot{F_{s}(1,1)}{F_{e}(1,1)}$
$\quad+2F_{s}(1,1)F_{e}(1,1)\cdot{F_{s}(0,2)}{F_{e}(0,2)}+4F_{s}(1,1)\cdot{F_{s}(1,2)}{F_{e}(0,1)}$
$\quad+F_{s}(1,1)^{2}\cdot{F_{s}(2,2)}{F_{e}(0,0)}\bigr{]},$
where this inequality follows from the fact that $a_{0}$ contains a term multiplied by $e^{2}$, $a_{1}$ contains a term multiplied by $e$ and $a_{2}$ does not contain any powers of $e$. As a result, $h(2)$ can be chosen as
$h(2)$ $=\frac{1}{(2+F_{s}(1,1)+{F_{e}(1,1)})^{2}}\bigl{[}F_{e}(1,1)^{2}\cdot{F_{s}(0,0)}{F_{e}(2,2)}$
$\quad+4F_{e}(1,1)\cdot{F_{s}(0,1)}{F_{e}(1,2)}+4\cdot{F_{s}(1,1)}{F_{e}(1,1)}$
$\quad+2F_{s}(1,1)F_{e}(1,1)\cdot{F_{s}(0,2)}{F_{e}(0,2)}$
$\quad+4F_{s}(1,1)\cdot{F_{s}(1,2)}{F_{e}(0,1)}+F_{s}(1,1)^{2}\cdot{F_{s}(2,2)}{F_{e}(0,0)}\bigr{]}.$
Consequently, $V_{2}\leq nh(2)V_{1}^{2}$.
Generalization
Finally, this procedure can be easily generalized in order to compute all the necessary values for $h$, therefore proving our bound in (5), i.e.,
$V_{l}\leq nh(l)V_{l-1}^{2},\hskip 28.45274ptl=1,\dots,L-1.$
5.2 Circuits with Modulo Switch
Figure 2: Reference circuit
In the present section, the reasoning previously introduced is adapted to circuits where the modulus switching technique is employed, in order to exploit the advantages of the method in practical scenarios.
Let $L$ denote the number of levels in the circuit, and $\bm{c}_{1},\dots,\bm{c}_{2^{L-1}}$ the initial fresh ciphertexts, generated by encrypting $2^{L-1}$ independent and randomly generated messages, using the same key. The reference circuit for our analysis is depicted in Fig. 2.
We consider a model in which, at each level, homomorphic multiplication of pairs of ciphertexts is carried out. The key difference is that at the beginning of each level $l\in\{1,\dots,L-1\}$ the input ciphertexts are switched to a smaller modulus, to maintain the error almost constant throughout the process. In particular, following the previously defined notion, let $q_{l}$ be the ciphertexts moduli of each level, defined as
$q_{l}=\prod_{j=0}^{l}p_{j},$
where $p_{j}$ are primes such that $\gcd(p_{i},p_{j})=1$ for $i\neq j$.
The process represented by the circuit can be resumed as follows:
  • At level 0, all the fresh ciphertexts are defined in $\mathcal{R}_{q_{L-1}}\times\mathcal{R}_{q_{L-1}}$;
  • At level 1, modulo switch to $q_{L-2}$ is applied to each of the initial fresh ciphertexts. After that, the first homomorphic multiplication is performed in $\mathcal{R}_{q_{L-2}}$, yielding $2^{L-2}$ resulting ciphertexts in $\mathcal{R}_{q_{L-2}}^{3}$. Finally, these ciphertexts are relinearized to obtain equivalent ones in $\mathcal{R}_{q_{L-2}}^{2}$, which constitutes the input of the next level;
  • At level $2$, each input ciphertext undergoes a modulo switch from $q_{L-2}$ to $q_{L-3}$, which is again followed by the homomorphic multiplication of pairs of them. To conclude, a relinearization step is carried out;
  • This process is repeated until level $L-1$, where a final ciphertext, which constitutes the output of the circuit, is returned;
It should be noted that, in the circuit represented in Fig. 2, the relinearization step is not explicated. In fact, as for the case of the circuit without modulus switching, we decided to omit its contribution from our error analysis, since it is negligible compared to the impact of modulus switch and multiplication.
Before discussing the results, it should be noted that, in contrast to the case without the modulus switching, in this scenario, $L$ ciphertext moduli are required for the operation of the circuit. Therefore, the selection of parameters will consist in choosing all the $q_{l}$ values, i.e., all the primes $p_{j}$, such that the correct functioning of the scheme is ensured while still choosing sufficiently small moduli in order to improve efficiency, which is crucial for practical usage of BGV.
For better understanding, we will divide our analysis according to the level $l$ considered. Specifically, an in-depth explanation of the noise estimates for the first three levels will be provided. Then, these results will be generalized for a fixed level $l$, thus delineating our bound.
Level 0
Let $\bm{c}\in\mathcal{R}_{q_{L-1}}\times\mathcal{R}_{q_{L-1}}$ be a fresh ciphertext, obtained encrypting a random message $m\in\mathcal{R}_{t}$.
Let $\nu_{\mathsf{clean}}=a_{0}+a_{1}s$ be the critical quantity associated with $\bm{c}$.
As for the circuits without modulo switching, it is possible to assume that
$\mathsf{Var}(a_{0}|_{i})\approx\mathsf{Var}((a_{1}s)|_{i})\approx\frac{V_{0}}{2},$ (8)
where the notation $V_{0}$ is used to denote the variance of the coefficients of the critical quantity $\nu_{\mathsf{clean}}$. In general, the notation $V_{l}$ will be used to represent the variance of the coefficients of the critical quantity at the end of level $l$. It is significant to observe that the assumption in (8) applies to all initial $2^{L-1}$ ciphertexts. In fact, the estimate of $V_{l}$ is the same for all ciphertexts belonging to the same level.
Level 1
As mentioned above, at the beginning of level $1$, all ciphertexts are subject to a modulus switch from $q_{L-1}$ to $q_{L-2}$.
Therefore, the error and the modulus of $\bm{c}$ are rescaled by a factor of ${q_{L-2}}/{q_{L-1}}={1}/{p_{L-1}}$. Thus, the critical quantity associated to the resulting ciphertext is given by
$\nu_{\mathsf{modswitch}}=\frac{\nu_{\mathsf{clean}}+\nu_{\mathsf{ms}}(p_{L-1})}{p_{L-1}},$ (9)
where $\nu_{\mathsf{ms}}(p_{L-1})$ is defined as
$\nu_{\mathsf{ms}}(p_{L-1})=\delta_{0}+\delta_{1}s\hskip 28.45274pt\delta_{i}=t[-c_{i}t^{-1}]_{p_{L-1}}$
In particular, for the following analysis we will use the notation
$\frac{\nu_{\mathsf{ms}}(p_{L-1})}{p_{L-1}}=a_{0}^{\mathsf{ms}}+a_{1}^{\mathsf{ms}}s,$
with
$\begin{cases}a_{0}^{\mathsf{ms}}=\frac{\delta_{0}}{p_{L-1}}\\ a_{1}^{\mathsf{ms}}=\frac{\delta_{1}}{p_{L-1}}\end{cases}$
Moreover, we will denote with $V_{\mathsf{ms}}$ the following variance
$V_{\mathsf{ms}}=\mathsf{Var}\left(\frac{\nu_{\mathsf{ms}}(p_{L-1})}{p_{L-1}}\middle|_{i}\right).$
For our analysis, it is possible to assume that $V_{\mathsf{ms}}\approx\mathsf{Var}(a_{1}^{\mathsf{ms}}s|_{i})$. This is a consequence of the fact that, observing the distributions of the coefficients in $a_{0}^{\mathsf{ms}}$ and $a_{1}^{\mathsf{ms}}$, it is evident that $\mathsf{Var}(a_{0}^{\mathsf{ms}}|_{i})\ll\mathsf{Var}(a_{1}^{\mathsf{ms}}s|_{i})$. In fact, noting that $c_{i}\leftarrow\mathcal{U}_{q_{L-1}}$, it is possible to estimate the variance of $a_{0}^{\mathsf{ms}}|_{i}$ and $a_{1}^{\mathsf{ms}}|_{i}$ as
$\begin{cases}\mathsf{Var}(a_{0}^{\mathsf{ms}}|_{i})=\mathsf{Var}(a_{1}^{\mathsf{ms}}|_{i})=t^{2}/12\\ \mathsf{Var}(a_{1}^{\mathsf{ms}}s|_{i})={t^{2}nV_{s}}/{12},\end{cases}$
where $n$ is typically in $\small\{2^{12},2^{13},2^{14},2^{15}\}$. With this in mind, we can proceed with the analysis of the noise growth related to the multiplication of two ciphertexts, whose associated critical quantity is as in (9).
We start by observing that the critical quantity in (9) can be written as
$\nu_{\mathsf{modswitch}}=\hat{a_{0}}+\hat{a_{1}}s=\frac{a_{0}}{p_{L-1}}+a_{0}^{\mathsf{ms}}+\left(\frac{a_{1}}{p_{L-1}}+a_{1}^{\mathsf{ms}}\right)s.$
Moreover, it is possible to further simplify the notation by using the fact that $\mathbb{E}[a_{0}^{\mathsf{ms}}]=0$ and its variance is negligible. Therefore, we rewrite this value as
$\nu_{\mathsf{modswitch}}=\hat{a_{0}}+\hat{a_{1}}s=\frac{a_{0}}{p_{L-1}}+\left(\frac{a_{1}}{p_{L-1}}+a_{1}^{\mathsf{ms}}\right)s.$
Then, it can be derived that
$\begin{cases}\mathsf{Var}(\hat{a_{0}}|_{i})\approx\frac{V_{0}}{2p_{L-1}^{2}}\\ \mathsf{Var}(\hat{a_{1}}s|_{i})\approx\frac{V_{0}}{2p_{L-1}^{2}}+V_{\mathsf{ms}}\end{cases}$
since
$\mathsf{Var}\left(\frac{a_{0}}{p_{L-1}}\middle|_{i}\right)=\mathsf{Var}\left(\frac{a_{1}}{p_{L-1}}s\middle|_{i}\right)=\frac{V_{0}}{2p_{L-1}^{2}}.$
In order to estimate the variance of the critical quantity resulting from multiplication, we will rely on Theorem 4.1. In particular, in this analysis, we will try to avoid explicitly writing the terms $b_{\mu}(\iota)$ involved in the expression of a generic $a_{\iota}=\sum_{\mu=0}^{K}b_{\mu}(\iota)e^{\mu}$, to avoid the notation becoming cumbersome. However, it is crucial to keep in mind the highest power of $e$, namely $K$, present in the terms $a_{\iota}$ involved in the multiplications, as the correction of the function $F$ depends on this value. Precisely, for fresh ciphertexts at level 00, it can be seen that for $a_{0}$ the associated $K$ is $1$, while for $a_{1}$ we have $K=0$. Instead, for the terms $a_{0}^{\mathsf{ms}},a_{1}^{\mathsf{ms}}$ the case is quite different. In fact, these values can be assumed to be randomly distributed over $\mathcal{R}_{t}$, i.e. $a_{0}^{\mathsf{ms}},a_{1}^{\mathsf{ms}}\leftarrow\mathcal{U}_{t}$ for every level of the circuit. This makes the analysis slightly more complicated, since, in order to derive tight estimates, it will be necessary to distinguish the contribution of ${a_{0},a_{1}}$ from the one of $a_{0}^{\mathsf{ms}},a_{1}^{\mathsf{ms}}$.
Now, assume that $\bm{c},\bm{c}^{\prime}\in\mathcal{R}_{q_{L-2}}\times\mathcal{R}_{q_{L-2}}$ are two ciphertexts at level 1, after modulus switching has been carried out. Their associated critical quantity is given, respectively, by
$\begin{cases}\nu_{\mathsf{modswitch}}=\frac{1}{p_{L-1}}a_{0}+\left(\frac{1}{p_{L-1}}a_{1}+a_{1}^{\mathsf{ms}}\right)s\\ \nu^{\prime}_{\mathsf{modswitch}}=\frac{1}{p_{L-1}}a^{\prime}_{0}+\left(\frac{1}{p_{L-1}}a^{\prime}_{1}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s\\ \end{cases}$
Therefore, the critical quantity obtained after their multiplication is of the form
$\nu_{\mathsf{mul}}^{(1)}=a_{0}^{\mathsf{mul}}+a_{1}^{\mathsf{mul}}s+a_{2}^{\mathsf{mul}}s^{2},$
namely,
$\nu_{\mathsf{mul}}^{(1)}$ $=\frac{1}{p_{L-1}^{2}}a_{0}a^{\prime}_{0}+\frac{1}{p^{2}_{L-1}}\left(a^{\prime}_{0}{a_{1}}+a_{0}{a^{\prime}_{1}}\right)s+\frac{1}{p_{L-1}}(a^{\prime}_{0}a_{1}^{\mathsf{ms}}+a_{0}{a^{\prime}_{1}}^{\mathsf{ms}})s\hskip 4.26773pt$
$\hskip 11.38092pt+\frac{1}{p^{2}_{L-1}}a_{1}a^{\prime}_{1}s^{2}+\frac{1}{p_{L-1}}(a_{1}{a^{\prime}_{1}}^{\mathsf{ms}}+a^{\prime}_{1}{a_{1}}^{\mathsf{ms}})s^{2}+{a^{\prime}_{1}}^{\mathsf{ms}}{a_{1}}^{\mathsf{ms}}s^{2}.$
So, we are now able to provide an estimate of the variance $V_{1}=\mathsf{Var}(\nu_{\mathsf{mul}}|_{i})$, applying Theorem 4.1, as follows
$V_{1}$ $\leq\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a^{\prime}_{0}|_{i})F_{e}(1,1)$
$+\frac{n}{p_{L-1}^{4}}V(a^{\prime}_{0}|_{i})V(a_{1}s|_{i})+\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a^{\prime}_{1}s|_{i})$
$+\frac{n}{p_{L-1}^{2}}V(a^{\prime}_{0}|_{i})V(a_{1}^{\mathsf{ms}}s|_{i})+\frac{n}{p_{L-1}^{2}}V(a_{0}|_{i})V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})$
$+\frac{n}{p_{L-1}^{4}}V(a_{1}s|_{i})V(a^{\prime}_{1}s|_{i})F_{s}(1,1)+\frac{n}{p_{L-1}^{2}}V(a_{1}s|_{i})V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})F_{s}(1,1)$
$+\frac{n}{p_{L-1}^{2}}V(a^{\prime}_{1}s|_{i})V({a_{1}}^{\mathsf{ms}}s|_{i})F_{s}(1,1)+{n}V({a_{1}}^{\mathsf{ms}}s|_{i})V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})F_{s}(1,1).$
which, recalling that $\operatorname{Var}(a_{\iota}|_{i})=\operatorname{Var}(a^{\prime}_{\iota}|_{i})$ and that $V({a_{1}}^{\mathsf{ms}}s|_{i})=V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})=V_{\mathsf{ms}}$, can be rewritten as
$V_{1}$ $\leq\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})^{2}F_{e}(1,1)+$
$+\frac{2n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a_{1}s|_{i})+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}V(a_{0}|_{i})$
$+\frac{n}{p_{L-1}^{4}}V(a_{1}s|_{i})^{2}F_{s}(1,1)+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}V(a_{1}s|_{i})+nF_{s}(1,1)V_{\mathsf{ms}}^{2}.$
By substituting to $V(a_{0}|_{i})$,$V(a_{1}s|_{i})$ the value $\frac{V_{0}}{2}$, our bound on $V_{1}$ is derived. To conclude, it is possible to observe that the following estimate holds
$\begin{cases}V(a_{0}^{\mathsf{mul}}|_{i})\approx\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})^{2}F_{e}(1,1)\\ V(a_{1}^{\mathsf{mul}}s|_{i})\approx\frac{2n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a_{1}s|_{i})+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}V(a_{0}|_{i})\\ V(a_{2}^{\mathsf{mul}}s^{2}|_{i})\approx\frac{n}{p_{L-1}^{4}}V(a_{1}s|_{i})^{2}F_{s}(1,1)+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}V(a_{1}s|_{i})+nF_{s}(1,1)V_{\mathsf{ms}}^{2}\end{cases}$ (10)
These values are needed in order to bound the variance $V_{2}$ at the end of the next level.
Level 2
The first computation at level 2 is the modulo switch from $q_{L-2}$ to $q_{L-3}$.
It should be noted that, for our theoretical analysis, we chose to avoid considering the key switching step. Therefore, the ciphertexts which will be subject to modulo switch have critical quantity of the form
$\nu^{(1)}=a_{0}+a_{1}s+a_{2}s^{2},$
with $a_{0},a_{1},a_{2}$ having variances as in (10).
After the modulo switch, the critical quantity associated to the resulting ciphertexts will be as follows
$\nu_{\mathsf{modswitch}}=\frac{\nu^{(1)}+\nu_{\mathsf{ms}}(p_{L-2})}{p_{L-2}}.$
Therefore, it is possible to express this quantity as
$\nu_{\mathsf{modswitch}}=\frac{a_{0}}{p_{L-2}}+\left(\frac{a_{1}}{p_{L-2}}+a_{1}^{\mathsf{ms}}\right)s+\frac{a_{2}}{p_{L-2}}s^{2}.$
For this level, the maximum power of $e$ contained in each $a_{i}$ has changed.
In fact, $a_{0}$ contains a term multiplied by $e^{2}$, $a_{1}$ a term multiplied by $e$, while $a_{2}$ does not contain any power of $e$ as for the term $a_{1}^{\mathsf{ms}}$.
This distinction is crucial in order to correctly apply Theorem 4.1.
In light of this, given two ciphertexts with critical quantities, respectively,
$\begin{cases}\nu_{\mathsf{modswitch}}=\frac{a_{0}}{p_{L-2}}+\left(\frac{a_{1}}{p_{L-2}}+{a_{1}}^{\mathsf{ms}}\right)s+\frac{a_{2}}{p_{L-2}}s^{2}\\ \nu^{\prime}_{\mathsf{modswitch}}=\frac{a^{\prime}_{0}}{p_{L-2}}+\left(\frac{a^{\prime}_{1}}{p_{L-2}}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s+\frac{a^{\prime}_{2}}{p_{L-2}}s^{2}\end{cases}$
the resulting $\nu_{\mathsf{mul}}^{(2)}$ that follows by their multiplication will be of the form
$\nu_{\mathsf{mul}}^{(2)}$ $=\nu_{\mathsf{modswitch}}\cdot\nu^{\prime}_{\mathsf{modswitch}}=a_{0}^{\mathsf{mul}}+a_{1}^{\mathsf{mul}}s+a_{2}^{\mathsf{mul}}s^{2}+a_{3}^{\mathsf{mul}}s^{3}+a_{4}^{\mathsf{mul}}s^{4},$
which can be written as
$\nu_{\mathsf{mul}}^{(2)}$ $=\frac{a_{0}a^{\prime}_{0}}{p^{2}_{L-2}}+\frac{a^{\prime}_{0}}{p_{L-2}}\left(\frac{a_{1}}{p_{L-2}}+{a_{1}}^{\mathsf{ms}}\right)s+\frac{a_{0}}{p_{L-2}}\left(\frac{a^{\prime}_{1}}{p_{L-2}}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s$
$\hskip 11.38092pt+\left(\frac{a_{1}}{p_{L-2}}+{a_{1}}^{\mathsf{ms}}\right)\left(\frac{a^{\prime}_{1}}{p_{L-2}}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s^{2}+\frac{a_{0}a^{\prime}_{2}}{p^{2}_{L-2}}s^{2}+\frac{a^{\prime}_{0}a_{2}}{p^{2}_{L-2}}s^{2}$
$\hskip 11.38092pt+\frac{a^{\prime}_{2}}{p_{L-2}}\left(\frac{a_{1}}{p_{L-2}}+{a_{1}}^{\mathsf{ms}}\right)s^{3}+\frac{a_{2}}{p_{L-2}}\left(\frac{a^{\prime}_{1}}{p_{L-2}}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s^{3}+\frac{a_{2}a^{\prime}_{2}}{p^{2}_{L-2}}s^{4}.$
A bound over $V_{2}$ can then be deduced, by applying Theorem 4.1 and recalling that $\mathsf{Var}(a_{\iota}|_{i})=\mathsf{Var}(a^{\prime}_{\iota}|_{i})$, resulting in
$V_{2}\leq$ $\,\frac{n}{p_{L-2}^{4}}\mathsf{Var}(a_{0}|_{i})^{2}F_{e}(2,2)+\frac{2n}{p_{L-2}^{4}}\mathsf{Var}(a_{0}|_{i})\mathsf{Var}(a_{1}s|_{i})F_{e}(1,2)$
$+\frac{n}{p_{L-2}^{4}}\mathsf{Var}(a_{1}s|_{i})^{2}F_{s}(1,1)F_{e}(1,1)+\frac{2n}{p_{L-2}^{4}}\mathsf{Var}(a_{0}|_{i})\mathsf{Var}(a_{2}s^{2}|_{i})$
$+\frac{2n}{p_{L-2}^{4}}\mathsf{Var}(a_{1}s|_{i})\mathsf{Var}(a_{2}s^{2}|_{i})F_{s}(1,2)+\frac{n}{p_{L-2}^{4}}\mathsf{Var}(a_{2}s^{2}|_{i})^{2}F_{s}(2,2)$
$+\frac{2n}{p_{L-2}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{0}|_{i})+\frac{2n}{p_{L-2}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{1}s|_{i})F_{s}(1,1)$
$+\frac{2n}{p_{L-2}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{2}s^{2}|_{i})F_{s}(1,2)+{n}V_{\mathsf{ms}}^{2}F_{s}(1,1),$
where the values for $V(a_{\iota}s^{\iota}|_{i})$ follows by (10). Moreover, similar to level $1$, it is possible to deduce the contribution, in terms of variance, of each $a_{\iota}^{\mathsf{mul}}$ of $\nu_{\mathsf{mul}}^{(2)}$, which will then be needed for the successive levels.
Level ${l}$
For a generic level $l$, with $l\in\{1,\dots,L-1\}$, the same reasoning applied in the first levels can be reiterated. Assume that two ciphertexts with critical quantity of the form
$\nu^{(l-1)}=\sum_{\iota=0}^{2^{l-1}}a_{\iota}s^{\iota},$
are given as input. Let $V_{l}$ be the variance of the coefficients of $\nu^{(l)}$, which is the critical quantity obtained after modulo switch and multiplication.
Following the notation adopted in the precedent levels, this quantity can be indicated as
$\nu^{(l)}=\sum_{\iota=0}^{2^{l}}a_{\iota}^{\mathsf{mul}}s^{\iota}.$
Then, $V_{l}$ is bounded by the sum of
$V_{l}$ $\leq\frac{n}{p_{L-l}^{4}}\sum_{\iota=0}^{2^{l}}\sum_{k+\mu=\iota}\mathsf{Var}(a_{k}s^{k}|_{i})\mathsf{Var}(a_{\mu}s^{\mu}|_{i}){F_{s}(\iota,\mu)}{F_{e}(2^{l-1}-\iota,2^{l-1}-\mu)}$
$\quad+\frac{2n}{p_{L-l}^{2}}\sum_{\iota=0}^{2^{l-1}}V_{\mathsf{ms}}\mathsf{Var}(a_{\iota}s^{\iota}|_{i}){F_{s}(1,\iota)}+nV_{\mathsf{ms}}^{2}{F_{s}(1,1)}$ (11)
Moreover, each term of $\nu^{(l)}$ has variance
$\begin{cases}\mathsf{Var}(a_{0}^{\mathsf{mul}}|_{i})=\frac{n}{p_{L-l}^{4}}\sum_{k+\mu=0}\mathsf{Var}(a_{k}s^{k}|_{i})\mathsf{Var}(a_{\mu}s^{\mu}|_{i}){F_{s}(k,\mu)}{F_{e}(2^{l-1}-k,2^{l-1}-\mu)}\\ \mathsf{Var}(a_{1}^{\mathsf{mul}}s|_{i})=\frac{n}{p_{L-l}^{4}}\sum_{k+\mu=1}\mathsf{Var}(a_{k}s^{k}|_{i})\mathsf{Var}(a_{\mu}s^{\mu}|_{i}){F_{s}(k,\mu)}{F_{e}(2^{l-1}-k,2^{l-1}-\mu)}\\ \hskip 65.44142pt+\frac{2n}{p_{L-l}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{0}|_{i}){F_{s}(1,0)}\\ \mathsf{Var}(a_{2}^{\mathsf{mul}}s^{2}|_{i})=\frac{n}{p_{L-l}^{4}}\sum_{k+\mu=2}\mathsf{Var}(a_{k}s^{k}|_{i})\mathsf{Var}(a_{\mu}s^{\mu}|_{i}){F_{s}(k,\mu)}{F_{e}(2^{l-1}-k,2^{l-1}-\mu)}\\ \hskip 65.44142pt+\frac{2n}{p_{L-l}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{1}s|_{i}){F_{s}(1,1)}+nF_{s}(1,1)V_{\mathsf{ms}}^{2}\\ \mathsf{Var}(a_{\iota}^{\mathsf{mul}}s^{\iota}|_{i})=\frac{n}{p_{L-l}^{4}}\sum_{k+\mu=\iota}\mathsf{Var}(a_{k}s^{k}|_{i})\mathsf{Var}(a_{\mu}s^{\mu}|_{i}){F_{s}(k,\mu)}{F_{e}(2^{l-1}-k,2^{l-1}-\mu)}\\ \hskip 65.44142pt+\frac{2n}{p_{L-l}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{\iota-1}s^{\iota-1}|_{i}){F_{s}(1,\iota-1)}\mathbb{I}_{\{0,\dots,2^{l-1}+1\}}(\iota)\end{cases}$
where $\mathbb{I}_{\{0,\dots,2^{l-1}+1\}}(\iota)$ is the indicator function, i.e.,
$\mathbb{I}_{\{0,\dots,2^{l-1}+1\}}(\iota)=\begin{cases}1,&\text{if }\iota\in\{0,\dots,2^{l-1}+1\}\\ 0,&\text{if }\iota\notin\{0,\dots,2^{l-1}+1\}\end{cases}$
6 Conclusions
In this work, we focused on improving the current average-case approach for tracking the error and its growth under homomorphic operations. In particular, we demonstrated that the primary reason for the underestimation of noise in existing average-case methods lies in the fact that the coefficients of the error polynomial are treated as independent. However, to obtain accurate bounds that never underestimate the actual values observed in practice, it is crucial to account for the dependencies among these coefficients.
Accurate noise estimation plays a key role in the proper selection of scheme parameters, especially the ciphertext modulus $q$, to ensure correctness, security, and improved efficiency. To evaluate the effectiveness of our analysis, we applied our results to specific homomorphic circuits, comparing our estimations with both experimental data and those obtained from the current average-case approach. The circuits were implemented using the MAGMA language.
Our findings show that the proposed method effectively overcomes the limitations of existing average-case techniques, particularly their tendency to underestimate noise variance. Furthermore, in the context of initial parameter selection, the high accuracy of our estimations, consistently close to experimental results, represents a significant improvement over the worst-case methods that currently dominate state-of-the-art parameter generation. This leads to tighter bounds and thus enables more efficient configurations of the scheme, which is especially critical for promoting the practical adoption of homomorphic encryption, and in particular of the BGV scheme, in real-world applications.
Future work will include extending our approach to other cryptographic schemes, such as CKKS, as well as studying scenarios in which ciphertexts are generated in a dependent manner. We are currently developing a comprehensive framework for modulus selection that is fully aligned with the new approach introduced in this paper.
Acknowledgment
The third author was partially supported by the Italian Ministry of University and Research in the framework of the Call for Proposals for scrolling of final rankings of the PRIN 2022 call - Protocol no. 2022RFAZCJ and by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.
Appendix 0.A Proof of Lemma 1
In order to prove this lemma, we will first demonstrate that these properties hold for the critical quantity of a fresh ciphertext $\nu_{\mathsf{clean}}$.
Then, we will show that any operation involved in the BGV circuit does not affect these properties.
Fresh ciphertexts
For $\nu_{\mathsf{clean}}$, the coefficients $b_{\iota}(\mu)$ are defined as
$\begin{cases}b_{0}(0)=m+te_{0}\\ b_{1}(0)=tu\end{cases}\begin{cases}b_{0}(1)=te_{1}\end{cases}$
Therefore, the first property follows immediately from the independence of $b_{\mu_{1}}(\iota_{1})|_{j_{1}},b_{\mu_{2}}(\iota_{2})|_{j_{2}}$ when $\mu_{1}\neq\mu_{2}\text{ or }j_{1}\neq j_{2}$.
As for the second property, it holds since
  • $\mathbb{E}[b_{0}(0)|_{i}]=\mathbb{E}[m|_{i}]+t\mathbb{E}[e_{0}|_{i}]=0$, due to the linearity of the expected value and the distributions considered, i.e. $m\leftarrow\mathcal{U}_{t},e_{0}\leftarrow\mathcal{DG}_{q}(\sigma^{2})$;
  • $\mathbb{E}[b_{1}(0)|_{i}]=t\mathbb{E}[u|_{i}]=0$, as $u\leftarrow\chi_{s}$;
  • $\mathbb{E}[b_{0}(1)|_{i}]=t\mathbb{E}[e_{1}|_{i}]=0$, as $e_{1}\leftarrow\mathcal{DG}_{q}(\sigma^{2})$;
We will therefore show that the remaining homomorphic operations do not alter these properties.
Let $\nu=\sum_{\iota_{1}}\sum_{\mu_{1}}b_{\mu_{1}}(\iota_{1})e^{\mu_{1}}s^{\iota_{1}}$ , $\nu^{\prime}=\sum_{\iota_{2}}\sum_{\mu_{2}}b^{\prime}_{\mu_{2}}(\iota_{2})e^{\mu_{2}}s^{\iota_{2}}$ be the respective critical quantities of two generic ciphertexts, for which the properties stated above are assumed to hold.
Addition of two ciphertexts
The critical quantity after the addition of the two BGV ciphertexts is given by
$\nu_{\text{add}}=\nu+\nu^{\prime}=\sum_{\iota}\sum_{\mu}b^{\text{add}}_{\mu}(\iota)e^{\mu}s^{\iota},$
where $b^{\text{add}}_{\mu}(\iota)=b_{\mu}(\iota)+b^{\prime}_{\mu}(\iota)$.
Therefore, if $\mathbb{E}[b_{\mu}(\iota)|_{i}]=\mathbb{E}[b^{\prime}_{\mu}(\iota)|_{i}]=0$ then
$\mathbb{E}[b^{\text{add}}_{\mu}(\iota)|_{i}]=0,$
according to the linearity of the expected value.
Moreover, using the bilinearity of the covariance, we have that, for $\mu_{1}\neq\mu_{2}\text{ or }j_{1}\neq j_{2}$
$\mathsf{Cov}(b_{\mu_{1}}^{\text{add}}(\iota_{1})|_{j_{1}},b_{\mu_{2}}^{\text{add}}(\iota_{2})|_{j_{2}})=0,$
since
$\mathsf{Cov}(b_{\mu_{1}}(\iota_{1})+b^{\prime}_{\mu_{1}}(\iota_{1})|_{j_{1}},b_{\mu_{2}}(\iota_{2})+b^{\prime}_{\mu_{2}}(\iota_{2})|_{j_{2}})$ $=\mathsf{Cov}(b_{\mu_{1}}(\iota_{1})|_{j_{1}},b_{\mu_{2}}(\iota_{2})|_{j_{2}})$
$\quad+\mathsf{Cov}(b_{\mu_{1}}(\iota_{1})|_{j_{1}},b^{\prime}_{\mu_{2}}(\iota_{2})|_{j_{2}})$
$\quad+\mathsf{Cov}(b^{\prime}_{\mu_{1}}(\iota_{1})|_{j_{1}},b_{\mu_{2}}(\iota_{2})|_{j_{2}})$
$\quad+\mathsf{Cov}(b^{\prime}_{\mu_{1}}(\iota_{1})|_{j_{1}},b^{\prime}_{\mu_{2}}(\iota_{2})|_{j_{2}}),$
where all the summands vanish because:
  • $\mathsf{Cov}(b_{\mu_{1}}(\iota_{1})|_{j_{1}},b_{\mu_{2}}(\iota_{2})|_{j_{2}})=\mathsf{Cov}(b^{\prime}_{\mu_{1}}(\iota_{1})|_{j_{1}},b^{\prime}_{\mu_{2}}(\iota_{2})|_{j_{2}})=0$ holds by assumption for $\mu_{1}\neq\mu_{2}$ or $j_{1}\neq j_{2}$;
  • $\mathsf{Cov}(b_{\mu_{1}}(\iota_{1})|_{j_{1}},b^{\prime}_{\mu_{2}}(\iota_{2})|_{j_{2}})=\mathsf{Cov}(b^{\prime}_{\mu_{1}}(\iota_{1})|_{j_{1}},b_{\mu_{2}}(\iota_{2})|_{j_{2}})=0$ since $b_{\mu_{1}}(\iota_{1})$ and $b^{\prime}_{\mu_{2}}(\iota_{2})$ are independent $\forall\mu_{1},\mu_{2}$;
Multiplication by a constant
Given a ciphertext with $\nu=\sum_{\iota}\sum_{\mu}b_{\mu}(\iota)e^{\mu}s^{\iota}$ and a constant $\alpha$, the critical quantity obtained after their homomorphic multiplication, according to the BGV scheme, is as follows
$\nu_{\text{const}}=\alpha\nu=\alpha\sum_{\iota}\sum_{\mu}b_{\mu}(\iota)e^{\mu}s^{\iota}=\sum_{\iota}\sum_{\mu}\alpha b_{\mu}(\iota)e^{\mu}s^{\iota}.$
Therefore, we can define $b^{\text{const}}_{\mu}(\iota)=\alpha b_{\mu}(\iota)$ from which, according to the linearity of the expected value
$\mathbb{E}[b^{\text{const}}_{\mu}(\iota)|_{i}]=\mathbb{E}[\alpha b_{\mu}(\iota)|_{i}]=\alpha\mathbb{E}[b_{\mu}(\iota)|_{i}]=0,$
which proves property $b)$.
Property $a)$ follows directly by assumption from the bilinearity of the covariance
$\mathsf{Cov}(b_{\mu_{1}}^{\text{const}}(\iota_{1})|_{j_{1}},b_{\mu_{2}}^{\text{const}}(\iota_{2})|_{j_{2}})$ $=\mathsf{Cov}(\alpha b_{\mu_{1}}(\iota_{1})|_{j_{1}},\alpha b_{\mu_{2}}(\iota_{2})|_{j_{2}})=$
$=\alpha^{2}\mathsf{Cov}(b_{\mu_{1}}(\iota_{1})|_{j_{1}},b_{\mu_{2}}(\iota_{2})|_{j_{2}})=0.$
Multiplication of two ciphertexts
The critical quantity arising from the multiplication of two ciphertexts, whose associated noise is defined as above, can be expressed as
$\nu_{\mathsf{mul}}=\nu\cdot\nu^{\prime}=\sum_{\iota}a^{\mathsf{mul}}_{\iota}s^{\iota}=\sum_{\iota}\sum_{\mu}b^{\mathsf{mul}}_{\mu}({\iota})e^{\mu}s^{\iota},$
where $a^{\mathsf{mul}}_{\iota}=\sum_{\iota_{1}+\iota_{2}=\iota}a_{\iota_{1}}a^{\prime}_{\iota_{2}}$.
Moreover
$\begin{cases}a_{\iota_{1}}=\sum_{\mu_{1}}b_{\mu_{1}}({\iota_{1}})e^{\mu_{1}}\\ a^{\prime}_{\iota_{2}}=\sum_{\mu_{2}}b^{\prime}_{\mu_{2}}({\iota_{2}})e^{\mu_{2}}\end{cases}$
From which it follows that
$b^{\mathsf{mul}}_{\mu}({\iota})=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}b_{\mu_{1}}({\iota_{1}})b^{\prime}_{\mu_{2}}({\iota_{2}}).$
From the independence of $b_{\mu_{1}}(\iota_{1})$ and $b^{\prime}_{\mu_{2}}(\iota_{2})$ $\forall\mu_{1},\mu_{2},\iota_{1},\iota_{2}$, and for the linearity of the expected value, one can deduce that
$\mathbb{E}[b^{\mathsf{mul}}_{\mu}({\iota})]$ $=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\mathbb{E}[b_{\mu_{1}}({\iota_{1}})b^{\prime}_{\mu_{2}}({\iota_{2}})]$
$=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\mathbb{E}[b_{\mu_{1}}({\iota_{1}})]\mathbb{E}[b^{\prime}_{\mu_{2}}({\iota_{2}})]=0,$
which easily proves property $b)$.
The expression of the covariance $\mathsf{Cov}(b^{\mathsf{mul}}_{\mu_{1}}({\iota_{1}})|_{i_{1}},b^{\mathsf{mul}}_{\mu_{2}}({\iota_{2}})|_{i_{2}})$ can be reduced, using its bilinearity, to a sum of terms of the form
$\mathsf{Cov}(b_{\mu_{1}}({\iota_{1}})|_{l_{1}}b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i_{1}-l_{1}},b_{\mu_{3}}({\iota_{3}})|_{l_{2}}b^{\prime}_{\mu_{4}}({\iota_{4}})|_{i_{2}-l_{2}}),$
which are all zero, using the property of the covariance stated below.
Property 1
Let $X_{1},X_{2},X_{3},X_{4}$ be some fixed random variables.
If $X_{2},X_{4}$ are independent with respect to $X_{1},X_{3}$, $\mathsf{Cov}(X_{2},X_{4})=0$ and $\mathbb{E}[X_{2}]=0$ then
$\mathsf{Cov}(X_{1}\cdot X_{2},X_{3}\cdot X_{4})=0.$
Thus, property $a)$ follows by observing that, for $\mu_{2}\neq\mu_{4}\text{ or }i_{2}\neq i_{4}$:
  • $\mathsf{Cov}(b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i_{1}-l_{1}},b^{\prime}_{\mu_{4}}({\iota_{4}})|_{i_{2}-l_{2}})=0$ e $\mathbb{E}[b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i_{1}-l_{1}}]=0$ based on the hypotheses made;
  • $b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i_{1}-l_{1}},b^{\prime}_{\mu_{4}}({\iota_{4}})|_{i_{2}-l_{2}}$ are independent with respect to $b_{\mu_{1}}({\iota_{1}})|_{l_{1}},b_{\mu_{3}}({\iota_{3}})|_{l_{2}}$;
Therefore, thanks to the property 1, this implies that
$\mathsf{Cov}(b_{\mu_{1}}({\iota_{1}})|_{l_{1}}b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i_{1}-l_{1}},b_{\mu_{3}}({\iota_{3}})|_{l_{2}}b^{\prime}_{\mu_{4}}({\iota_{4}})|_{i_{2}-l_{2}})=0.$
Modulus and Key Switching
Let $\bm{c}=(c_{0},c_{1})$ be a ciphertext in $R_{q_{l}}\times R_{q_{l}}$ and suppose that the modulus switch to $q_{l^{\prime}}$ is applied, in order to reduce the error.
The resulting ciphertext is defined as
$\bm{c^{\prime}}=\frac{q_{l^{\prime}}}{q_{l}}(\bm{c}+\bm{\delta})\mod q_{l^{\prime}},$
where $\bm{\delta}=t[-\bm{c}t^{-1}]_{\frac{q_{l}}{q_{l^{\prime}}}}$.
The critical quantity associated to $\bm{c^{\prime}}$ can be expressed as
$\nu^{\prime}=\frac{q_{l^{\prime}}}{q_{l}}(\nu+\nu_{\mathsf{ms}})\quad\text{ where}\quad\nu_{\mathsf{ms}}=\delta_{0}+\delta_{1}s$
It is possible to observe that the ciphertext components ${c}_{0},c_{1}$ can be thought as randomly distributed over $R_{q_{l}}$, ${c_{0}},c_{1}\leftarrow\mathcal{U}_{q_{l}}$, and therefore the $\delta_{i}$ can be treated as independent polynomials with coefficients chosen randomly over $I={(-\frac{tq_{l}}{2q_{l^{\prime}}},\frac{tq_{l}}{2q_{l^{\prime}}})}$, i.e. $\delta_{0},\delta_{1}\leftarrow\mathcal{U}_{I}$.
Moreover, it should be noted that the values $\delta_{i}$ exclusively influence $b_{0}(0),b_{0}(1)$, and that they have an expected value equal to zero, because of their distributions.
Therefore, referring back to the case of the homomorphic sum, we can deduce that the expected value of $b^{\prime}_{\mu}(\iota)$ for the new ciphertext $\bm{c^{\prime}}$ remains zero, as do the covariances.
In the same way, by reducing the problem to the case of homomorphic addition, it is possible to show that these properties remain valid also after the relinearization process.
We decided not to report all the technical details but to provide only the key underlying idea, as there are multiple relinearization variants and including them would have required too much space. However, all the calculations can be derived in a very straightforward manner by simply adapting the approach in [2].
Appendix 0.B Proof of Lemma 2
In order to prove the statement, we start by writing the term $a_{\iota}s^{\iota}|_{i}$, according to 1, as
$a_{\iota}s^{\iota}|_{i}=\sum_{\mu}(b_{\mu}(\iota)e^{\mu}s^{\iota})|_{i}=\sum_{\mu}\sum_{j=0}^{n-1}\xi(i,j)b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j}.$
Recall that, given two random variables X and Y, the following properties hold:
[…]
$\hskip 68.28644pt-(\mathsf{Cov}(X,Y)+\mathbb{E}[X]\mathbb{E}[Y])^{2}$
[…]
where the second property can be generalized for $k$ random variables $\{X_{i}\}_{i=0}^{k}$ as follows
$\mathsf{Var}\left(\sum_{i=0}^{k}X_{i}\right)=\sum_{i=0}^{k}\mathsf{Var}(X_{i})+\sum_{i_{1}\neq i_{2}}\mathsf{Cov}(X_{i_{1}},X_{i_{2}}).$
Then, it is possible to compute the variance of $a_{\iota}s^{\iota}|_{i}$ as
$\mathsf{Var}(a_{\iota}s^{\iota}|_{i})$ $=\sum_{\mu}\sum_{j=0}^{n-1}\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j})$
$+\sum_{\mu_{1}\neq\mu_{2}\text{ or }j_{1}\neq j_{2}}\xi(i,j_{1})\xi(i,j_{2})\mathsf{Cov}(b_{\mu_{1}}(\iota)e^{\mu_{1}}|_{j_{1}}s^{\iota}|_{i-j_{1}},b_{\mu_{2}}(\iota)e^{\mu_{2}}|_{j_{2}}s^{\iota}|_{i-j_{2}}),$
where the covariances vanishes based on property 1.
Thus, the following equality holds
$\mathsf{Var}(a_{\iota}s^{\iota}|_{i})=\sum_{\mu}\sum_{j=0}^{n-1}\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j}).$ (12)
Moreover, according to property (a.), it follows that
$\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j})$ $=(\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j})+\mathbb{E}[b_{\mu}(\iota)e^{\mu}|_{j}]^{2})(\mathsf{Var}(s^{\iota}|_{i-j})+\mathbb{E}[s^{\iota}|_{i-j}]^{2})$
$\quad+\mathsf{Cov}(b_{\mu}(\iota)e^{\mu}|_{j}^{2},s^{\iota}|_{i-j}^{2})$
$\quad-\left(\mathsf{Cov}(b_{\mu}(\iota)e^{\mu}|_{j},s^{\iota}|_{i-j})+\mathbb{E}[b_{\mu}(\iota)e^{\mu}|_{j}]\mathbb{E}[s^{\iota}|_{i-j}]\right)^{2}.$
At this point, it should be noted that
  • $\mathbb{E}[b_{\mu}(\iota)e^{\mu}|_{j}]=0$ according to lemma 1;
  • $\mathsf{Cov}(b_{\mu}(\iota)e^{\mu}|_{j},s^{\iota}|_{i-j})=\mathsf{Cov}(b_{\mu}(\iota)e^{\mu}|_{j}^{2},s^{\iota}|_{i-j}^{2})=0$ as $b_{\mu}(\iota)e^{\mu}|_{j},s^{\iota}|_{i-j}$ are independent;
This results in
$\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j})=\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j})(\mathsf{Var}(s^{\iota}|_{i-j})+\mathbb{E}[s^{\iota}|_{i-j}]^{2}).$
In addition, for a random variable $X$, it holds that
$\mathsf{Var}(X)=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}.$
Therefore,
$\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j})=\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j})\mathbb{E}[s^{\iota}|_{i-j}^{2}].$ (13)
Finally, the same reasoning can be applied in order to derive
$\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j})=\sum_{k=0}^{n-1}\mathsf{Var}(b_{\mu}(\iota)|_{k})\mathbb{E}[e^{\mu}|_{j-k}^{2}].$ (14)
In fact, using (1) and property (b.), it follows that
$\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j})$ $=\mathsf{Var}\left(\sum_{k=0}^{n-1}b_{\mu}(\iota)|_{k}e^{\mu}|_{j-k}\right)$
$=\sum_{k=0}^{n-1}\mathsf{Var}(b_{\mu}(\iota)|_{k}e^{\mu}|_{j-k})$
$\hskip 11.38092pt+\sum_{k_{1}\neq k_{2}}\xi(j,k_{1})\xi(j,k_{2})\mathsf{Cov}(b_{\mu}(\iota)|_{k_{1}}e^{\mu}|_{j-k_{1}},b_{\mu}(\iota)|_{k_{2}}e^{\mu}|_{j-k_{2}}),$
where the covariances are null thanks to property 1.
Moreover, according to property (a.), it follows that
$\mathsf{Var}(b_{\mu}(\iota)|_{k}e^{\mu}|_{j-k})$ $=(\mathsf{Var}(b_{\mu}(\iota)|_{k})+\mathbb{E}[b_{\mu}(\iota)|_{k}]^{2})(\mathsf{Var}(e^{\mu}|_{j-k})+\mathbb{E}[e^{\mu}|_{j-k}]^{2})$
$\quad+\mathsf{Cov}(b_{\mu}(\iota)|_{k}^{2},e^{\mu}|_{j-k}^{2})$
$\quad-(\mathsf{Cov}(b_{\mu}(\iota)|_{k},e^{\mu}|_{j-k})+\mathbb{E}[b_{\mu}(\iota)|_{k}]\mathbb{E}[e^{\mu}|_{j-k}])^{2}.$
Thus, (13) is proven observing that $\mathbb{E}[b_{\mu}(\iota)|_{k}]=0$, according to lemma 1, and that $b_{\mu}(\iota)|_{k}$ and $e^{\mu}|_{j-k}$ are independent.
By substituting (13) and (14) in (12), it follows that
$\mathsf{Var}(a_{\iota}s^{\iota}|_{i})$ $=\sum_{\mu}\sum_{j=0}^{n-1}\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j})=\sum_{\mu}\sum_{j=0}^{n-1}\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j})\mathbb{E}[s^{\iota}|_{i-j}^{2}]$
$=\sum_{\mu}\sum_{j=0}^{n-1}\sum_{k=0}^{n-1}\mathsf{Var}(b_{\mu}(\iota)|_{k})\mathbb{E}[e^{\mu}|_{j-k}^{2}]\mathbb{E}[s^{\iota}|_{i-j}^{2}].$
Finally, observing that $\mathsf{Var}(b_{\mu}(\iota)|_{i}),\mathbb{E}[e^{\mu}|_{i}^{2}]$ and $\mathbb{E}[s^{\iota}|_{i}^{2}]$ do not depend on $i$, the thesis is demonstrated, i.e.,
$\mathsf{Var}(a_{\iota}s^{\iota}|_{i})=\sum_{\mu}\mathsf{Var}(b_{\mu}(\iota)|_{i})\sum_{k=0}^{n-1}\mathbb{E}[e^{\mu}|_{k}^{2}]\sum_{j=0}^{n-1}\mathbb{E}[s^{\iota}|_{j}^{2}].$

### v3

@@
 For BGV, this threshold strictly depends on the ciphertext modulus, which is one of the initial parameters whose selection heavily affects both the efficiency and security of the scheme.
-In this paper, we provide a new method to estimate noise growth, closely aligning with experimental results and forming the basis for parameter selection that ensures correctness and improves efficiency.
+For an optimal parameter choice, it is crucial to accurately estimate the noise growth, particularly that arising from multiplication, which is the most complex operation.
+In this work, we propose a novel average-case approach that precisely models noise evolution and guides the selection of initial parameters, improving efficiency while ensuring security.
+The key innovation of our method lies in accounting for the dependencies among ciphertext errors generated with the same key, and in providing general guidelines for accurate parameter selection that are library-independent.
 1 Introduction
@@
 Since then, several FHE constructions have been proposed, such as BGV [6], BFV [8, 22], FHEW [21], TFHE [13, 14], and CKKS [12, 11].
-The homomorphic encryption schemes currently in use base their security on the presumed intractability of the Learning with Errors (LWE) problem, [26], and its ring variant (RLWE) [23].
+The homomorphic encryption schemes currently in use base their security on the presumed intractability of the Learning with Errors (LWE) problem [36], and its ring variant (RLWE) [33].
 Informally, the decisional version of RLWE consists of distinguishing polynomial equations $(a,b=s\cdot a+e)\in\mathcal{R}_{q}\times\mathcal{R}_{q}$, perturbed by small noise $e$ (also called error), from uniform random tuples from $\mathcal{R}_{q}\times\mathcal{R}_{q}$, where $\mathcal{R}_{q}=\mathbb{Z}_{q}[x]/\langle x^{n}+1\rangle$ and $q$ is a positive integer.
@@
 The prevailing trend in the current literature adopts the average-case analysis, which involves treating the noise coefficients as random variables distributed according to a Gaussian distribution and studying their expected value and variance.
-Interest in this method, initially applied in the TFHE scheme [9], and subsequently in the CKKS [12], BGV [14, 25] and BFV [2] schemes, grew due to a recognized discrepancy between the estimates based on worst-case technique and experimental data, as highlighted in [13].
+Interest in this method, initially applied in the TFHE scheme [13], and subsequently in the CKKS [16, 5], BGV [18, 35] and BFV [4] schemes, grew due to a recognized discrepancy between the estimates based on worst-case technique and experimental data, as highlighted in [17].
 The introduction of the average-case approach, as seen in [4, 18], offers a potential resolution to these disparities, indeed, with this method, it is possible to compute a tight probabilistic upper bound.
-However, the heuristics used for the BGV [25] and CKKS [12] schemes often underestimate the noise growth due to the assumption of the noises independence, leading to imprecise bounds, as also pointed out in [1, 12, 25].
+However, a recent work [24] criticizes current approaches by showing that the noise coefficients may follow a heavy-tailed distribution, potentially invalidating the assumptions underlying average-case bounds.
+In particular, focusing on BGV without modulus switching, the authors observe a heavy-tailed behavior of the noise.
+In the following, we confirm this behavior for the same setting, but we further prove that when modulus switching is applied, this is no longer the case.
+Specifically, under typical parameter choices used by all major libraries, the noise coefficients in BGV with modulus switching follow a Gaussian distribution, and therefore the average-case noise analysis remains valid.
+Similarly to our demonstration for BGV, a recent work [5] confirms that in CKKS the average-noise analysis remains valid in practice, in line with the observations of [24].
+It is worth noting that, the heuristics used for the BGV [35] and CKKS [16] schemes often underestimate the noise growth due to the assumption of the noises independence, leading to imprecise bounds, as also pointed out in [3, 16, 20, 35].
 Such underestimates lead to two potential issues: first, the ciphertext is not correctly decrypted with non-negligible probability due to excessive noise and, second, the scheme is exposed to security vulnerabilities, as shown in recent papers [9, 10].
@@
 This, in turn, enables the definition of more accurate operational parameters, making the scheme both more secure and efficient, which is essential for the widespread adoption of FHE.
-In this paper, we propose the first average-case noise analysis for BGV that does not provide underestimates, taking in account the dependencies introduced by the common secret and public key.
-We extend the approach of BFV [2], where the authors consider the fact that the errors are not independent among each other, and introduce a function $F$ to “correct” the product of the variances for homomorphic multiplication.
-We notice the additional dependencies given by the public key and employ a second correcting function to address its contribution.
+In this paper, we propose the first average-case noise analysis for BGV that does not provide underestimates, taking into account the dependencies introduced by the common secret and public key.
+We extend the approach of BFV [4], where the authors introduced a correction function $F$ to adjust the product of variances in homomorphic multiplication, by incorporating the dependency effects of the secret key.
+Similarly, we introduce a correction function which, unlike in [4], is no longer heuristic but derived from formal results presented in our work.
+Moreover, we observed that in BGV it is necessary to account for additional dependencies introduced by the public key, requiring the correction function to also compensate for these effects.
 The results obtained in this study suggest that this approach leads to significant improvements in noise analysis.
@@
 Note that these terms are common to all ciphertexts calculated using the same public key and are responsible for the dependence of the noise.
-To account for these dependencies, we applied the correction function $F$ introduced in [2].
-It is important to note that the case of BGV is more complex than that of BFV.
-In BFV, the dependency due to $e$ appears only in a negligible term.
-In contrast, for BGV, we must take into account the dependencies on both $s$ and $e$.
-This requires the use of two distinct correction functions to accurately model and mitigate these dependencies.
 The paper is structured as follows.
-Section 2 introduces essential definitions and fundamental properties that are instrumental for understanding both our contribution and the scheme more broadly.
+Section 2 introduces essential definitions and fundamental properties that are necessary for understanding both our contribution and the functioning of the scheme.
 Section 3 provides a concise overview of the main features and structure of the BGV scheme.
 In Section 4, we present our key results concerning the behavior of the error term and its growth under homomorphic operations.
-Section 5 then demonstrates how the findings from Section 4 can be leveraged to estimate error growth in fixed-operation circuits, laying the groundwork for novel approaches to selecting the scheme’s initial parameters.
+Section 5 provides a discussion on why, in the BGV scheme, the error coefficients can be modeled as samples from a discrete Gaussian distribution.
+Section 6 then demonstrates how the findings from Section 4 and Section 5 can be leveraged to estimate error growth in fixed-operation circuits and to properly select the ciphertext moduli.
+Section 7 compares our approach with state-of-the-art methods, showing how our parameter selection leads to a significant improvement over those currently adopted in major libraries such as OpenFHE and HElib.
 Finally, Section 8 concludes the paper and outlines possible directions for future research inspired by our results.
 2 Preliminaries
-In this section, we define the general notations that we will use in the remainder of the work.
-In our analysis we will describe the leveled version of the BGV scheme, noting that this implies that the parameters of the scheme depend (polynomially) on the depth of the circuits that the scheme is capable of evaluating [4].
-It should be noted that this version is also the most widely adopted because it does not require the usage of the bootstrapping technique, which is highly complex from a computational perspective.
-Before recalling the BGV scheme, some preliminary parameters must be fixed.
-Let $L$ denote the number of levels of the arithmetic circuit that the scheme must be able to evaluate.
-Moreover, assume that $m$ is a suitably chosen integer.
-The notation $\mathcal{R}_{d}$, for a fixed $d\in\mathbb{Z}$, will denote the ring $\mathcal{R}_{d}=\mathbb{Z}_{d}[x]/\langle\Phi_{m}(x)\rangle$, where $\Phi_{m}(x)$ is the $m$-th cyclotomic polynomial, a definition of which is briefly recalled below.
-Definition 1
-Given a positive integer $m$, the $m$-th cyclotomic polynomial is defined as
-$\Phi_{m}(x)=\prod_{\begin{subarray}{c}1\leq j<m\\ \gcd(j,m)=1\end{subarray}}(x-\zeta^{j}),$
-where $\zeta$ is a primitive $m$-th root of unity. This polynomial has degree $\phi(m)$, where $\phi$ denotes the Euler function.
-Eventually, we will denote with $\mathcal{R}$ the ring $\mathcal{R}=\mathbb{Z}[x]/\langle\Phi_{m}(x)\rangle$. To define the plaintext and ciphertext spaces, $L+1$ moduli are selected as follows:
-  • An integer $t$, denoted as the plaintext modulus, is chosen such that $t\equiv 1\mod m$.
-  • $L$ integers $q_{l}$, for $l\in\{0,\dots,L-1\}$, are defined as:
-    $q_{l}=\prod_{j=0}^{l}p_{j},$
-    where the $p_{j}$ are primes satisfying $p_{j}\equiv 1\mod m$ for $j=0,\dots,L-1$ and $p_{1},...,p_{L-2}$ have approximately the same size.
-Each $q_{L-1-l}$ is referred to as the ciphertext modulus for the $l$-th level.
-According to these notations, we can define the plaintext and ciphertext spaces, respectively, as $P=\mathcal{R}_{t}$, $C=\mathcal{R}_{q_{l}}\times R_{q_{l}}$.
-Due to the ciphertext coefficients being reduced according to different moduli at each level, BGV is commonly referred to as scale-dependent.
-We use lowercase letters such as $x$ to denote polynomials, and bold symbols like $\bm{x}$ to represent vectors of polynomials.
-Given a polynomial $x\in\mathcal{R}_{n}$, we use the notation $[x]_{n}$ to refer to the centered representation of $x$ modulo $n$ i.e., the representative of $x$ with coefficients in $[-\frac{n}{2},\frac{n}{2})$.
-In general, unless otherwise specified, we always assume the coefficients of polynomials in $\mathcal{R}_{d}$ to be centered with respect to the modulus $d$.
-Moreover, given a random polynomial $p\in\mathcal{R}$ and a probabilistic distribution $X$, the notation $p\leftarrow X$ is used to indicate that each coefficient of $p$ is chosen independently according to $X$.
+In this section, we define the general notation and provide the mathematical background that we will use throughout the paper.
+2.1 Notation
+Let $\mathbb{Z}$ be the ring of integers, and for $d\in\mathbb{Z}_{>0}$ we denote by $\mathbb{Z}_{d}=\nicefrac{{\mathbb{Z}}}{{d\mathbb{Z}}}$.
+Let $n$ be a power of 2.
+We define $\mathcal{R}$ as the ring $\mathcal{R}=\mathbb{Z}[x]/\langle x^{n}+1\rangle$ and $\mathcal{R}_{d}=\mathbb{Z}_{d}[x]/\langle x^{n}+1\rangle$.
+We use $t$ and $q$ to represent the plaintext and the ciphertext modulus, respectively, and $\mathcal{R}_{t}$ defines the plaintext space, where $t$ is chosen such that $t\equiv 1\mod 2n$.
+Moreover, to define the ciphertext space, we need to select $L=M+1$ moduli, where $M$ is the multiplicative depth of the circuit.
+Then, for each level $\ell\in\{0,\dots,L-1\}$, we denote
+$q_{\ell}=\prod_{j=0}^{\ell}p_{j},$
+as the ciphertext modulus at level L1L-1-$\ell$. Sometimes, the ciphertext modulo $q_{L-1}$ at level 0 will be indicated simply by $q$.
+We use lowercase letters such as $a$ for polynomials and bold letters, like $\bm{a}$, for vectors of polynomials.
+Moreover, we denote by $a|_{i}$ the coefficient of $x^{i}$ of the polynomial $a$.
+For $a\in\mathcal{R}_{d}$, we let $a]_{d}$ define its centered reduction modulo $d$, with coefficients in [d/2,d/2)[-d/2,d/2).
+Sometimes, we will also write $a\mod d$.
+Unless stated otherwise, we assume that the coefficients of the polynomials in $\mathcal{R}_{d}$ are always centered modulo $d$.
+Finally, we recall that for $a,b\in\mathcal{R}$ the $i$-th coefficient of their product is given by (see, e.g., [4])
+$ab|_{i}=\sum_{j=0}^{n-1}\xi(i,j)a|_{j}b|_{i-j},\quad\mbox{ where }\quad\xi(i,j)=\left\{\begin{array}[]{rl}1&\mbox{ for }i-j\in[0,n)\\ -1&\mbox{ otherwise}.\end{array}\right$. (1)
+2.2 Probabilistic distributions
+Given a polynomial $a\in\mathcal{R}$ and a probabilistic distribution $\chi$, the notation $a\leftarrow\chi$ is used to indicate that each coefficient of $a$ is randomly and independently sampled according to $\chi$.
 Some distributions that will be frequently considered are the following:
-  • $\mathcal{U}_{q}$ as the uniform distribution over $\nicefrac{{\mathbb{Z}}}{{q\mathbb{Z}}}$;
-  • $\mathcal{N}(0,\sigma^{2})$ as the normal distribution, also referred to as the Gaussian distribution, over $\mathbb{R}$, with mean 00 and variance $\sigma^{2}$;
+  • $\mathcal{U}_{q}$ as the uniform distribution over $\mathbb{Z}_{q}$ where the representatives modulo $q$ are taken in the interval $\left[-\frac{q}{2},\frac{q}{2}\right)$;
+  • $\mathcal{N}(0,\sigma^{2})$ as the normal distribution, also referred to as the Gaussian distribution, over $\mathbb{R}$, with mean 0 and variance $\sigma^{2}$;
   • $\mathcal{DG}_{q}(\sigma^{2})$ as the discrete Gaussian distribution, which involves sampling a value according to $\mathcal{N}(0,\sigma^{2})$, rounding it to the nearest integer and then reducing it modulo $q$.
-    Moreover, the representative modulo $q$ is taken in the interval $\left[-\frac{q}{2},\frac{q}{2}\right)$;
-For the BGV scheme the notation $\chi_{e},\chi_{s}$ will be adopted in order to indicate the distribution of the error for a RLWE instance and the secret key coefficients, respectively.
+Moreover, the representative modulo $q$ is taken in the interval $\left[-\frac{q}{2},\frac{q}{2}\right)$.
+For the BGV scheme, the notation $\chi_{e}$ and $\chi_{s}$ will be adopted in order to indicate the distribution of the error for a RLWE instance and the secret key coefficients, respectively.
 Typically, $\chi_{e}$ is a discrete Gaussian distribution with a suitable standard deviation, while for the secret key, the ternary uniform distribution $\mathcal{U}_{3}$ is usually considered.
-However, in some special cases, where bootstrapping is needed, the choice for the secret key distribution falls on the Hamming Weight distribution, a definition of which is recalled below [2].
-Definition 2
-The Hamming Weight distribution $\mathcal{HWT}_{n,h}$ consists of selecting uniformly at random a vector of $n$ entries in $\{-1,0,1\}$, where exactly $h<n$ entries are non-zero.
-To ensure the correctness of the scheme, it is important to underline that the error added to the plaintext during encryption must be sufficiently small [4].
-This will become more apparent through the presentation of the BGV functions.
-For the purposes of our subsequent analysis, it is helpful to recall that for two polynomials $r,s\in\mathcal{R}$, and denoting by $r|_{i}$ and $s|_{i}$ their coefficients with respect to $x^{i}$, the $i$-th coefficient of their product is given by [2]
-$rs|_{i}=\sum_{j=0}^{\phi(m)-1}\xi(i,j)r|_{j}s|_{i-j},$ (1)
-where
-$\xi(i,j)=\left\{\begin{matrix}1&\text{\hskip 51.21504ptfor }i-j\in[0,\phi(m))\\[5.0pt] -1&\text{otherwise}\end{matrix}\right.$
-Furthermore, the following notation will be adopted.
-Notation 1
-Given a polynomial $p\in\mathcal{R}_{q}$, for a fixed integer $q$, the variance of its coefficients, namely $\mathsf{Var}(p|_{i})$, is denoted as $V_{p}$.
-Proposition 1
-Let $p,q$ be two independent polynomials in $\mathcal{R}_{q}$, where $q$ is a fixed integer, and let $k$ be an integer in $\nicefrac{{\mathbb{Z}}}{{q\mathbb{Z}}}$.
-Moreover, assume that the coefficients of each polynomial are independent and identically distributed, with 00-mean.
-Then,
-  • a.
-    $V_{p+q}=V_{p}+V_{q}$
-  • b.
-    $V_{kp}=k^{2}V_{p}$
-  • c.
-    $V_{p\cdot q}=\phi(m)V_{p}\cdot V_{q}$
-Finally, the variance values for some common distributions, which will often be employed in the BGV scheme, are:
+However, in some special cases, where bootstrapping is needed, the choice for the secret key distribution falls on the Hamming Weight distribution where the secret key is sampled uniformly among sparse ternary vectors with a fixed number of non-zero entries.
+Moreover, we denote by $V_{a}$ the variance of the coefficients of the polynomial $a$, namely $\mathsf{Var}(a|_{i})$.
+Given $\gamma\in\mathbb{Z}$, if $a,b\in\mathcal{R}$ are two independent polynomials whose coefficients are independent, identically distributed, and have zero mean, then [17]:
+  • $V_{a+b}=V_{a}+V_{b}$
+  • $V_{\gamma a}=\gamma^{2}V_{a}$
+  • $V_{a\cdot b}=nV_{a}\cdot V_{b}$.
+Finally, the values of the variance for some common distributions, which will often be employed in the BGV scheme, are
   • $V_{\mathcal{DG}_{q}(\sigma^{2})}=\sigma^{2}$ for the discrete Gaussian distribution centered at 0 with standard deviation $\sigma$;
@@
   • $V_{q}=\frac{q^{2}-1}{12}\approx\frac{q^{2}}{12}$ for the uniform distribution over integer values in $\left[-\frac{q}{2},\frac{q}{2}\right)$;
+2.3 Infinity and canonical norms
+To conclude this section, we recall the definitions of the infinity and the canonical norm, along with some of their properties.
+Definition 1
+The infinity norm of a polynomial $a\in\mathcal{R}$ is defined as
+$\|a\|_{\infty}=\max_{0\leq i<n}|\hskip 2.84544pta|_{i}\hskip 2.84544pt|.$
+If $a\in\mathcal{R}_{q}$ and its coefficients are well-approximated by identically distributed independent Gaussian variables centered at zero, then
+$\mathbb{P}\left(\|a\|_{\infty}>T\right)\leq n\left(1-\text{erf}\left(\frac{T}{\sqrt{2V_{a}}}\right)\right),$ (2)
+where $\text{erf}(z)$ is the error function, defined as $\text{erf}(z)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}e^{-t^{2}}\,dt$, see, e.g., [4].
+Definition 2
+The canonical embedding norm of a polynomial $a\in\mathcal{R}$ is
+$\|a\|_{can}=\max_{\begin{subarray}{c}1\leq j<2n\\ \gcd(j,2n)=1\end{subarray}}|a(\zeta^{j})|,$
+where $\zeta$ is a primitive $2n$-th root of unity. Essentially, it corresponds to the infinity norm of the canonical embedding of $a$, denoted as $\sigma(a)$.
+[…]
+[…]
+[…]
+We know that if $a\in\mathcal{R}_{q}$ is a random polynomial with coefficient variance $V_{a}$ and $\zeta$ be a primitive $2n^{th}$ root of unity, then the distribution of $a(\zeta)$ is well approximated by a centred Gaussian distribution with variance $nV_{a}$ [20].
+This immediately translates into a bound on the canonical norm of $a$:
+$a||^{can}<D\sqrt{nV_{a}}\,$, (3)
+which holds with probability $(1-e^{-D^{2}/2})^{n}\approx 1-ne^{-D^{2}/2}$. This means that, for a suitable choice of $D$, the bound fails only with negligible probability [20].
 3 The BGV scheme
-Once the parameters have been specified, the framework of the BGV scheme can be defined according to three main functions: Key Generation, Encryption and Decryption.
-3.1 Key Generation
-The key generation function is responsible for producing:
-  • the secret key, represented by a polynomial in $\mathcal{R}_{q_{L-1}}$;
-  • the public key, which consists of a pair of polynomials, each belonging to $\mathcal{R}_{q_{L-1}}$;
-To achieve this, three polynomials in $\mathcal{R}_{q_{L-1}}$ are generated according to the following distributions: $s\leftarrow\chi_{s}$ , $a\leftarrow\mathcal{U}_{q_{L-1}}$ and $e\leftarrow\chi_{e}$.
-The resulting keys are given by
-$\begin{cases}sk&=s\\ pk&=(b,a)\equiv(-a\cdot s+te,a)\mod q_{L-1}\end{cases}$
-3.2 Encryption
-Given the plaintext $m\in\mathcal{R}_{t}$ and the public key $pk=(b,a)$, the encryption function returns as output $(\bm{c},q_{l},\nu_{clean})$ where:
-  • $\bm{c}\in\mathcal{R}_{q_{l}}\times\mathcal{R}_{q_{l}}$ is the actual ciphertext defined as
-    $\bm{c}=(c_{0},c_{1})\equiv(b\cdot u+te_{0}+m,a\cdot u+te_{1})\mod q_{l},$
-    where $u,e_{0},e_{1}\in\mathcal{R}_{q_{l}}$, with coefficients distributed as $u\leftarrow\chi_{s}$ and $e_{0},e_{1}\leftarrow\chi_{e}$.
-  • $q_{l}$ represents the ciphertext modulus for the level at which computations are performed.
-Specifically, as the encryption function is computed in the first level 00, the corresponding $q_{l}$ is $q_{L-1}$.
-  • $\nu$ is the critical quantity associated to the ciphertext and is defined as
-    $\nu=[c_{0}+c_{1}\cdot s]_{q_{l}}.$
-We refer to the triple $\tilde{c}=(\bm{c},q_{l},\nu_{clean})$ with the term extended ciphertext.
-Broadly speaking, the encryption of the plaintext message results in a pair of ciphertext values: the first encodes the masked plaintext using the public key, while the second supplies auxiliary information necessary for the receiver to recover the original message, provided they possess the secret key.
-Regarding the critical quantity, it is important to emphasize that this value is strictly related to the error introduced by the encryption operator.
-Thus, it essentially determines whether the ciphertext can be correctly decrypted or not.
-For this reason, its behavior is analyzed to gain a clear understanding of the functioning of the scheme.
-3.3 Decryption
-At any level, it should be possible to recover the original plaintext using the decryption function.
-Given the extended ciphertext $(\bm{c},q_{l},\nu)$ and the secret key $sk=s$, the plaintext is recovered performing the following computations
-$m=\left[\left[c_{0}+c_{1}\cdot s\right]_{q_{l}}\right]_{t}.$
-Using the previously established notation, the critical quantity can be rewritten as
-$[c_{0}+c_{1}\cdot s]_{q_{l}}=\left[m+t(e\cdot u+e_{1}\cdot s+e_{0})\right]_{q_{l}}=\left[m+t\epsilon\right]_{q_{l}},$
-where $t\epsilon$ denotes the error introduced during encryption.
-By considering the reduction modulo $t$ of the critical quantity, it is possible to verify that the plaintext is successfully recovered.
-However, if the error is too large, the value $m+t(e\cdot u+e_{1}\cdot s+e_{0})$ could wrap around the modulus, resulting in an incorrect decryption.
-So, decryption is correct only if the error magnitude remains below a certain threshold, as previously emphasized.
-In addition, the error associated to the ciphertext $\bm{c}$ at level $l$ increases each time the homomorphic circuit is evaluated.
-Therefore, for constructing a leveled homomorphic encryption scheme, it is crucial to estimate and manage this error properly.
-To deal with the error effectively, the magnitude of the critical quantity, often referred to as noise, is typically analyzed using its norm.
-In the context of homomorphic encryption, the infinity and the canonical norm, a definition of which are briefly recalled below [19], assume a central role.
+In this section, we recall the three basic encryption functions of the BGV scheme and the homomorphic operations.
+3.1 Basic encryption functions
+Key Generation.
+The key generation function generates $s\leftarrow\chi_{s}$ , $a\leftarrow\mathcal{U}_{q_{L-1}}$ and $e\leftarrow\chi_{e}$ and outputs the secret key: $\mathsf{sk}=s\in\mathcal{R}_{q_{L-1}}$, and the public key $\mathsf{pk}=(b,a)\equiv(-a\cdot s+te,a)\mod q_{L-1}$.
+Encryption.
+Given the plaintext $m\in\mathcal{R}_{t}$ and the public key $\mathsf{pk}=(b,a)$, the encryption function outputs the ciphertext $\bm{c}\in\mathcal{R}_{q_{L-1}}^{2}$ defined as
+$\bm{c}=(c_{0},c_{1})\equiv(b\cdot u+te_{0}+m,a\cdot u+te_{1})\mod q_{L-1},$
+where $u,e_{0},e_{1}\in\mathcal{R}_{q_{L-1}}$, with coefficients distributed as $u\leftarrow\chi_{s}$ and $e_{0},e_{1}\leftarrow\chi_{e}$.
+Decryption.
+Given a ciphertext $\bm{c}\in\mathcal{R}_{q_{\ell}}^{2}$ and the secret key $\mathsf{sk}=s$, the plaintext is recovered performing the following computations $m=\left[\left[c_{0}+c_{1}\cdot s\right]_{q_{\ell}}\right]_{t}$.
+We denote by $\nu=[c_{0}+c_{1}\cdot s]_{q_{\ell}}$ the critical quantity corresponding to $\bm{c}$. In particular, for a fresh ciphertext, it can be rewritten as
+$\nu_{\mathsf{clean}}=[c_{0}+c_{1}\cdot s]_{q_{L-1}}=[m+t(e\cdot u+e_{1}\cdot s+e_{0})]_{q_{L-1}}=[m+t\epsilon]_{q_{L-1}},$
+where $\epsilon$ denotes the error introduced during encryption.
+Considering the reduction modulo $t$ of the critical quantity, it is possible to verify that the plaintext is successfully recovered.
+However, if the error is too large, the value $m+t\epsilon$ could wrap around the modulus, resulting in an incorrect decryption.
+So, the decryption is correct only if the coefficients of $m+t\epsilon$ remain below a certain threshold.
+In addition, the error associated to the ciphertext increases through homomorphic operations [15].
+Therefore, it is crucial to estimate the magnitude of the critical quantity (called also noise) which is typically analyzed using its norm.
+In light of this, to guarantee the correctness of the decryption, the condition on the critical quantity can be expressed as follows [34]:
+[…]
+Naturally, in order to bound the noise, any type of norm could be used.
+Finally, another concept that is often introduced for the estimation of error growth is the noise budget, which represents the number of bits remaining before wrap-around would occur.
+We will use the term extended ciphertext to refer to the tuple $\mathfrak{c}=(\bm{c},q_{\ell},\nu)$ where $\bm{c}$ is the actual ciphertext, $q_{\ell}$ is the ciphertext modulus, and $\nu$ is the associated critical quantity.
 Definition 3
-The infinity norm of a polynomial $p\in\mathcal{R}$, of the form $p=p|_{0}+p|_{1}x+...+p|_{\phi(m)-1}x^{\phi(m)-1}$ is defined as
-$\|p\|_{\infty}=\max_{0\leq i<\phi(m)}|\hskip 2.84544ptp|_{i}\hskip 2.84544pt|.$
-Definition 4
-The canonical embedding norm of a polynomial $p\in\mathcal{R}$ is defined as
-$\|p\|_{can}=\max_{\begin{subarray}{c}1\leq j<m\\ \gcd(j,m)=1\end{subarray}}|p(\zeta^{j})|,$
-where $\zeta$ is a primitive $m$-th root of unity.
-Essentially, it corresponds to the infinity norm of the canonical embedding of $p$, denoted as $\sigma(p)$, which is the $\phi(m)$-vector defined as
-$\sigma(p)=(p(\zeta^{i}))_{i},$
-where $1\leq i<m$ and $\gcd(j,m)=1$.
-A relationship between these two norms is given by the following property
-$||p||_{\infty}\leq\alpha_{m}||p||_{can}\hskip 28.45274pt\forall p\in\mathcal{R},$
-where $\alpha_{m}$ is a constant related to the ring $\mathcal{R}$, which depends exclusively on $m$.
-Another property, which will be useful in the context of the noise growth analysis, is recalled below [15].
-Proposition 2
-Let $p,q$ be two polynomials in $\mathcal{R}$. Then,
-$||p\cdot q||_{\infty}\hskip 3.41418pt\leq\phi(m)||p||_{\infty}\cdot||q||_{\infty}$
-$||p\cdot q||_{can}\leq||p||_{can}\cdot||q||_{can}$
-In light of this, to guarantee the correctness of the decryption, the condition on the critical quantity can be expressed as follow
-$||\nu||_{\infty}\leq\alpha_{m}||\nu||_{can}<\frac{q_{l}}{2},$
-which allows to avoid the error wrapping around the modulus [24].
-Naturally, in order to bound the noise, any type of norm could be used.
-Nevertheless, in the state-of-the-art analysis of noise growth for homomorphic schemes, the canonical norm is typically preferred, due to its properties, as it permits obtaining more accurate estimates.
-Finally, another concept that is often introduced for the error growth estimation is the noise budget.
-Definition 5
-Let $(\bm{c},q_{l},\nu)$ be an extended ciphertext. The noise budget associated to $\bm{c}$ is the quantity
-$\log_{2}(q_{l})-\log_{2}(||\nu||)-1,$
-where |||||$\cdot$||| | ⋅ | | refers to a fixed norm.
-The key operations involved in the BGV scheme are briefly summarized in the following.
-3.4 Ciphertext Addition and Constant Multiplication
-Assume that $\bm{c}=(c_{0},c_{1})$ and $\bm{c^{\prime}}=(c_{0}^{\prime},c_{1}^{\prime})$ are two ciphertexts defined with respect to the same modulus $q_{l}$ and computed encrypting two plaintexts $m$ and $m’$ ’, respectively, using the same key.
-Their homomorphic sum is defined as
-$\mathsf{Add}(\bm{c},\bm{c^{\prime}}):=(c_{0}+c_{0}^{\prime},\;c_{1}+c_{1}^{\prime})\mod q_{l}.$
-The resulting ciphertext encrypts the sum $m+m^{\prime}$, and the corresponding noise grows additively. In fact, the resulting critical quantity is defined as
-$\nu_{\mathsf{add}}=\nu+\nu^{\prime},$
-where $\nu=c_{0}+c_{1}s$ and $\nu^{\prime}=c_{0}^{\prime}+c_{1}^{\prime}s$ denote the critical quantities of $\bm{c}$ and $\bm{c^{\prime}}$, respectively.
-Similarly, let $\bm{c}=(c_{0},c_{1})$ be a ciphertext and $k\in R_{t}$ a constant polynomial. The homomorphic multiplication by $k$ is performed as follows
-$\mathsf{Mul}_{k}(\bm{c}):=(kc_{0},\;kc_{1})\mod q_{l},$
-and the corresponding critical quantity scales by $k$, yielding
-$\nu_{\mathsf{const}}=k\cdot\nu.$
-In both cases, decryption correctness is preserved as long as the resulting noise does not cause the message to wrap around the modulus.
-3.5 Homomorphic multiplication and Key switching
-Homomorphic multiplication is one of the fundamental operations supported by the BGV scheme.
-Given two ciphertexts $\bm{c}=(c_{0},c_{1})$ and $\bm{c^{\prime}}=(c_{0}^{\prime},c_{1}^{\prime})$, and assuming that both are defined with respect to the same modulus $q_{l}$, their product is given by
-$\mathsf{Mul}(\bm{c},\bm{c^{\prime}}):=(c_{0}^{\mathsf{mul}},c_{1}^{\mathsf{mul}},c_{2}^{\mathsf{mul}})=(c_{0}\cdot c_{0}^{\prime},c_{0}\cdot c_{1}^{\prime}+c_{1}\cdot c_{0}^{\prime},c_{1}\cdot c_{1}^{\prime})\mod q_{l}.$
+[…]
+3.2 Homomorphic Operations
+Ciphertext Addition.
+Let $(\bm{c},q_{\ell},\nu)$ and $(\bm{c}^{\prime},q_{\ell},\nu^{\prime})$ be two extended ciphertexts. Their homomorphic sum is
+$\mathsf{Add}(\bm{c},\bm{c^{\prime}}):=(c_{0}+c_{0}^{\prime},\;c_{1}+c_{1}^{\prime})\mod q_{\ell}.$
+with resulting critical quantity $\nu_{\mathsf{add}}=\nu+\nu^{\prime}$.
+Constant Multiplication.
+Let $(\bm{c},q_{\ell},\nu)$ be an extended ciphertext and $k\in\mathcal{R}_{t}$ a fixed polynomial. The homomorphic multiplication by $k$ is
+$\mathsf{Mul}_{k}(\bm{c}):=(kc_{0},\;kc_{1})\mod q_{\ell},$
+and the corresponding critical quantity scales by $k$, yielding $\nu_{\mathsf{const}}=k\cdot\nu$.
+Homomorphic multiplication.
+Let $\bm{c},\bm{c^{\prime}}$ be two ciphertexts defined in $\mathcal{R}_{q_{\ell}}$, then
+$\mathsf{Mul}(\bm{c},\bm{c^{\prime}}):=(d_{0},d_{1},d_{2})=(c_{0}\cdot c_{0}^{\prime},c_{0}\cdot c_{1}^{\prime}+c_{1}\cdot c_{0}^{\prime},c_{1}\cdot c_{1}^{\prime})\mod q_{\ell}.$
 As a result, the ciphertext expands from two to three polynomials, which violates the compactness property and makes subsequent operations more costly.
-This phenomenon becomes even more problematic when performing multiple multiplications.
-Recovering the message from $\mathsf{Mul}(\bm{c},\bm{c^{\prime}})$ requires computing the reduction modulo $t$ of the resulting critical quantity given by $\nu_{\mathsf{mul}}=c_{0}^{\mathsf{mul}}+c_{1}^{\mathsf{mul}}s+c_{2}^{\mathsf{mul}}s^{2}$.
-It is significant to note that, unlike other operations, multiplication causes the noise to grow multiplicatively, making it the most critical case to handle among the basic homomorphic operations in terms of error growth.
-To address this, a technique known as relinearization, or key switching, is employed.
-The key idea is to reduce the ciphertext back to its original size while preserving its correctness.
-In practice, this means converting a ciphertext of the form $(c_{0},c_{1},c_{2})$ into a new, equivalent one of the form $(\hat{c}_{0},\hat{c}_{1})$, using auxiliary data called the key switching key.
-More intuitively, key switching allows the expression $c_{0}+c_{1}s+c_{2}s^{2}$ to be re-expressed as a linear polynomial in $s$, restoring the standard noise form.
-Although the relinearization step introduces additional noise, it is essential for maintaining the practicality of the scheme.
+Recovering the message from $\mathsf{Mul}(\bm{c},\bm{c^{\prime}})$ requires computing the reduction modulo $t$ of the resulting critical quantity given by $\nu_{\mathsf{mul}}=d_{0}+d_{1}s+d_{2}s^{2}$.
+To modify the ciphertext polynomial $d_{0}+d_{1}s+d_{2}s^{2}$ back to another polynomial $\bar{c}_{0}+\bar{c}_{1}\cdot s$ encrypting the same plaintext, a technique known as relinearization, or key switching, is employed.
+Key switching.
+Intuitively, the key switching technique converts $d_{2}s^{2}$ into $\hat{c}_{0}+\hat{c}_{1}s$ using somehow the encryption of $s^{2}$ under $s$. Indeed,
+$\mathsf{Enc}_{s}(s^{2})=(\beta,\alpha)=(-us+te+s^{2},u+te_{1})\approx(\alpha s+s^{2},-\alpha).$
+Thus, $s^{2}\approx\beta-\alpha s$ and then $d_{0}+d_{1}s+d_{2}s^{2}\approx d_{0}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+d_{1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s+d_{2}(\beta-\alpha s)=\bar{c}_{0}+\bar{c}_{1}s$.
+As expected, while the relinearization step introduces additional noise, it is crucial for ensuring the practicality of the scheme.
 Several key switching techniques have been proposed in the literature, each aiming to optimize this trade-off between correctness and noise growth.
-The most commonly used are the Brakerski Vaikuntanathan (BV) variant, the Gentry Halevi Smart (GHS) variant and the Hybrid variant, which can be considered as a combination of the previous ones.
-3.6 Modulus switching
-The primary aim of the modulus switching technique is to reduce the noise resulting from homomorphic evaluations.
-In practical circuits, modulus switching is typically applied only after homomorphic multiplications, due to their higher cost in terms of error growth.
-Similarly to the bootstrapping technique, the main purpose is to transition from an extended ciphertext $(\bm{c},q_{l},\nu)$ to another $(\bm{c}^{\prime},q_{l^{\prime}},\nu^{\prime})$ such that the error associated to the latter is smaller, while preserving the fact that, using the same key, both can be decrypted to the same original message.
-The key difference lies in how this is achieved, and it is in this aspect that modulus switching demonstrates greater efficiency compared to bootstrapping.
-Let $(\bm{c},q_{l},\nu)$ be the extended ciphertext, whose error we want to reduce, and let $l^{\prime}$ be an integer, such that $q_{l^{\prime}}<q_{l}$.
-The new ciphertext produced by the modulus switch is given by
-$\bm{c}^{\prime}=\frac{q_{l^{\prime}}}{q_{l}}(\bm{c}+\bm{\delta})\mod q_{l^{\prime}},$
-where $\bm{\delta}=t[-\bm{c}t^{-1}]_{\frac{q_{l}}{q_{l^{\prime}}}}$.
-The $\bm{\delta}$ value can be interpreted as a correction required to ensure that the ciphertext is divisible by $\frac{q_{l}}{q_{l^{\prime}}}$ and does not affect the original message.
+The most commonly used are the Brakerski Vaikuntanathan (BV) variant [7], the Gentry Halevi Smart (GHS) variant [25], and the Hybrid variant [25], which can be considered as a combination of the previous ones.
+Herein, we do not delve into the details of each method and refer the reader to [34] for further information.
+Modulus switching.
+The primary aim of the modulus switching technique is to reduce the noise resulting from homomorphic operations.
+Let $(\bm{c},q_{\ell},\nu)$ be the extended ciphertext whose error we aim to reduce, and let $\ell^{\prime}$ be an integer such that $q_{\ell^{\prime}}<q_{\ell}$.
+The modulus switching procedure outputs $(\bm{c}^{\prime},q_{\ell^{\prime}},\nu^{\prime})$, where
+$\bm{c}^{\prime}=\frac{q_{\ell^{\prime}}}{q_{\ell}}(\bm{c}+\bm{\delta})\mod q_{\ell^{\prime}},$
+with $\bm{\delta}=t[-\bm{c}t^{-1}]_{{q_{\ell}}/{q_{\ell^{\prime}}}}$.
+The $\bm{\delta}$ value can be interpreted as a correction required to ensure that the ciphertext is divisible by ${q_{\ell}}/q_{\ell^{\prime}}$ and does not affect the original message.
 In fact, it only influences the error since $\bm{\delta}\equiv 0\mod t$.
-Therefore, the new ciphertext $\bm{c}^{\prime}$ will still decrypt to the original plaintext (scaled by a factor of $\frac{q_{l}}{q_{l^{\prime}}}$), provided that the necessary conditions are met.
-The critical quantity associated to the new ciphertext $\bm{c^{\prime}}$ can be expressed in terms of that of $\bm{c}$, namely $\nu$, as
-$\nu_{\mathsf{modswitch}}$ $=[c_{0}^{\prime}+c_{1}^{\prime}\cdot s]_{q_{l^{\prime}}}=\frac{q_{l^{\prime}}}{q_{l}}([c_{0}+c_{1}\cdot s]_{q_{l}}+\delta_{0}+\delta_{1}\cdot s)=\frac{q_{l^{\prime}}}{q_{l}}(\nu+\delta_{0}+\delta_{1}\cdot s).$
-To conclude, it is worth noting that this correctness verification implicitly relies on the common value $h$ in the two equalities
-$\begin{split}[c_{0}+c_{1}\cdot s]_{q_{l}}\hskip 2.84544pt&=c_{0}+c_{1}\cdot s-hq_{l}\\ [c_{0}^{\prime}+c_{1}^{\prime}\cdot s]_{q_{l^{\prime}}}&=c_{0}^{\prime}+c_{1}^{\prime}\cdot s-hq_{l^{\prime}}\end{split}$
-The reason why the value $h$ is the same for both the expressions can be found in [4, Lemma 1].
+Therefore, the new ciphertext $\bm{c}^{\prime}$ will still decrypt to the original plaintext (scaled by a factor of ${q_{\ell}}/q_{\ell^{\prime}}$).
+The critical quantity associated to the new ciphertext $\bm{c^{\prime}}$ can be expressed in terms of that of $\bm{c}$, namely,
+$\nu_{\mathsf{ms}}$ $=[c_{0}^{\prime}+c_{1}^{\prime}\cdot s]_{q_{\ell^{\prime}}}=\frac{q_{\ell^{\prime}}}{q_{\ell}}([c_{0}+c_{1}\cdot s]_{q_{\ell}}+\delta_{0}+\delta_{1}\cdot s)=\frac{q_{\ell^{\prime}}}{q_{\ell}}(\nu+\delta_{0}+\delta_{1}\cdot s).$
 4 Average-Case Noise Analysis for BGV
 The aim of this section is to investigate the behavior of the noise resulting from the main homomorphic operations supported by the BGV scheme.
-Throughout this analysis, we consider ciphertexts that are mutually independent, obtained by encrypting independently generated random messages under the same public key.
+Throughout this analysis, we consider ciphertexts that are mutually independent, obtained by encrypting independently generated random messages using the same public key.
 As previously mentioned, the novel approach introduced in this paper seeks to analyze noise growth by accounting for dependencies among the coefficients of the critical quantities involved.
 Before delving into the details, it is important to highlight that in BGV, the critical quantity resulting from homomorphic operations can be affected by such dependencies, even when the ciphertexts involved are independent.
-These dependencies stem from the fact that the noise in the ciphertexts includes terms involving powers of the secret key $s$ and powers of the error term $e$, which makes it necessary to explicitly consider these contributions when analyzing the variance of the noise.
+These dependencies arise because the noise in the ciphertexts includes terms involving powers of the secret key $s$ and powers of the error term $e$, which makes it necessary to explicitly consider these contributions when analyzing the variance of the noise.
 To study the impact of $s$ and $e$, we isolate their contribution in the expression of the critical quantity $\nu$, using the following notation:
@@
 where $a_{\iota}=\sum_{\mu}b_{\mu}(\iota)e^{\mu}$, and $b_{\mu}(\iota)$ contains no powers of $s$ or $e$.
-To enhance clarity, for the critical quantity $\nu$ of a fresh ciphertext $\bm{c}$, this notation yields
-$\nu=a_{0}+a_{1}s=b_{0}(0)+b_{1}(0)e+b_{0}(1)s,$
+To enhance clarity, for the critical quantity $\nu_{\mathsf{clean}}$ of a fresh ciphertext $\bm{c}$, this notation yields
+$\nu_{\mathsf{clean}}=a_{0}+a_{1}s=b_{0}(0)+b_{1}(0)e+b_{0}(1)s,$
 where
@@
 We then proceed to describe how these properties are used to estimate the variance of the error coefficients after homomorphic multiplications, and finally how such estimates can be applied to circuits consisting of fixed sequences of operations.
-From this point onward, we adopt the notation $n$ to denote $\phi(m)$, in order to simplify the presentation. It is interesting to point out that when $n$ is a power of two, $m=2n$.
-4.1 Distribution Analysis
-As previously mentioned, the noise coefficients are treated as random variables.
-We observed that their distributions are well-approximated by identically distributed, dependent Gaussian variables centered at zero.
-This empirical observation was validated through experiments using the Python fitter package, which confirmed the Gaussian nature of the noise coefficients.
-This property is particularly advantageous, as it allows us to bound the maximum absolute value of the noise coefficients with high probability by simply controlling their variance $V$.
-In particular, we recall the following fundamental property of Gaussian random variables.
-Proposition 3
-Let $Z$ be a real-valued, zero-mean Gaussian random variable with variance $V$. Then, for any fixed value $z>0$, the probability that $Z$ lies within the interval $(-z,z)$ is given by
-$\mathbb{P}(|Z|<z)=\text{erf}\left(\frac{z}{\sqrt{2V}}\right),$
-where $\text{erf}(z)$ is the error function, defined as
-[…]
-Extending this result to a vector $\bm{Z}\in\mathbb{R}^{n}$, whose entries are distributed as $Z$, we obtain the following bound on the tail probability of the infinity norm
-$\mathbb{P}(\|\bm{Z}\|_{\infty}>z)\leq n\left(1-\text{erf}\left(\frac{z}{\sqrt{2V}}\right)\right).$ (2)
-This implies that, in order to ensure that a noise vector $\nu$ satisfies the correctness condition $\|\nu\|_{\infty}<{q}/{2},$ we can use Proposition 3 to bound the failure probability as
-$\mathbb{P}\left(\|\nu\|_{\infty}>\frac{q}{2}\right)\leq n\left(1-\text{erf}\left(\frac{q}{2\sqrt{2V}}\right)\right),$
-where $V$ denotes the estimated variance of the noise coefficients. To express this bound more conveniently, we introduce a security parameter $D$ such that
-$D\leq\frac{q}{2\sqrt{2V}},$
-from which it follows, using the monotonicity of the error function, that
-$\mathbb{P}\left(\|\nu\|_{\infty}>\frac{q}{2}\right)\leq n\left(1-\text{erf}(D)\right).$
-Thus, by appropriately choosing the security parameter $D$, we can ensure that the probability of decryption failure remains negligibly small.
-For instance, setting $D=6$ and $n=2^{13}$ yields a failure probability of approximately $2^{-42}$.
-For practical applications, $D=8$ should be preferred, resulting in a probability of approximately $2^{-83}$, when the ring dimension is $n=2^{13}$.
-Consequently, the ciphertext modulus $q$ can be selected according to the inequality
-$q\geq 2D\sqrt{2V},$ (3)
-which highlights the importance of accurately estimating $V$.
-It is important to emphasize that the bounds derived in our analysis provide insight into the minimum ciphertext modulus $q$ required to guarantee the correctness of the scheme.
-Although larger values of $q$ may be used in practice, identifying the minimal admissible value is essential for optimizing performance and efficiency.
-This is particularly relevant in the context of the BGV scheme, where our results can serve as a guideline for carefully selecting the setting parameters.
-By determining tight lower bounds on $q$, it is possible to lay the foundation for sound and efficient parameter selection that ensures correctness without over-provisioning resources.
 4.1 Mean and Variance Analysis
-As previously discussed, we have demonstrated that the coefficients of the error term are centered at zero.
-Furthermore, we have shown that the coefficients of the $b_{\mu}$ terms are uncorrelated, meaning that their pairwise covariance is zero.
+In the following, we prove that the coefficients of the error term are centered at zero.
+Furthermore, we show that the coefficients of the $b_{\mu}$ terms are uncorrelated, meaning that their pairwise covariance is zero.
 Lemma 1
@@
 Let $\nu=\sum_{\iota}a_{\iota}s^{\iota}$ represent the critical quantity associated with a given ciphertext, where, for a fixed $\iota$, $a_{\iota}=\sum_{\mu}b_{\mu}(\iota)e^{\mu}$.
-Then, the following equivalence holds
-$\mathsf{Var}(a_{\iota}s^{\iota}|_{i})=\sum_{\mu\geq 0}\mathsf{Var}(b_{\mu}(\iota)|_{i})\sum_{k=0}^{n-1}\mathbb{E}[e^{\mu}|_{k}^{2}]\sum_{j=0}^{n-1}\mathbb{E}[s^{\iota}|_{j}^{2}]$
+Then, the following identity holds
+$\mathsf{Var}(a_{\iota}s^{\iota}|_{i})=\sum_{\mu\geq 0}\mathsf{Var}(b_{\mu}(\iota)|_{i})\sum_{k=0}^{n-1}\mathbb{E}[e^{\mu}|_{k}^{2}]\sum_{j=0}^{n-1}\mathbb{E}[s^{\iota}|_{j}^{2}].$
+Moreover, $\mathsf{Var}(\nu|_{i})=\sum_{\iota,\mu\geq 0}\mathsf{Var}(b_{\mu}(\iota)|_{i})\sum_{k=0}^{n-1}\mathbb{E}[e^{\mu}|_{k}^{2}]\sum_{j=0}^{n-1}\mathbb{E}[s^{\iota}|_{j}^{2}].$
 A proof of Lemma 2 can be found in Appendix 0.B.
+4.2 Homomorphic Operations
+We can state our results on the variance computation for operations, especially focusing on the multiplication in the next section.
+Proposition 1(Encryption)
+The invariant noise $\nu_{\mathsf{clean}}$ of a fresh ciphertext has coefficient variance
+$V_{\mathsf{clean}}=\mathsf{Var}(\nu_{\mathsf{clean}}|_{i})=\tfrac{t^{2}}{q^{2}}\left(\tfrac{1}{12}+nV_{e}V_{u}+V_{e}+nV_{e}V_{s}\right).$ (4)
+Proof
+The fresh error $\nu_{\mathsf{clean}}$ can be written as $a_{0}+a_{1}s=b_{0}(0)+b_{1}(0)e+b_{0}(1)s$, where
+$\begin{cases}a_{0}=b_{0}(0)+b_{1}(0)\cdot e=(m+te_{0})+tu\cdot e\\ a_{1}=b_{0}(1)=te_{1}.\end{cases}$
+The proof is thus concluded by applying Lemma 2 and observing that $\mathbb{E}[s|_{j}^{2}]=\mathsf{Var}(s|_{j})=V_{s}$ and $\mathbb{E}[e|_{j}^{2}]=\mathsf{Var}(e|_{j})=V_{e}$ since $\mathbb{E}[s|_{j}]=\mathbb{E}[e|_{j}]=0$.
+Proposition 2(Addition & Constant Multiplication)
+Let $\alpha\in\mathcal{R}_{t}$ and $\mathbf{c}$, $\mathbf{c}^{\prime}$ be two independently-computed ciphertexts with invariant noises $\nu,$ and $\nu^{\prime}$, respectively.
+Then, the variance of the error coefficients
+  • resulting from the addition of $\mathbf{c}$ and $\mathbf{c}^{\prime}$ is
+    $\mathsf{Var}(\nu_{\mathsf{add}}|_{i})=\mathsf{Var}(\nu|_{i})+\mathsf{Var}(\nu^{\prime}|_{i}).$ (5)
+  • after a multiplication between $\alpha$ and $\mathbf{c}$ is
+    $\mathsf{Var}(\nu_{\mathsf{const}}|_{i})=\tfrac{(t^{2}-1)n}{12}\mathsf{Var}(\nu|_{i}).$ (6)
+Proposition 3(Modulo Switch)
+Let $\mathfrak{c}=(\mathbf{c},q_{\ell},\nu)$ be an extended ciphertext. The variance of the error coefficients after the modulo switch to the target modulo $q^{\prime}_{\ell}$ is
+$V(\nu_{\mathsf{ms}}|_{i})=\tfrac{q_{\ell}^{\prime 2}}{q_{\ell}^{2}}\mathsf{Var}(\nu|_{i})+\tfrac{t^{2}}{12}(1+nV_{s}).$ (7)
+Propositions 2 and 3 follow directly from Lemma 2 by arguments analogous to those in Proposition 1. The proofs are therefore omitted.
+We do not explicitly address key switching here, since its negligible impact is discussed in Appendix 0.C.
 4.3 Homomorphic multiplications
-The main goal of our analysis is to provide estimates on the growth of the error resulting from homomorphic multiplications, which are the most complex operations and therefore the most interesting to study.
-It thus becomes crucial to account for the dependencies that arise among the noise coefficients as a consequence of these multiplications.
-To this end, the approach we adopted relies on the correction functions introduced in [2].
-Therefore, before stating the theorem, some of the key properties of $F$ will be recalled.
-Unlike the BFV scheme, the BGV scheme requires additional care in the use of the correction function $F$, as not only the impact of the term $s$ must be taken into account, but also that of the term $e$, which is negligible in BFV.
-Consequently, our analysis requires the introduction of two distinct correction functions, which will be denoted as $F_{e}$ and $F_{s}$.
-To define these two functions, we need the following
-Heuristic 1
-Let $x$ be a polynomial whose coefficients are independent and identically distributed according to a fixed distribution. The function $f_{x}$ defined as
-$f_{x}(\iota)=-e^{\alpha-\beta\iota-\gamma\iota^{2}}+\delta,$ (4)
-where $\alpha,\beta,\gamma,\delta$ depend only on the coefficient distribution of $x$ and the ring dimension $n$, satisfies the following
-$f_{x}(\iota)\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[x^{\iota}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[x^{\iota-1}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[x|_{i_{2}}^{2}]}.$
-In Table 1, we report the values of $\alpha,\beta,\gamma,\delta$ required to define the function $f_{s}$, where $s$ denotes the secret key with coefficients sampled from the ternary distribution $\chi_{s}=\mathcal{U}_{3}$, as in [2].
-        $n$         $\alpha$         $\beta$         $\gamma$         $\delta$
-        $2^{12}$         2.8732         0.0160         0.0049         19.1895
-        $2^{13}$         2.9644         0.0196         0.0046         20.4747
-        $2^{14}$         2.9578         0.0386         0.0032         19.5755
-        $2^{15}$         2.9765         0.0197         0.0043         20.7760
-Table 1: Parameters for $\chi_{s}=\mathcal{U}_{3}$
-Similarly, it is possible to define the function $f_{e}$ for the error term $e$, whose coefficients are distributed according to a discrete Gaussian distribution.
-For completeness, a table containing the parameters depending on the distribution $\chi_{e}$ can be found in Table 2 [2].
-        $n$         $\alpha$         $\beta$         $\gamma$         $\delta$
-        $2^{12}$         2.9000         0.0157         0.0051         19.5356
-        $2^{13}$         2.9340         0.0042         0.0055         20.7063
-        $2^{14}$         2.9138         0.0290         0.0039         19.2973
-        $2^{15}$         2.9511         0.0129         0.0046         20.7263
-Table 2: Parameters for $\chi_{e}=\mathcal{DG}_{q}(\sigma^{2})$, $\sigma=3.19$
+The main goal of our analysis is to estimate the error growth caused by homomorphic multiplications, the most critical operations to handle.
+This requires accounting for the dependencies among noise coefficients induced by such multiplications.
+Our method relies on a correction function $F$, whose construction is detailed in this section.
+We recall that the function $F$ was introduced in [4] to “correct” the product of the variances, since these are not independent.
+Unlike BFV, the BGV scheme demands particular care, as both terms $s$ and $e$ must be considered, whereas the contribution of $e$ is negligible in BFV.
+Furthermore, unlike [4], our construction of $F$ is not heuristic but is based on Lemma 3.
+We recall the Isserlis theorem which will be exploited in the proof of the Lemma.
+Theorem 4.1
+[32, Chapter 8] Let $(X_{1},\ldots,X_{n})$ be a zero-mean multivariate normal random vector, then
+$\mathbb{E}[X_{1}\cdots X_{n}]=\sum\prod_{i,j}\mathbb{E}[X_{i}X_{j}]\,$,
+where the sum is over all the partition of $\{1,\ldots,n\}$ into pairs and the $(i,j)$ ranges in these pairs.
+Lemma 3
+Let $a(x)=a|_{0}+a|_{1}x+\dots+a|_{n-1}x^{n-1}\in\mathcal{R}$, where $a|_{i}$’s are i.i.d. random variables with $\mathbb{E}[a|_{i}]=0$, $\mathbb{E}[a|_{i}^{2}]=V_{a}$.
+Then
+$\mathbb{E}[a^{k}|_{i}]=0,\quad\mathbb{E}[(a^{k}|_{i})^{2}]=k!n^{k-1}V_{a}^{k}\,$,
+for all $0\leq i\leq n-1$, where $a^{k}|_{i}$ denotes the ii–th coefficient of the polynomial $a(x)^{k}$, for $n$ sufficiently large.
+Proof
+Without loss of generality, we prove the statement for $a^{k}|_{0}$, since the same argument applies to all other terms.
+Let $\omega_{1},\dots,\omega_{n}$ be the roots of $x^{n}+1$, which is a cyclotomic polynomial since $n$ is a power of two.
+Then, we know that $\sum_{i=1}^{n}\omega_{i}^{m}=0$ for $1\leq m\leq n-1$, and $a(\omega_{1})^{k}+\dots+a(\omega_{n})^{k}=na^{k}|_{0}+a^{k}|_{1}(\omega_{1}+\dots+\omega_{n})+\dots+a^{k}|_{n-1}(\omega_{1}^{n-1}+\dots+\omega_{n}^{n-1})$, thus
+$a^{k}|_{0}=\frac{1}{n}(a(\omega_{1})^{k}+\dots+a(\omega_{n})^{k}).$
+Now, we define $Z_{i}:=a(\omega_{i})=a_{0}+a_{1}\omega_{i}+\dots+a_{n-1}\omega_{i}^{n-1},$ for $i=1,\ldots,n,$ where, for $n$ sufficiently large, thanks to [20, Theorem 9] we have that $Z_{i}$ follows a Gaussian distribution centered at zero.
+Then
+$\mathbb{E}[Z_{m}\cdot Z_{l}]$ $=\mathbb{E}[(a|_{0}+a|_{1}\omega_{m}+\dots+a|_{n-1}\omega_{m}^{n-1})(a|_{0}+a|_{1}\omega_{l}+\dots+a|_{n-1}\omega_{l}^{n-1})]=$
+$=\mathbb{E}[a|_{0}^{2}]+\mathbb{E}[a|_{1}^{2}]\omega_{m}\omega_{l}+\dots+\mathbb{E}[a|_{n-1}^{2}]\omega_{m}^{n-1}\omega_{l}^{n-1}=$
+$=V_{a}(1+\omega_{m}\omega_{l}+\dots+\omega_{m}^{n-1}\omega_{l}^{n-1})=$
+$=\begin{cases}nV_{a}\quad\quad\text{if}\quad\omega_{m}=\omega_{l}^{-1}\\ 0\quad\quad\text{otherwise}.\end{cases}$
+Therefore, for $m=l$, $\mathbb{E}[Z_{m}^{2}]=0$ and, since every linear combination of $Z_{i}$’s is normally distributed, we can apply the Isserlis theorem (Theorem 4.1) to obtain $\mathbb{E}[Z_{m}^{k}]=0$.
+Consequently
+$\mathbb{E}[a^{k}|_{0}]=\frac{1}{n}(\mathbb{E}[Z_{1}^{k}]+\dots+\mathbb{E}[Z_{n}^{k}])=0.$
+Moreover, $\mathbb{E}[a^{k}|_{0}^{2}]=\frac{1}{n^{2}}\sum_{m}\mathbb{E}[Z_{m}^{k}Z_{l}^{k}]=\frac{1}{n^{2}}\sum_{m,l}\mathbb{E}[Z_{m}^{k}Z_{m^{-1}}^{k}]=k!n^{k-1}V_{a}^{k}$, where the last equality follows again from the Isserlis theorem.
+Indeed,
+$\mathbb{E}[Z_{m}^{k}Z_{m^{-1}}^{k}]$ $=\mathbb{E}[Z_{m}\cdot\dots\cdot Z_{m}\cdot Z_{m^{-1}}\cdot\dots\cdot Z_{m^{-1}}]=$
+$=\mathbb{E}[X_{1}\cdot\dots\cdot X_{m}\cdot X_{m+1}\cdot\dots\cdot X_{2m}]=$
+$=\sum\prod\mathbb{E}[X_{i}X_{j}]=$
+$=\begin{cases}0\quad\quad\text{if}\,\,\,0\leq i,j\leq k\,\,\,\text{or}\,\,\,k+1\leq i,j\leq 2k+1\\ nV_{a}\quad\quad\text{otherwise}.\end{cases}$
+Therefore, $\mathbb{E}[Z_{m}^{k}Z_{m^{-1}}^{k}]$ is given by the sum of $k!$ addends whose value is $n^{k}V_{a}^{k}$, i.e., $\mathbb{E}[Z_{m}^{k}Z_{m^{-1}}^{k}]=k!n^{k}V_{a}^{k}$.
+Finally, $\mathbb{E}[a^{k}|_{0}^{2}]=\frac{1}{n^{2}}\cdot n\cdot k!\cdot n^{k}\cdot V_{a}^{k}=k!n^{k-1}V_{a}^{k}$.
+∎
+Remark 1
+This result extends some of [24, Theorem 4.3], where the authors proved the same statement, using a different approach, but only for the case where the $a|_{i}$’s are normal independent random variables.
+In our case, the coefficients of the polynomial $a(x)$ follow any distribution (with zero mean and finite variance).
+Moreover, a proof of the specific case $k=2$ is also provided in [5, Lemma 2].
 Definition 4
-Let $g_{x}(\iota)=\prod_{i=0}^{\iota}f_{x}(i)$ with $f_{x}$ defined as in (4) and $f_{x}(0)=f_{x}(1)=~{}1$. The correction function $F_{x}$ is
-$F_{x}(\iota_{1},\iota_{2}):=\frac{g_{x}(\iota_{1}+\iota_{2})}{g_{x}(\iota_{1})\cdot g_{x}(\iota_{2})}$
-Moreover, this function satisfies
-$F_{x}(\iota_{1},\iota_{2})\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[x^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[x^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[x^{\iota_{2}}|_{i_{2}}^{2}]}$
-A proof of this result can be found in [2].
-For the purposes of our study, we denote by $F_{e}$ the correction function required to account for the dependencies induced by the term $e$, and by $F_{s}$ the correction function associated with the term $s$.
+Let $a\in\mathcal{R}$. The correction function $F_{a}$ is defined as
+$F_{a}(\iota_{1},\iota_{2})=\frac{\sum_{i=0}^{n-1}\mathbb{E}[a^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[a^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[a^{\iota_{2}}|_{i_{2}}^{2}]}\,.$ (8)
+From Lemma 3, it follows that for polynomials with coefficients following distributions of that form, it holds that
+$F_{a}(\iota_{1},\iota_{2})=\frac{(\iota_{1}+\iota_{2})!}{\iota_{1}!\,\iota_{2}!}\,.$ (9)
+We observe that the value of $F_{a}$ is the same for all polynomials as in Lemma 3.
+However, for the sake of clarity in presenting our results, we will keep separate notations, writing $F_{s}$ for the secret key $s$ and $F_{e}$ for the public key $e$.
 We can now introduce our main theorem.
@@
 $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})\leq n\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}})|_{i})\mathsf{Var}((a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})F_{s}(\iota_{1},\iota_{2})F_{e}(K_{1},K_{2}),$
-where $K_{1},K_{2}$ arise from the expansions
-$a_{\iota_{1}}=\sum_{\mu_{1}=0}^{K_{1}}b_{\mu_{1}}(\iota_{1})e^{\mu_{1}},\hskip 7.11317pta^{\prime}_{\iota_{2}}=\sum_{\mu_{2}=0}^{K_{2}}b^{\prime}_{\mu_{2}}(\iota_{2})e^{\mu_{2}}$
-and represent the highest power of $e$ appearing in $a_{\iota_{1}},a^{\prime}_{\iota_{2}}$, respectively.
+where $K_{1},K_{2}$ represent the highest power of $e$ appearing in $a_{\iota_{1}},a^{\prime}_{\iota_{2}}$, respectively.
 Proof
@@
 Moreover, it can be noted that $n\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})\mathsf{Var}(a^{\prime}_{\iota_{2}}s^{\iota_{2}}|_{i})$ can be written as
-$n\left(\sum_{\mu_{1}=0}^{K_{1}}\mathsf{Var}(b_{\mu_{1}}(\iota_{1})|_{i})\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{2}}]\right)\cdot$
-$\cdot\left(\sum_{\mu_{2}=0}^{K_{2}}\mathsf{Var}(b^{\prime}_{\mu_{2}}(\iota_{2})|_{i})\sum_{j_{3}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{3}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}]\right),$
-which can be reordered as
 $n\sum_{\mu}\sum_{\mu_{1}+\mu_{2}=\mu}\mathsf{Var}(b_{\mu_{1}}(\iota_{1})|_{i})\mathsf{Var}(b^{\prime}_{\mu_{2}}(\iota_{2})|_{i})$
 $\hskip 5.69046pt\cdot\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{3}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{3}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{2}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}],$
-Recalling that the correction function $F$ approximates the following quantities
-$\begin{cases}F_{e}(\iota_{1},\iota_{2})\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[e^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[e^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[e^{\iota_{2}}|_{i_{2}}^{2}]}\\ F_{s}(\iota_{1},\iota_{2})\approx\frac{\sum_{i=0}^{n-1}\mathbb{E}[s^{\iota_{1}+\iota_{2}}|_{i}^{2}]}{\sum_{i_{1}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|_{i_{1}}^{2}]\sum_{i_{2}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|_{i_{2}}^{2}]}\par\par\end{cases}$
-It is possible to use these functions in order to express $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})$ in terms of $\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})\mathsf{Var}(a^{\prime}_{\iota_{2}}s^{\iota_{2}}|_{i})$.
+By (8), we express $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})$ in terms of $\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})\mathsf{Var}(a^{\prime}_{\iota_{2}}s^{\iota_{2}}|_{i})$.
 In fact, $\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})$ can be written as
@@
 $\cdot$ $\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{2}}]\sum_{j_{3}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{3}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}]$
-[…]
+$\approx\quad$ $nF_{s}(\iota_{1},\iota_{2})\sum_{\mu}\sum_{\mu_{1}+\mu_{2}=\mu}\mathsf{Var}\left(b_{\mu_{1}}(\iota_{1})|_{i}\right)\mathsf{Var}\left(b^{\prime}_{\mu_{2}}(\iota_{2})|_{i}\right)F_{e}(\mu_{1},\mu_{2})\cdot$
 $\cdot$ $\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{2}}]\sum_{j_{3}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{3}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}].$
-Then, by exploiting the monotonicity of $F_{e}$ and $F_{s}$ [2], it is possible to derive an upper bound given by
+Then, by exploiting the monotonicity of $F_{e}$ and $F_{s}$, it is possible to derive an upper bound given by
 $nF_{s}(\iota_{1},\iota_{2})F_{e}(K_{1},K_{2})$ $\sum_{\mu}\sum_{\mu_{1}+\mu_{2}=\mu}\mathsf{Var}\left(b_{\mu_{1}}(\iota_{1})|_{i}\right)\mathsf{Var}\left(b^{\prime}_{\mu_{2}}(\iota_{2})|_{i}\right)\cdot$
-[…]
+$\quad\cdot$ $\sum_{j_{1}=0}^{n-1}\mathbb{E}[e^{\mu_{1}}|^{2}_{j_{1}}]\sum_{j_{2}=0}^{n-1}\mathbb{E}[e^{\mu_{2}}|^{2}_{j_{2}}]\sum_{j_{3}=0}^{n-1}\mathbb{E}[s^{\iota_{1}}|^{2}_{j_{3}}]\sum_{j_{4}=0}^{n-1}\mathbb{E}[s^{\iota_{2}}|^{2}_{j_{4}}],$
 where $K_{1},K_{2}$ represent the highest power of $e$ appearing in $a_{\iota_{1}},a^{\prime}_{\iota_{2}}$, respectively.
 It is straightforward to verify that this concludes our proof, yielding
-$\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})=nF_{s}(\iota_{1},\iota_{2})F_{e}(K_{1},K_{2})\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})\mathsf{Var}(a^{\prime}_{\iota_{2}}s^{\iota_{2}}|_{i}).$
-Remark 1
-Theorem 4.1 provides an upper bound on the variance of the product $a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}}$.
-However, it is significant to point out that this estimate can be regarded as a faithful approximation, i.e.,
-$\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})\approx n\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}})|_{i})\mathsf{Var}((a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})F_{s}(\iota_{1},\iota_{2})F_{e}(K_{1},K_{2}).$
-The reason behind this statement can be immediately found in the distributions of the $b_{\mu}$-terms in the noise of a fresh ciphertext.
-In fact, for a fresh ciphertext, the critical quantity can be expressed as
-$\nu=a_{0}+a_{1}s=(b_{0}(0)+b_{1}(0)\cdot e)+b_{0}(1)s,$
-where
-$\begin{cases}a_{0}=b_{0}(0)+b_{1}(0)\cdot e=(m+te_{0})+tu\cdot e\\ a_{1}=b_{0}(1)=te_{1}\end{cases}$
-Then, it can be observed that the variance of the $b_{\mu}$-values is given by
-$\begin{cases}\mathsf{Var}(b_{0}(0)|_{i})=V_{m}+t^{2}V_{e}\\ \mathsf{Var}(b_{1}(0)|_{i})=nt^{2}V_{e}V_{u}\\ \mathsf{Var}(b_{0}(1)|_{i})=nt^{2}V_{e}V_{s}\end{cases}$
-where $n$ typically assumes values in $\{2^{13},2^{14},2^{15},2^{16}\}$. Consequently,
-$\mathsf{Var}(b_{0}(0)|_{i})\ll\mathsf{Var}(b_{1}(0)|_{i}),\mathsf{Var}(b_{0}(1)|_{i}).$
-Therefore, $b_{0}(0)$ becomes negligible in our analysis, and we can assume $a_{0}\approx b_{1}(0)\cdot e$, obtaining the approximated version of the theorem.
-However, in order to maintain our results as general as possible, our next analysis will refer to the upper bound version of the theorem, still knowing that it corresponds to a reliable approximation.
-5 Application of Our Variance Estimation Method to Fixed Circuits
-This section employs the results of Section 4 to propose an effective method for tracking the error growth in circuits with a fixed number of operation.
-As previously highlighted, obtaining bounds sufficiently tight with respect to the experimental values of the error coefficients’ variance is crucial for development and real use of homomorphic encryption.
-These bounds, in fact, allow one to construct new techniques for efficiently finding initial scheme parameters that improve performance and security simultaneously.
-Naturally, the behavior of the error is very dependent on the specific circuit under consideration, that is, the sequence of operations that the scheme must accommodate.
-We analyze a circuit in which pairs of ciphertexts are progressively multiplied.
-We focus on circuits that primarily involves homomorphic multiplication, since, as previously mentioned, it is the most complex and therefore the most significant operation to study.
-For clarity, we divide the analysis into two distinct scenarios.
-In the first scenario, only theoretical, the modulus switch technique is not applied, resulting in a very rapid increase of the error.
-Instead, in the second case, the modulus switch is performed after each multiplication.
-Typically, the first scenario is not taken into account, as the error magnitude, without modulus switching, quickly becomes unmanageable, making the scheme impractical for real-world applications.
-In contrast, with modulus switching, error growth is significantly diminished, and the noise remains within a manageable range.
-This is, in fact, one of the primary goals pursued in designing practical FHE schemes.
-However, the application of modulus switch drastically reduces the differences in estimates derived from even significantly distant approaches.
-Instead, by studying circuits without the modulus switch, we believe that the necessity of considering the dependencies in the coefficients of the error polynomial becomes unequivocal.
-For this reason, although the absence of modulus switching does not reflect a realistic scenario, we decided to start our analysis in this simplified context.
-This allows us to highlight the key structural aspects of error growth, which are then carried over and extended to the more realistic case involving modulus switching.
-Again, we want to stress that circuits without modulo switch are not practicable and do not have real world applications.
-The choice to describe this scenario is purely to complete the explanation of our method, further emphasizing the differences between our approach and others considered.
-The remainder of this section is structured as follows.
-In the first part, we present circuits without modulus switching.
-Subsequently, we focus on practical circuits that employ modulus switching, showing how the results from Section 4 can be applied in this setting.
-Based on the results, we are confident that our estimates represent a tangible improvement over the current state of the art.
-Our bounds closely match the experimental results, while consistently avoiding underestimation, which affects instead the existing approaches and must be avoided for the proper functioning of the scheme together with its security.
-5.1 Circuits without Modulo Switch
+$\mathsf{Var}((a_{\iota_{1}}s^{\iota_{1}}\cdot a_{\iota_{2}}^{\prime}s^{\iota_{2}})|_{i})^{\prime}\leq n\mathsf{Var}(a_{\iota_{1}}s^{\iota_{1}}|_{i})\mathsf{Var}(a^{\prime}_{\iota_{2}}s^{\iota_{2}}|_{i})F_{s}(\iota_{1},\iota_{2})F_{e}(K_{1},K_{2}).$
+5 The Shape of Noise: Gaussian Distributions in BGV
+In order to provide accurate bounds for the selection of initial parameters, it is necessary to first discuss the distribution of the error coefficients.
+Our analysis relies on the assumption that these coefficients follow a normal distribution.
+Under this assumption, tight average-case bounds can be derived, facilitating efficient and reliable parameter selection.
+However, assuming a Gaussian distribution when this does not hold may result in a significant underestimation of the error bounds, causing computation failures, as highlighted in [24].
+It is worth noting that [24] claims that variance-based methods offer no theoretical guarantee on the failure probability.
+This statement can be misleading, as variance-based analyses remain valid once an appropriate bound on the actual distribution is established.
+Furthermore, we emphasize that the results proven in [24] concern the heavy-tailed nature of the error coefficient distributions in the absence of modulus switching, which is not practical, since BGV is always used with modulus switching technique.
+As noted in the article itself, applying modulus switching ensures that the dominant terms are effectively distributed according to Gaussian distributions.
+Therefore, assuming Gaussianity remains reasonable when analyzing BGV parameters, in line with other works such as [18, 35].
+Indeed, as the authors of [24] themselves note, no evidence of decryption failures has been observed in practical BGV usage scenarios, nor in those of CKKS [5].
+This section aims to rigorously justify why the Gaussian assumption for the error coefficients is reasonable and to identify the conditions on the choice of initial parameters that are required to ensure it.
+In particular, we will show that the values of the primes $p_{\ell}$ must satisfy a certain lower bound.
+This requirement is reasonable: the primes used in the most common libraries already meet this criterion.
+Moreover, while imposing a minimum size on these primes, our result still allows for improvements and a reduction in the overall modulus size.
+5.1 Gaussian Distribution
+In this section, we provide both theoretical arguments and empirical evidence that the use of modulus switching ensures that the error coefficients are, in practice, Gaussian distributed.
+We show that the coefficients of the critical quantity can be reasonably assumed to follow a Gaussian distribution at three distinct stages: immediately after the initial encryption, after modulus switching, and after the multiplication of two ciphertexts that have both undergone modulus switching beforehand.
+Specifically, we provide a formal proof for the first two cases, while for the latter we verify this behavior experimentally.
+Gaussianity of the Fresh Error.
+The critical quantity associated with a fresh ciphertext is given by
+$m+t(e\cdot u+e_{1}\cdot s+e_{0}),$
+where $e,e_{0},e_{1}\sim\mathcal{DG}_{q}(\sigma^{2})$ and $m,s,u\sim\mathcal{U}_{3}$. We are thus interested in the distribution of the coefficients
+$m|_{i}+t(\sum_{j=0}^{n-1}e|_{j}u|_{i-j}+\sum_{j=0}^{n-1}e_{1}|_{j}s|_{i-j}+e_{0}|_{i}),$
+where $m|_{i}\longleftarrow\mathcal{U}_{t}$, $te_{0}|_{i}\longleftarrow\mathcal{DG}_{q}(t^{2}\sigma^{2})$ and $t\sum_{j=0}^{n-1}e|_{j}u|_{i-j},t\sum_{j=0}^{n-1}e_{1}|_{j}s|_{i-j}\longleftarrow\mathcal{DG}_{q}(t^{2}nV_{e}V_{s})$.
+Since the sum of independent Gaussian random variables is still Gaussian with variance equal to the sum of individual variances [23], the error term can be written as
+$m|_{i}+K|_{i},$
+where $m|_{i}\longleftarrow\mathcal{U}_{t}$ and $K|_{i}\longleftarrow\mathcal{DG}_{q}(t^{2}V_{e}(2nV_{s}+1))$.
+Therefore, the Gaussianity of the final critical quantity follows from the fact that the first term is negligible compared to the second.
+In Figure 1, we present the results for circuits of multiplicative depth 3 (Figure 1(a)) and 6 (Figure 1(b)), focusing on the distribution of the first coefficient of the fresh error.
+We used OpenFHE library [1] to generate 50,000 error samples, and we analyzed their coefficients using the Python fitter package [^2] https://fitter.readthedocs.io/en/latest/.
+The code used to generate these samples and to perform the estimates in the following sections is publicly available [^3] https://github.com/nadirmur/openFHE.
+As the figures show, the distribution of the fresh error coefficient is well approximated by a Gaussian.
+In particular, the Kolmogorov–Smirnov test produces a $p$-value >0.05>0.05 (kspval in the caption) [37], the Anderson-Darling test statistic (adstat) is below the critical value (adcrit) [2, 38] at the $15\%$ significance level, and the kurtosis is approximately $3$.
+The parameters used are: $t=65537$, $n=2^{13}$, ciphertext modulus $q$ chosen by the library to ensure at least 128-bit security, $\chi_{s}=\chi_{u}=\mathcal{U}_{3}$, and $\chi_{e}=\mathcal{DG}(\sigma^{2})$ with $\sigma=3.19$.
+We employed hybrid key switching and the HPSPOVERQ multiplication method [1].
+(a) adstat 0.517 << adcrit 0.576 ($\alpha=0.15$), kspval 0.896 and kurtosis 2.999. Circuit of multiplicative depth 3.
+(b) adstat 0.304 << adcrit 0.576 ($\alpha=0.15$), kspval 0.940 and kurtosis 3.00. Circuit of multiplicative depth 6.
+Figure 1: Distribution of the first coefficient of the fresh error
+Gaussianity after Modulus Switching.
+Given $\mathfrak{c}=(\mathbf{c},q_{\ell},\nu)$ an extended ciphertext, we recall that the variance of the error coefficients after modulo switching to the target modulo $q^{\prime}_{\ell}$, where $q_{\ell}=p_{\ell}q^{\prime}_{\ell}$, is
+$V(\nu_{\mathsf{ms}}|_{i})=\tfrac{1}{p_{\ell}^{2}}(\mathsf{Var}(\nu|_{i})+V_{\delta_{0}}+nV_{\delta_{1}}V_{s})\approx\tfrac{1}{p_{\ell}^{2}}(\mathsf{Var}(\nu|_{i})+nV_{\delta_{1}}V_{s}).$ (10)
+as $\mathsf{Var}({\delta_{0}}/{p_{\ell}}|_{i})=t^{2}/12\ll\mathsf{Var}({\delta_{1}}s/{p_{\ell}}|_{i})=nt^{2}V_{s}/12$.
+In our analysis, we will assume that the dominant component in (10) is the second term.
+This assumption is quite common in the study of BGV, as can be seen in [18].
+We will later clarify how our choice of moduli is specifically designed to ensure that this condition is satisfied.
+It is precisely this constraint that allows us to conclude the Gaussianity of the error coefficients after modulus switching.
+By making the first term negligible, the error after modulus switching reduces, according to the previous notation, to ${\delta_{1}}/{p_{\ell}}s,$ with ${\delta_{1}}/{p_{\ell}}\longleftarrow\mathcal{U}_{t}$.
+Thus, we obtain
+$\frac{\delta_{1}}{p_{\ell}}s|_{i}=\sum_{j=0}^{n-1}\xi(i,j)\hskip 2.84544pt\frac{\delta_{1}}{p_{\ell}}|_{j}\hskip 2.84544pts|_{i-j}.$
+Therefore, we can observe that each coefficient is given by the sum of i.i.d.
+random variables with finite mean and variance.
+Since $n$ is sufficiently large, by applying the Central Limit Theorem, we can conclude that
+$\frac{\delta_{1}}{p_{\ell}}s\longleftarrow\mathcal{DG}_{q}(n\frac{t^{2}}{12}V_{s}).$
+Gaussianity after a multiplication.
+In Figure 2, we present the results for circuits of multiplicative depth 3 (Figure 2(a)) and 4 (Figure 2(b)), focusing on the distribution of the first coefficient of the error just after the last multiplication.
+Moreover, Figures 2(c) and 2(d) show the multiplication error after the third and fifth multiplications, respectively, for a circuit of depth 6.
+We used OpenFHE library [1] (with the same setting as before) to generate 50,000 error samples and analyzed their coefficients using the Python fitter package.
+As the figures show, the distribution of the error coefficients is well approximated by a Gaussian.
+(a) adstat $0.267<$ adcrit $0.576$ ($\alpha=0.15$), kspval $0.867$ and kurtosis $2.987$. After the third multiplication.
+(b) adstat $0.306<$ adcrit $0.576$ ($\alpha=0.15$), kspval $0.937$ and kurtosis $3.03$. After the fourth multiplication.
+(c) adstat 0.233 << adcrit $0.576$ ($\alpha=0.15$), kspval $0.985$ and kurtosis $2.980$. After the 3rd multiplication in a circuit of multiplicative depth 6.
+(d) adstat 0.215 << adcrit $0.576$ ($\alpha=0.15$), kspval $0.842$ and kurtosis $2.985$. After the 5th multiplication in a circuit of multiplicative depth 6.
+Figure 2: Distribution of the first coefficient just after a multiplication.
+The fact that the noise coefficients follow a Gaussian distribution is particularly advantageous, as it allows us to bound the maximum absolute value of the noise coefficients with high probability simply by controlling their variance $V$.
+This implies that, to satisfy the correctness condition, a noise vector $\nu$ must satisfy $\|\nu\|_{\infty}<{q}/{2}.$ We can use Equation 2 to bound the failure probability
+$\mathbb{P}\left(\|\nu\|_{\infty}>\frac{q}{2}\right)\leq n\left(1-\text{erf}\left(\frac{q}{2\sqrt{2V}}\right)\right),$
+where $V$ denotes the estimated variance of the noise coefficients.
+To express this bound more conveniently, we introduce a security parameter $D$ such that $D\leq{q}/{2\sqrt{2V}}$, from which it follows, using the monotonicity of the error function, that
+$\mathbb{P}\left(\|\nu\|_{\infty}>\frac{q}{2}\right)\leq n\left(1-\text{erf}(D)\right).$
+Thus, by appropriately choosing the security parameter $D$, we can ensure that the probability of decryption failure remains negligibly small.
+For instance, setting $D=6$ and $n=2^{13}$ yields a failure probability of approximately $2^{-42}$.
+For practical applications, $D=8$ should be preferred, resulting in a probability of approximately $2^{-83}$, when the ring dimension is $n=2^{13}$.
+Consequently, the ciphertext modulus $q$ can be selected as
+$q\geq 2D\sqrt{2V}.$ (11)
+It is worth noting that the bounds derived in our analysis provide insight into the minimum ciphertext modulus $q$ required to guarantee the correctness of the scheme, a key factor for optimizing performance and efficiency.
+5.2 On the Role of Modulus Switching in Preserving Gaussianity
+In this section, we highlight the crucial role of modulus switching in the studied circuits.
+This mechanism is essential not only to control the noise growth induced by homomorphic operations, but also to shape the error distribution.
+Our experiments confirm this behavior: when modulus switching is omitted, the Gaussianity of the coefficients could no longer hold, making it necessary to explicitly characterize their distribution.
+Specifically, Figure 3 shows the error distribution after three multiplications without modulus switching for $n=2^{13}$, where the tails clearly deviate from Gaussian behavior.
+(a) asymmetry 0.5240-0.5240 and kurtosis $14.3769$
+(b) when the distribution is Gaussian, the points lie along the diagonal.
+Figure 3: Error distribution after 3 multiplications without modulus switching.
+It is worth noting that this deviation can be observed even without plotting the distribution.
+Indeed, when the Gaussian bound introduced in the previous section is applied to select the ciphertext modulus $q$ according to Equation 11, the resulting parameters quickly lead to an unexpectedly high failure rate.
+For instance, when targeting a decryption failure probability of approximately $2^{-83}$ (as obtained by setting $D=8$ in our bound), experiments without modulus switching exhibit failure rates of around 3%.
+This discrepancy provides strong evidence that the Gaussian assumption no longer holds and that the tails of the distribution are significantly heavier, in agreement with [24].
+The validity of our approach is therefore directly linked to the use of modulus switching.
+As previously shown, the error coefficients after encryption, modulus switching, and multiplication preserve their Gaussian distribution.
+For this property to hold, as highlighted in the previous sections, it is necessary to ensure that the dominant terms after modulus switching and after multiplication are, respectively, $\delta_{1}s/p_{\ell}$ and $(\delta_{1}/p_{\ell}\cdot\delta_{1}^{\prime}/p_{\ell})s^{2}$.
+For this reason, in our moduli selection, we explicitly require that
+$\frac{V_{\ell-1}}{p_{L-\ell}^{2}}<\alpha V_{\mathsf{ms}}.$ (12)
+for every level $\ell$, with $\alpha=\nicefrac{{1}}{{100}}$.
+This condition imposes a lower bound on the moduli $p_{\ell}$, which must be selected accordingly.
+Nevertheless, it does not compromise the modulus-size reduction achieved by our approach compared to the state-of-the-art.
+In fact, the primes generated by most existing libraries already satisfy this requirement.
+6 Variance Estimation in Circuits with Fixed Depth
+This section builds on the results of Section 4 and 5 to propose a method for tracking the error growth in circuits with a fixed number of operations.
+As previously noted, obtaining sufficiently tight bounds with respect to the experimental variance of the error coefficients is crucial for identifying initial scheme parameters that simultaneously improve both performance and security.
+We analyze the circuit depicted in Figure 4 in which pairs of ciphertexts are progressively multiplied.
+Note that we focus on circuits that involve homomorphic multiplication since it is the most complex and, therefore, the most significant operation to study.
+It is worth noting that the circuit described in this work closely resembles the one adopted by default in OpenFHE [1], where multiplications are always preceded by a modulus switching operation.
 Figure 4: Reference circuit
-Figure 1 provides a schematic representation of the circuit used as a reference for our analysis.
-Let $M$ denote the multiplicative depth of the circuit, that is, the number of multiplications to be supported.
-To facilitate the analysis, we divide the circuit into $L=M+1$ levels.
-For the considered circuit, we assume the independent and random generation of $2^{L-1}$ plaintexts, which are then encrypted using the same public key, resulting in $2^{L-1}$ fresh ciphertexts, denoted by $\bm{c}_{1},\dots,\bm{c}_{2^{L-1}}$.
-At each level $l\in[0,\dots,L-1]$, pairs of ciphertexts are multiplied, yielding the corresponding ciphertexts of the $(l+1)$-th level, and then the process is repeated until the final ciphertext is computed.
-Our study aims to estimate the variance of the coefficients of the critical quantity $\nu^{(l)}|_{i}$, denoted as $V_{l}$, which corresponds to the ciphertexts of the circuit at level $l$.
-According to the current average-case approach, the variance at level $l$ can be approximated as
-$V_{l}\leq nV_{l-1}^{2}.$
-However, as pointed out by the authors of [25], this estimation tends to underestimate the variance observed in experimental evaluations.
-Such underestimation may lead to security vulnerabilities, as previously discussed, and, therefore, must be avoided.
-In contrast, in our investigation, we proved that
-$V_{l}\leq nh(l)V_{l-1}^{2},$ (5)
-where $h$ is a function specifically designed to account for the dependencies among the coefficients of the error polynomial.
-This construction is based on the correction function $F$ introduced in [2] and represents the crucial difference compared to the current approach [25].
-Section 4 has laid the groundwork for deriving this bound.
-Thus, the construction of our bound can now be presented, beginning with the initial levels and subsequently extending the reasoning to the general case for a level $l$.
-First Multiplication
-Let $\bm{c},\bm{c^{\prime}}$ be two fresh ciphertexts with critical quantities $\nu=a_{0}+a_{1}s$, $\nu^{\prime}=a^{\prime}_{0}+a^{\prime}_{1}s$, respectively, where
-$\begin{cases}a_{0}=b_{0}(0)+b_{1}(0)\cdot e=(m+te_{0})+tu\cdot e\\ a_{1}=b_{0}(1)=te_{1}\end{cases}$
-$\begin{cases}a^{\prime}_{0}=b^{\prime}_{0}(0)+b^{\prime}_{1}(0)\cdot e=(m^{\prime}+te^{\prime}_{0})+tu^{\prime}\cdot e\\ a^{\prime}_{1}=b^{\prime}_{0}(1)=te^{\prime}_{1}\end{cases}$
-Let $V_{0}$ be the variance of the critical quantity associated with the fresh ciphertexts. Then, it can be observed that
-$\begin{cases}V(a_{0}|_{i})=V(a^{\prime}_{0}|_{i})=t^{2}(\frac{1}{12}+V_{e}+nV_{e}V_{u})\\ V(a_{1}s|_{i})=V(a^{\prime}_{1}s|_{i})=t^{2}nV_{e}V_{s}\end{cases}$
-Therefore, it is possible to assume that
-$V(a_{0}|_{i})\approx V(a_{1}s|_{i})\approx\frac{V_{0}}{2},$ (6)
-since $V_{u}=V_{s}$, due to the choice of the distributions of $s$ and $u$, and
-$\frac{1}{12}+V_{e}\ll nV_{e}V_{u},$
-which is then negligible, as highlighted in 1.
-Clearly, the same reasoning can be applied to $V(a^{\prime}_{0}|_{i}),V(a^{\prime}_{1}s|_{i})$, obtaining
-$V(a^{\prime}_{0}|_{i})\approx V(a^{\prime}_{1}s|_{i})\approx\frac{V_{0}}{2}.$ (7)
-Our goal is to estimate the coefficients variance of $\nu_{\mathsf{mul}}=\nu\cdot\nu^{\prime}$, which will be referred to as $V_{1}$, following the level notation introduced at the beginning.
-Moreover, according to the previously introduced notation, it is possible to write $\nu_{\mathsf{mul}}$ as
-$\nu_{\mathsf{mul}}=a_{0}a^{\prime}_{0}+(a_{0}a^{\prime}_{1}+a_{1}a^{\prime}_{0})\cdot s+a_{1}a^{\prime}_{1}\cdot s^{2}.$
-Consequently, when the variance is considered, applying Theorem 4.1 and using that the covariance of this addends is zero, it follows that
-$V_{1}:=\mathsf{Var}(\nu_{\mathsf{mul}}|_{i})$ $\leq n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i}){F_{s}(0,0)}{F_{e}(1,1)}$
-$+n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}((a^{\prime}_{1}s)|_{i}){F_{s}(0,1)}{F_{e}(1,0)}$
-$+n\mathsf{Var}((a_{1}s)|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i}){F_{s}(1,0)}{F_{e}(0,1)}$
-$+n\mathsf{Var}((a_{1}s)|_{i}))\mathsf{Var}((a^{\prime}_{1}s)|_{i}){F_{s}(1,1)}{F_{e}(0,0)}.$
-Moreover, this bound can be simplified, observing that $F_{s}(0,0)=F_{s}(1,0)=F_{s}(0,1)=1$ and $F_{e}(0,0)=F_{e}(1,0)=F_{e}(0,1)=1$. Therefore, it follows that
-$V_{1}\leq$ $\,n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i}){F_{e}(1,1)}+n\mathsf{Var}(a_{0}|_{i})\mathsf{Var}((a^{\prime}_{1}s)|_{i})$
-$+n\mathsf{Var}((a_{1}s)|_{i})\mathsf{Var}(a^{\prime}_{0}|_{i})+n\mathsf{Var}((a_{1}s)|_{i}))\mathsf{Var}((a^{\prime}_{1}s)|_{i}){F_{s}(1,1)}.$
-Again, this can be reformulated using (6) and (7) as
-$V_{1}$ $\leq\frac{nV_{0}^{2}}{4}\left(F_{e}(1,1)+{F_{s}(1,1)}+2\right).$
-Hence, the first value of the function $h$ we aim to construct, can be fixed as
-$h(1)=\frac{F_{s}(1,1)+{F_{e}(1,1)}+2}{4}.$
-Thus, for the first level, the following inequality has been proven
-$V_{1}\leq nh(1)V_{0}^{2}.$
-Second Multiplication
-In the second multiplication, the ciphertexts involved will have an associated critical quantity of the form
-$\nu=a_{0}+a_{1}\cdot s+a_{2}\cdot s^{2},$
-where, according to our study, each coefficient has variance
-$\begin{cases}\mathsf{Var}(a_{0}|_{i})\approx\frac{F_{e}(1,1)V_{1}}{F_{s}(1,1)+{F_{e}(1,1)}+2}\\ \mathsf{Var}(a_{1}s|_{i})\approx\frac{2V_{1}}{F_{s}(1,1)+{F_{e}(1,1)}+2}\\ \mathsf{Var}(a_{2}s^{2}|_{i})\approx\frac{F_{s}(1,1)V_{1}}{F_{s}(1,1)+{F_{e}(1,1)}+2}\end{cases}$
-Therefore, by repeating the same procedure depicted above, the critical quantity after the second multiplication will be of the form
-$\nu\cdot\nu^{\prime}=a_{0}a^{\prime}_{0}+(a_{0}a^{\prime}_{1}+a^{\prime}_{0}a_{1})s+(a_{1}a^{\prime}_{1}+a_{0}a^{\prime}_{2}+a^{\prime}_{0}a_{2})s^{2}+(a_{1}a^{\prime}_{2}+a^{\prime}_{1}a_{2})s^{3}+a_{2}a^{\prime}_{2}s^{4}$
-Thus, its coefficients’ variance can be estimated as
-$V_{2}$ $\leq\frac{nV_{1}^{2}}{(2+F_{s}(1,1)+{F_{e}(1,1)})^{2}}\bigl{[}F_{e}(1,1)^{2}\cdot{F_{s}(0,0)}{F_{e}(2,2)}$
-$\quad+4F_{e}(1,1)\cdot{F_{s}(0,1)}{F_{e}(1,2)}+4\cdot{F_{s}(1,1)}{F_{e}(1,1)}$
-$\quad+2F_{s}(1,1)F_{e}(1,1)\cdot{F_{s}(0,2)}{F_{e}(0,2)}+4F_{s}(1,1)\cdot{F_{s}(1,2)}{F_{e}(0,1)}$
-$\quad+F_{s}(1,1)^{2}\cdot{F_{s}(2,2)}{F_{e}(0,0)}\bigr{]},$
-where this inequality follows from the fact that $a_{0}$ contains a term multiplied by $e^{2}$, $a_{1}$ contains a term multiplied by $e$ and $a_{2}$ does not contain any powers of $e$.
-As a result, $h(2)$ can be chosen as
-$h(2)$ $=\frac{1}{(2+F_{s}(1,1)+{F_{e}(1,1)})^{2}}\bigl{[}F_{e}(1,1)^{2}\cdot{F_{s}(0,0)}{F_{e}(2,2)}$
-$\quad+4F_{e}(1,1)\cdot{F_{s}(0,1)}{F_{e}(1,2)}+4\cdot{F_{s}(1,1)}{F_{e}(1,1)}$
-$\quad+2F_{s}(1,1)F_{e}(1,1)\cdot{F_{s}(0,2)}{F_{e}(0,2)}$
-$\quad+4F_{s}(1,1)\cdot{F_{s}(1,2)}{F_{e}(0,1)}+F_{s}(1,1)^{2}\cdot{F_{s}(2,2)}{F_{e}(0,0)}\bigr{]}.$
-Consequently, $V_{2}\leq nh(2)V_{1}^{2}$.
-Generalization
-Finally, this procedure can be easily generalized in order to compute all the necessary values for $h$, therefore proving our bound in (5), i.e.,
-$V_{l}\leq nh(l)V_{l-1}^{2},\hskip 28.45274ptl=1,\dots,L-1.$
-5.2 Circuits with Modulo Switch
-Figure 2: Reference circuit
-In the present section, the reasoning previously introduced is adapted to circuits where the modulus switching technique is employed, in order to exploit the advantages of the method in practical scenarios.
-Let $L$ denote the number of levels in the circuit, and $\bm{c}_{1},\dots,\bm{c}_{2^{L-1}}$ the initial fresh ciphertexts, generated by encrypting $2^{L-1}$ independent and randomly generated messages, using the same key.
-The reference circuit for our analysis is depicted in Fig.
-2.
-We consider a model in which, at each level, homomorphic multiplication of pairs of ciphertexts is carried out.
-The key difference is that at the beginning of each level $l\in\{1,\dots,L-1\}$ the input ciphertexts are switched to a smaller modulus, to maintain the error almost constant throughout the process.
-In particular, following the previously defined notion, let $q_{l}$ be the ciphertexts moduli of each level, defined as
-$q_{l}=\prod_{j=0}^{l}p_{j},$
-where $p_{j}$ are primes such that $\gcd(p_{i},p_{j})=1$ for $i\neq j$.
-The process represented by the circuit can be resumed as follows:
-  • At level 0, all the fresh ciphertexts are defined in $\mathcal{R}_{q_{L-1}}\times\mathcal{R}_{q_{L-1}}$;
+Let $L$ denote the number of levels in the circuit, and $\bm{c}_{1},\dots,\bm{c}_{2^{L-1}}$ the initial fresh ciphertexts, generated by encrypting $2^{L-1}$ independent and randomly generated messages using the same key.
+We consider a model where, at each level, homomorphic multiplication of pairs of ciphertexts is carried out.
+At the beginning of each level $\ell\in\{1,\dots,L-1\}$ the input ciphertexts are switched to a smaller modulus, to maintain the error almost constant throughout the process.
+We recall that, $q_{\ell}$’s are the ciphertext moduli of each level L1L-1-$\ell$, defined as $q_{\ell}=\prod_{j=0}^{\ell}p_{j}$, where $p_{j}$ are primes such that $\gcd(p_{i},p_{j})=1$ for $i\neq j$.
+We denote by $\nu_{\ell}^{\mathsf{ms}}$ the critical quantity obtained after performing the modulus switching at the beginning of level $\ell$, and by $\nu_{\ell}$ the critical quantity at the end of level $\ell$, namely just after the multiplication (and key switching technique).
+Similarly, $V_{\ell}^{\mathsf{ms}}$ and $V_{\ell}$ denote the variance of the coefficients of these two critical quantities, respectively.
+Finally, for each level $\ell$ we will use the notation $V_{\mathsf{ms}}$ to denote $\mathsf{Var}(\delta_{1}s/p_{\ell}|_{i})$.
+The process represented by the circuit can be summarized as follows:
+  • At level 0, all the fresh ciphertexts are defined in $\mathcal{R}_{q_{L-1}}^{2}$.
   • At level 1, modulo switch to $q_{L-2}$ is applied to each of the initial fresh ciphertexts.
 After that, the first homomorphic multiplication is performed in $\mathcal{R}_{q_{L-2}}$, yielding $2^{L-2}$ resulting ciphertexts in $\mathcal{R}_{q_{L-2}}^{3}$.
-Finally, these ciphertexts are relinearized to obtain equivalent ones in $\mathcal{R}_{q_{L-2}}^{2}$, which constitutes the input of the next level;
-  • At level $2$, each input ciphertext undergoes a modulo switch from $q_{L-2}$ to $q_{L-3}$, which is again followed by the homomorphic multiplication of pairs of them.
-To conclude, a relinearization step is carried out;
-  • This process is repeated until level $L-1$, where a final ciphertext, which constitutes the output of the circuit, is returned;
+Finally, these ciphertexts are relinearized to obtain the equivalent ones in $\mathcal{R}_{q_{L-2}}^{2}$, which are the inputs of the next level.
+  • This process is repeated until level L1L-1, where a final ciphertext, which is the output of the circuit, is returned.
 It should be noted that, in the circuit represented in Fig.
-2, the relinearization step is not explicated.
-In fact, as for the case of the circuit without modulus switching, we decided to omit its contribution from our error analysis, since it is negligible compared to the impact of modulus switch and multiplication.
-Before discussing the results, it should be noted that, in contrast to the case without the modulus switching, in this scenario, $L$ ciphertext moduli are required for the operation of the circuit.
-Therefore, the selection of parameters will consist in choosing all the $q_{l}$ values, i.e., all the primes $p_{j}$, such that the correct functioning of the scheme is ensured while still choosing sufficiently small moduli in order to improve efficiency, which is crucial for practical usage of BGV.
-For better understanding, we will divide our analysis according to the level $l$ considered.
-Specifically, an in-depth explanation of the noise estimates for the first three levels will be provided.
-Then, these results will be generalized for a fixed level $l$, thus delineating our bound.
-Level 0
-Let $\bm{c}\in\mathcal{R}_{q_{L-1}}\times\mathcal{R}_{q_{L-1}}$ be a fresh ciphertext, obtained encrypting a random message $m\in\mathcal{R}_{t}$.
-Let $\nu_{\mathsf{clean}}=a_{0}+a_{1}s$ be the critical quantity associated with $\bm{c}$.
-As for the circuits without modulo switching, it is possible to assume that
-$\mathsf{Var}(a_{0}|_{i})\approx\mathsf{Var}((a_{1}s)|_{i})\approx\frac{V_{0}}{2},$ (8)
-where the notation $V_{0}$ is used to denote the variance of the coefficients of the critical quantity $\nu_{\mathsf{clean}}$.
-In general, the notation $V_{l}$ will be used to represent the variance of the coefficients of the critical quantity at the end of level $l$.
-It is significant to observe that the assumption in (8) applies to all initial $2^{L-1}$ ciphertexts.
-In fact, the estimate of $V_{l}$ is the same for all ciphertexts belonging to the same level.
-Level 1
-As mentioned above, at the beginning of level $1$, all ciphertexts are subject to a modulus switch from $q_{L-1}$ to $q_{L-2}$.
-Therefore, the error and the modulus of $\bm{c}$ are rescaled by a factor of ${q_{L-2}}/{q_{L-1}}={1}/{p_{L-1}}$. Thus, the critical quantity associated to the resulting ciphertext is given by
-$\nu_{\mathsf{modswitch}}=\frac{\nu_{\mathsf{clean}}+\nu_{\mathsf{ms}}(p_{L-1})}{p_{L-1}},$ (9)
-where $\nu_{\mathsf{ms}}(p_{L-1})$ is defined as
-$\nu_{\mathsf{ms}}(p_{L-1})=\delta_{0}+\delta_{1}s\hskip 28.45274pt\delta_{i}=t[-c_{i}t^{-1}]_{p_{L-1}}$
-In particular, for the following analysis we will use the notation
-$\frac{\nu_{\mathsf{ms}}(p_{L-1})}{p_{L-1}}=a_{0}^{\mathsf{ms}}+a_{1}^{\mathsf{ms}}s,$
-with
-$\begin{cases}a_{0}^{\mathsf{ms}}=\frac{\delta_{0}}{p_{L-1}}\\ a_{1}^{\mathsf{ms}}=\frac{\delta_{1}}{p_{L-1}}\end{cases}$
-Moreover, we will denote with $V_{\mathsf{ms}}$ the following variance
-$V_{\mathsf{ms}}=\mathsf{Var}\left(\frac{\nu_{\mathsf{ms}}(p_{L-1})}{p_{L-1}}\middle|_{i}\right).$
-For our analysis, it is possible to assume that $V_{\mathsf{ms}}\approx\mathsf{Var}(a_{1}^{\mathsf{ms}}s|_{i})$.
-This is a consequence of the fact that, observing the distributions of the coefficients in $a_{0}^{\mathsf{ms}}$ and $a_{1}^{\mathsf{ms}}$, it is evident that $\mathsf{Var}(a_{0}^{\mathsf{ms}}|_{i})\ll\mathsf{Var}(a_{1}^{\mathsf{ms}}s|_{i})$.
-In fact, noting that $c_{i}\leftarrow\mathcal{U}_{q_{L-1}}$, it is possible to estimate the variance of $a_{0}^{\mathsf{ms}}|_{i}$ and $a_{1}^{\mathsf{ms}}|_{i}$ as
-$\begin{cases}\mathsf{Var}(a_{0}^{\mathsf{ms}}|_{i})=\mathsf{Var}(a_{1}^{\mathsf{ms}}|_{i})=t^{2}/12\\ \mathsf{Var}(a_{1}^{\mathsf{ms}}s|_{i})={t^{2}nV_{s}}/{12},\end{cases}$
-where $n$ is typically in $\small\{2^{12},2^{13},2^{14},2^{15}\}$.
-With this in mind, we can proceed with the analysis of the noise growth related to the multiplication of two ciphertexts, whose associated critical quantity is as in (9).
-We start by observing that the critical quantity in (9) can be written as
-$\nu_{\mathsf{modswitch}}=\hat{a_{0}}+\hat{a_{1}}s=\frac{a_{0}}{p_{L-1}}+a_{0}^{\mathsf{ms}}+\left(\frac{a_{1}}{p_{L-1}}+a_{1}^{\mathsf{ms}}\right)s.$
-Moreover, it is possible to further simplify the notation by using the fact that $\mathbb{E}[a_{0}^{\mathsf{ms}}]=0$ and its variance is negligible. Therefore, we rewrite this value as
-$\nu_{\mathsf{modswitch}}=\hat{a_{0}}+\hat{a_{1}}s=\frac{a_{0}}{p_{L-1}}+\left(\frac{a_{1}}{p_{L-1}}+a_{1}^{\mathsf{ms}}\right)s.$
-Then, it can be derived that
-$\begin{cases}\mathsf{Var}(\hat{a_{0}}|_{i})\approx\frac{V_{0}}{2p_{L-1}^{2}}\\ \mathsf{Var}(\hat{a_{1}}s|_{i})\approx\frac{V_{0}}{2p_{L-1}^{2}}+V_{\mathsf{ms}}\end{cases}$
-since
+4, the relinearization step is not presented.
+In fact, we decided to omit its contribution from our error analysis, since it is negligible compared to the impact of modulus switch and multiplication.
+A detailed justification for this choice is provided in Appendix 0.C.
+Level 0.
+Let $\bm{c}\in\mathcal{R}_{q_{L-1}}^{2}$ be a fresh ciphertext, obtained encrypting a random message $m\in\mathcal{R}_{t}$.
+Let $\nu_{\mathsf{clean}}=a_{0}+a_{1}s$ be the critical quantity associated with $\bm{c}$ and let $V_{0}$ be the variance of $\nu_{\mathsf{clean}}$.
+Then, we have
+$\begin{cases}V(a_{0}|_{i})=t^{2}(\frac{1}{12}+V_{e}+nV_{e}V_{u})\\ V(a_{1}s|_{i})=t^{2}nV_{e}V_{s}.\end{cases}$
+Therefore, it is possible to assume
+$\mathsf{Var}(a_{0}|_{i})\approx\mathsf{Var}((a_{1}s)|_{i})\approx\frac{V_{0}}{2}.$ (13)
+It should be noted that the assumption in (13) applies to all initial $2^{L-1}$ ciphertexts.
+In fact, the estimate of variance $V_{\ell}$ of the coefficients of the critical quantity at the end of level $\ell$, is the same for all ciphertexts belonging to the same level.
+Level $1$.
+At the beginning of this level, all ciphertexts are subject to a modulus switching from $q_{L-1}$ to $q_{L-2}$.
+Therefore, the error and the modulus of $\bm{c}$ are rescaled by a factor ${q_{L-2}}/{q_{L-1}}={1}/{p_{L-1}}$.
+Thus, the critical quantity associated with the resulting ciphertext is
+$\nu_{1}^{\mathsf{ms}}=\frac{\nu_{\mathsf{clean}}+\delta_{0}+\delta_{1}s}{p_{L-1}},\quad\delta_{i}=t[-c_{i}t^{-1}]_{p_{L-1}}$ (14)
+Since $\delta_{i}/p_{L-1}\leftarrow\mathcal{U}_{t}$, it follows that $\mathsf{Var}(\delta_{0}/p_{L-1}|_{i})=t^{2}/12$ and $\mathsf{Var}(\delta_{1}s/p_{L-1}|_{i})={t^{2}nV_{s}}/{12}$.
+In BGV, $n$ is typically larger than $2^{12}$, making the contribution of $\delta_{0}/p_{L-1}$ negligible.
+The critical quantity in (14) can be written as
+$\nu_{1}^{\mathsf{ms}}\approx\frac{a_{0}}{p_{L-1}}+\left(\frac{a_{1}}{p_{L-1}}+\frac{\delta_{1}}{p_{L-1}}\right)s,$
+By Equation 13 we have
 $\mathsf{Var}\left(\frac{a_{0}}{p_{L-1}}\middle|_{i}\right)=\mathsf{Var}\left(\frac{a_{1}}{p_{L-1}}s\middle|_{i}\right)=\frac{V_{0}}{2p_{L-1}^{2}}.$
-In order to estimate the variance of the critical quantity resulting from multiplication, we will rely on Theorem 4.1.
-In particular, in this analysis, we will try to avoid explicitly writing the terms $b_{\mu}(\iota)$ involved in the expression of a generic $a_{\iota}=\sum_{\mu=0}^{K}b_{\mu}(\iota)e^{\mu}$, to avoid the notation becoming cumbersome.
-However, it is crucial to keep in mind the highest power of $e$, namely $K$, present in the terms $a_{\iota}$ involved in the multiplications, as the correction of the function $F$ depends on this value.
-Precisely, for fresh ciphertexts at level 00, it can be seen that for $a_{0}$ the associated $K$ is $1$, while for $a_{1}$ we have $K=0$.
-Instead, for the terms $a_{0}^{\mathsf{ms}},a_{1}^{\mathsf{ms}}$ the case is quite different.
-In fact, these values can be assumed to be randomly distributed over $\mathcal{R}_{t}$, i.e.
-$a_{0}^{\mathsf{ms}},a_{1}^{\mathsf{ms}}\leftarrow\mathcal{U}_{t}$ for every level of the circuit.
-This makes the analysis slightly more complicated, since, in order to derive tight estimates, it will be necessary to distinguish the contribution of ${a_{0},a_{1}}$ from the one of $a_{0}^{\mathsf{ms}},a_{1}^{\mathsf{ms}}$.
-Now, assume that $\bm{c},\bm{c}^{\prime}\in\mathcal{R}_{q_{L-2}}\times\mathcal{R}_{q_{L-2}}$ are two ciphertexts at level 1, after modulus switching has been carried out.
+Therefore, the variance after the modulus switch can be written as
+$V_{1}^{\mathsf{ms}}=\frac{V_{0}}{p_{L-1}^{2}}+V_{\mathsf{ms}}.$
+As discussed previously, to ensure that the error follows a Gaussian distribution, we rely on Equation 12, i.e., we can assume ${V_{0}}/{p_{L-1}^{2}}<V_{\mathsf{ms}}/100$.
+In order to estimate the variance of the critical quantity resulting from multiplication, we rely on Theorem 4.2.
+In particular, we avoid explicitly writing the terms $b_{\mu}(\iota)$ in the expression $a_{\iota}=\sum_{\mu=0}^{K}b_{\mu}(\iota)e^{\mu}$ to keep the notation manageable.
+However, it is crucial to keep track of the highest power of $e$, which is $K$, appearing in the coefficients $a_{\iota}$ involved in the multiplications, since the correction applied by the function $F$ depends on this value.
+Precisely, given the fresh ciphertexts, it can be seen that for $a_{0}$ the associated $K$ is $1$, while for $a_{1}$ we have $K=0$.
+Instead, for the terms $\delta_{0}/p_{\ell},\delta_{1}/p_{\ell}$ the case is quite different.
+In fact, these values can be assumed to be randomly distributed over $\mathcal{R}_{t}$ , for every level of the circuit.
+This makes the analysis slightly more complicated, since, in order to derive tight estimates, it will be necessary to distinguish the contribution of $a_{0},a_{1}$} from the one of $\delta_{0},\delta_{1}$.
+Now, assume that $\bm{c},\bm{c}^{\prime}\in\mathcal{R}_{q_{L-2}}^{2}$ are two ciphertexts at level 1, after modulus switching has been carried out.
 Their associated critical quantity is given, respectively, by
-$\begin{cases}\nu_{\mathsf{modswitch}}=\frac{1}{p_{L-1}}a_{0}+\left(\frac{1}{p_{L-1}}a_{1}+a_{1}^{\mathsf{ms}}\right)s\\ \nu^{\prime}_{\mathsf{modswitch}}=\frac{1}{p_{L-1}}a^{\prime}_{0}+\left(\frac{1}{p_{L-1}}a^{\prime}_{1}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s\\ \end{cases}$
+[…]
 Therefore, the critical quantity obtained after their multiplication is of the form
-$\nu_{\mathsf{mul}}^{(1)}=a_{0}^{\mathsf{mul}}+a_{1}^{\mathsf{mul}}s+a_{2}^{\mathsf{mul}}s^{2},$
-namely,
-$\nu_{\mathsf{mul}}^{(1)}$ $=\frac{1}{p_{L-1}^{2}}a_{0}a^{\prime}_{0}+\frac{1}{p^{2}_{L-1}}\left(a^{\prime}_{0}{a_{1}}+a_{0}{a^{\prime}_{1}}\right)s+\frac{1}{p_{L-1}}(a^{\prime}_{0}a_{1}^{\mathsf{ms}}+a_{0}{a^{\prime}_{1}}^{\mathsf{ms}})s\hskip 4.26773pt$
-$\hskip 11.38092pt+\frac{1}{p^{2}_{L-1}}a_{1}a^{\prime}_{1}s^{2}+\frac{1}{p_{L-1}}(a_{1}{a^{\prime}_{1}}^{\mathsf{ms}}+a^{\prime}_{1}{a_{1}}^{\mathsf{ms}})s^{2}+{a^{\prime}_{1}}^{\mathsf{ms}}{a_{1}}^{\mathsf{ms}}s^{2}.$
-So, we are now able to provide an estimate of the variance $V_{1}=\mathsf{Var}(\nu_{\mathsf{mul}}|_{i})$, applying Theorem 4.1, as follows
-$V_{1}$ $\leq\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a^{\prime}_{0}|_{i})F_{e}(1,1)$
-$+\frac{n}{p_{L-1}^{4}}V(a^{\prime}_{0}|_{i})V(a_{1}s|_{i})+\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a^{\prime}_{1}s|_{i})$
-$+\frac{n}{p_{L-1}^{2}}V(a^{\prime}_{0}|_{i})V(a_{1}^{\mathsf{ms}}s|_{i})+\frac{n}{p_{L-1}^{2}}V(a_{0}|_{i})V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})$
-$+\frac{n}{p_{L-1}^{4}}V(a_{1}s|_{i})V(a^{\prime}_{1}s|_{i})F_{s}(1,1)+\frac{n}{p_{L-1}^{2}}V(a_{1}s|_{i})V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})F_{s}(1,1)$
-$+\frac{n}{p_{L-1}^{2}}V(a^{\prime}_{1}s|_{i})V({a_{1}}^{\mathsf{ms}}s|_{i})F_{s}(1,1)+{n}V({a_{1}}^{\mathsf{ms}}s|_{i})V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})F_{s}(1,1).$
-which, recalling that $\operatorname{Var}(a_{\iota}|_{i})=\operatorname{Var}(a^{\prime}_{\iota}|_{i})$ and that $V({a_{1}}^{\mathsf{ms}}s|_{i})=V({a^{\prime}_{1}}^{\mathsf{ms}}s|_{i})=V_{\mathsf{ms}}$, can be rewritten as
-$V_{1}$ $\leq\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})^{2}F_{e}(1,1)+$
-$+\frac{2n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a_{1}s|_{i})+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}V(a_{0}|_{i})$
-$+\frac{n}{p_{L-1}^{4}}V(a_{1}s|_{i})^{2}F_{s}(1,1)+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}V(a_{1}s|_{i})+nF_{s}(1,1)V_{\mathsf{ms}}^{2}.$
-By substituting to $V(a_{0}|_{i})$,$V(a_{1}s|_{i})$ the value $\frac{V_{0}}{2}$, our bound on $V_{1}$ is derived. To conclude, it is possible to observe that the following estimate holds
-$\begin{cases}V(a_{0}^{\mathsf{mul}}|_{i})\approx\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})^{2}F_{e}(1,1)\\ V(a_{1}^{\mathsf{mul}}s|_{i})\approx\frac{2n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a_{1}s|_{i})+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}V(a_{0}|_{i})\\ V(a_{2}^{\mathsf{mul}}s^{2}|_{i})\approx\frac{n}{p_{L-1}^{4}}V(a_{1}s|_{i})^{2}F_{s}(1,1)+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}V(a_{1}s|_{i})+nF_{s}(1,1)V_{\mathsf{ms}}^{2}\end{cases}$ (10)
-These values are needed in order to bound the variance $V_{2}$ at the end of the next level.
-Level 2
-The first computation at level 2 is the modulo switch from $q_{L-2}$ to $q_{L-3}$.
-It should be noted that, for our theoretical analysis, we chose to avoid considering the key switching step.
-Therefore, the ciphertexts which will be subject to modulo switch have critical quantity of the form
-$\nu^{(1)}=a_{0}+a_{1}s+a_{2}s^{2},$
-with $a_{0},a_{1},a_{2}$ having variances as in (10).
-After the modulo switch, the critical quantity associated to the resulting ciphertexts will be as follows
-$\nu_{\mathsf{modswitch}}=\frac{\nu^{(1)}+\nu_{\mathsf{ms}}(p_{L-2})}{p_{L-2}}.$
-Therefore, it is possible to express this quantity as
-$\nu_{\mathsf{modswitch}}=\frac{a_{0}}{p_{L-2}}+\left(\frac{a_{1}}{p_{L-2}}+a_{1}^{\mathsf{ms}}\right)s+\frac{a_{2}}{p_{L-2}}s^{2}.$
-For this level, the maximum power of $e$ contained in each $a_{i}$ has changed.
-In fact, $a_{0}$ contains a term multiplied by $e^{2}$, $a_{1}$ a term multiplied by $e$, while $a_{2}$ does not contain any power of $e$ as for the term $a_{1}^{\mathsf{ms}}$.
-This distinction is crucial in order to correctly apply Theorem 4.1.
-In light of this, given two ciphertexts with critical quantities, respectively,
-$\begin{cases}\nu_{\mathsf{modswitch}}=\frac{a_{0}}{p_{L-2}}+\left(\frac{a_{1}}{p_{L-2}}+{a_{1}}^{\mathsf{ms}}\right)s+\frac{a_{2}}{p_{L-2}}s^{2}\\ \nu^{\prime}_{\mathsf{modswitch}}=\frac{a^{\prime}_{0}}{p_{L-2}}+\left(\frac{a^{\prime}_{1}}{p_{L-2}}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s+\frac{a^{\prime}_{2}}{p_{L-2}}s^{2}\end{cases}$
-the resulting $\nu_{\mathsf{mul}}^{(2)}$ that follows by their multiplication will be of the form
-$\nu_{\mathsf{mul}}^{(2)}$ $=\nu_{\mathsf{modswitch}}\cdot\nu^{\prime}_{\mathsf{modswitch}}=a_{0}^{\mathsf{mul}}+a_{1}^{\mathsf{mul}}s+a_{2}^{\mathsf{mul}}s^{2}+a_{3}^{\mathsf{mul}}s^{3}+a_{4}^{\mathsf{mul}}s^{4},$
-which can be written as
-$\nu_{\mathsf{mul}}^{(2)}$ $=\frac{a_{0}a^{\prime}_{0}}{p^{2}_{L-2}}+\frac{a^{\prime}_{0}}{p_{L-2}}\left(\frac{a_{1}}{p_{L-2}}+{a_{1}}^{\mathsf{ms}}\right)s+\frac{a_{0}}{p_{L-2}}\left(\frac{a^{\prime}_{1}}{p_{L-2}}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s$
-$\hskip 11.38092pt+\left(\frac{a_{1}}{p_{L-2}}+{a_{1}}^{\mathsf{ms}}\right)\left(\frac{a^{\prime}_{1}}{p_{L-2}}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s^{2}+\frac{a_{0}a^{\prime}_{2}}{p^{2}_{L-2}}s^{2}+\frac{a^{\prime}_{0}a_{2}}{p^{2}_{L-2}}s^{2}$
-$\hskip 11.38092pt+\frac{a^{\prime}_{2}}{p_{L-2}}\left(\frac{a_{1}}{p_{L-2}}+{a_{1}}^{\mathsf{ms}}\right)s^{3}+\frac{a_{2}}{p_{L-2}}\left(\frac{a^{\prime}_{1}}{p_{L-2}}+{a^{\prime}_{1}}^{\mathsf{ms}}\right)s^{3}+\frac{a_{2}a^{\prime}_{2}}{p^{2}_{L-2}}s^{4}.$
-A bound over $V_{2}$ can then be deduced, by applying Theorem 4.1 and recalling that $\mathsf{Var}(a_{\iota}|_{i})=\mathsf{Var}(a^{\prime}_{\iota}|_{i})$, resulting in
-$V_{2}\leq$ $\,\frac{n}{p_{L-2}^{4}}\mathsf{Var}(a_{0}|_{i})^{2}F_{e}(2,2)+\frac{2n}{p_{L-2}^{4}}\mathsf{Var}(a_{0}|_{i})\mathsf{Var}(a_{1}s|_{i})F_{e}(1,2)$
-$+\frac{n}{p_{L-2}^{4}}\mathsf{Var}(a_{1}s|_{i})^{2}F_{s}(1,1)F_{e}(1,1)+\frac{2n}{p_{L-2}^{4}}\mathsf{Var}(a_{0}|_{i})\mathsf{Var}(a_{2}s^{2}|_{i})$
-$+\frac{2n}{p_{L-2}^{4}}\mathsf{Var}(a_{1}s|_{i})\mathsf{Var}(a_{2}s^{2}|_{i})F_{s}(1,2)+\frac{n}{p_{L-2}^{4}}\mathsf{Var}(a_{2}s^{2}|_{i})^{2}F_{s}(2,2)$
-$+\frac{2n}{p_{L-2}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{0}|_{i})+\frac{2n}{p_{L-2}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{1}s|_{i})F_{s}(1,1)$
-$+\frac{2n}{p_{L-2}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{2}s^{2}|_{i})F_{s}(1,2)+{n}V_{\mathsf{ms}}^{2}F_{s}(1,1),$
-where the values for $V(a_{\iota}s^{\iota}|_{i})$ follows by (10).
-Moreover, similar to level $1$, it is possible to deduce the contribution, in terms of variance, of each $a_{\iota}^{\mathsf{mul}}$ of $\nu_{\mathsf{mul}}^{(2)}$, which will then be needed for the successive levels.
-Level ${l}$
-For a generic level $l$, with $l\in\{1,\dots,L-1\}$, the same reasoning applied in the first levels can be reiterated. Assume that two ciphertexts with critical quantity of the form
-$\nu^{(l-1)}=\sum_{\iota=0}^{2^{l-1}}a_{\iota}s^{\iota},$
-are given as input. Let $V_{l}$ be the variance of the coefficients of $\nu^{(l)}$, which is the critical quantity obtained after modulo switch and multiplication.
-Following the notation adopted in the precedent levels, this quantity can be indicated as
-$\nu^{(l)}=\sum_{\iota=0}^{2^{l}}a_{\iota}^{\mathsf{mul}}s^{\iota}.$
-Then, $V_{l}$ is bounded by the sum of
-$V_{l}$ $\leq\frac{n}{p_{L-l}^{4}}\sum_{\iota=0}^{2^{l}}\sum_{k+\mu=\iota}\mathsf{Var}(a_{k}s^{k}|_{i})\mathsf{Var}(a_{\mu}s^{\mu}|_{i}){F_{s}(\iota,\mu)}{F_{e}(2^{l-1}-\iota,2^{l-1}-\mu)}$
-$\quad+\frac{2n}{p_{L-l}^{2}}\sum_{\iota=0}^{2^{l-1}}V_{\mathsf{ms}}\mathsf{Var}(a_{\iota}s^{\iota}|_{i}){F_{s}(1,\iota)}+nV_{\mathsf{ms}}^{2}{F_{s}(1,1)}$ (11)
-Moreover, each term of $\nu^{(l)}$ has variance
-$\begin{cases}\mathsf{Var}(a_{0}^{\mathsf{mul}}|_{i})=\frac{n}{p_{L-l}^{4}}\sum_{k+\mu=0}\mathsf{Var}(a_{k}s^{k}|_{i})\mathsf{Var}(a_{\mu}s^{\mu}|_{i}){F_{s}(k,\mu)}{F_{e}(2^{l-1}-k,2^{l-1}-\mu)}\\ \mathsf{Var}(a_{1}^{\mathsf{mul}}s|_{i})=\frac{n}{p_{L-l}^{4}}\sum_{k+\mu=1}\mathsf{Var}(a_{k}s^{k}|_{i})\mathsf{Var}(a_{\mu}s^{\mu}|_{i}){F_{s}(k,\mu)}{F_{e}(2^{l-1}-k,2^{l-1}-\mu)}\\ \hskip 65.44142pt+\frac{2n}{p_{L-l}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{0}|_{i}){F_{s}(1,0)}\\ \mathsf{Var}(a_{2}^{\mathsf{mul}}s^{2}|_{i})=\frac{n}{p_{L-l}^{4}}\sum_{k+\mu=2}\mathsf{Var}(a_{k}s^{k}|_{i})\mathsf{Var}(a_{\mu}s^{\mu}|_{i}){F_{s}(k,\mu)}{F_{e}(2^{l-1}-k,2^{l-1}-\mu)}\\ \hskip 65.44142pt+\frac{2n}{p_{L-l}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{1}s|_{i}){F_{s}(1,1)}+nF_{s}(1,1)V_{\mathsf{ms}}^{2}\\ \mathsf{Var}(a_{\iota}^{\mathsf{mul}}s^{\iota}|_{i})=\frac{n}{p_{L-l}^{4}}\sum_{k+\mu=\iota}\mathsf{Var}(a_{k}s^{k}|_{i})\mathsf{Var}(a_{\mu}s^{\mu}|_{i}){F_{s}(k,\mu)}{F_{e}(2^{l-1}-k,2^{l-1}-\mu)}\\ \hskip 65.44142pt+\frac{2n}{p_{L-l}^{2}}V_{\mathsf{ms}}\mathsf{Var}(a_{\iota-1}s^{\iota-1}|_{i}){F_{s}(1,\iota-1)}\mathbb{I}_{\{0,\dots,2^{l-1}+1\}}(\iota)\end{cases}$
-where $\mathbb{I}_{\{0,\dots,2^{l-1}+1\}}(\iota)$ is the indicator function, i.e.,
-$\mathbb{I}_{\{0,\dots,2^{l-1}+1\}}(\iota)=\begin{cases}1,&\text{if }\iota\in\{0,\dots,2^{l-1}+1\}\\ 0,&\text{if }\iota\notin\{0,\dots,2^{l-1}+1\}\end{cases}$
+$\nu_{1}$ $=\frac{1}{p_{L-1}^{2}}a_{0}a^{\prime}_{0}+\frac{1}{p^{2}_{L-1}}\left(a^{\prime}_{0}{a_{1}}+a_{0}{a^{\prime}_{1}}\right)s+\frac{1}{p_{L-1}}(a^{\prime}_{0}\frac{\delta_{1}}{p_{L-1}}+a_{0}\frac{\delta^{\prime}_{1}}{p_{L-1}})s\hskip 4.26773pt$
+$\hskip 11.38092pt+\frac{1}{p^{2}_{L-1}}a_{1}a^{\prime}_{1}s^{2}+\frac{1}{p_{L-1}}(a_{1}\frac{\delta^{\prime}_{1}}{p_{L-1}}+a^{\prime}_{1}\frac{\delta_{1}}{p_{L-1}})s^{2}+\frac{\delta^{\prime}_{1}}{p_{L-1}}\frac{\delta_{1}}{p_{L-1}}s^{2}.$
+So, we are now able to provide an estimate of the variance $V_{1}$, applying Theorem 4.2, as follows
+$V_{1}$ $\leq\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a^{\prime}_{0}|_{i})F_{e}(1,1)+\frac{n}{p_{L-1}^{4}}V(a^{\prime}_{0}|_{i})V(a_{1}s|_{i})+\frac{n}{p_{L-1}^{4}}V(a_{0}|_{i})V(a^{\prime}_{1}s|_{i})$
+$+\frac{n}{p_{L-1}^{2}}V(a^{\prime}_{0}|_{i})V(\frac{\delta_{1}}{p_{L-1}}s|_{i})+\frac{n}{p_{L-1}^{2}}V(a_{0}|_{i})V(\frac{\delta^{\prime}_{1}}{p_{L-1}}s|_{i})$
+$+\frac{n}{p_{L-1}^{4}}V(a_{1}s|_{i})V(a^{\prime}_{1}s|_{i})F_{s}(1,1)+\frac{n}{p_{L-1}^{2}}V(a_{1}s|_{i})V(\frac{\delta^{\prime}_{1}}{p_{L-1}}s|_{i})F_{s}(1,1)$
+$+\frac{n}{p_{L-1}^{2}}V(a^{\prime}_{1}s|_{i})V(\frac{\delta_{1}}{p_{L-1}}s|_{i})F_{s}(1,1)+{n}V(\frac{\delta_{1}}{p_{L-1}}s|_{i})V(\frac{\delta^{\prime}_{1}}{p_{L-1}}s|_{i})F_{s}(1,1).$
+which, recalling that $\operatorname{Var}(a_{\iota}|_{i})=\operatorname{Var}(a^{\prime}_{\iota}|_{i})={V_{0}}/{2}$ and that $V(\frac{\delta_{1}}{p_{L-1}}s|_{i})=V(\frac{\delta^{\prime}_{1}}{p_{L-1}}s|_{i})=V_{\mathsf{ms}}$ can be rewritten as
+$V_{1}$ $\leq\frac{n}{p_{L-1}^{4}}\frac{V_{0}^{2}}{4}F_{e}(1,1)+\frac{2n}{p_{L-1}^{4}}\frac{V_{0}^{2}}{4}+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}\frac{V_{0}}{2}$
+$+\frac{n}{p_{L-1}^{4}}\frac{V_{0}^{2}}{4}F_{s}(1,1)+\frac{2n}{p_{L-1}^{2}}V_{\mathsf{ms}}\frac{V_{0}}{2}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}F_{s}(1,1)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+nF_{s}(1,1)V_{\mathsf{ms}}^{2}.$
+By observing that the condition in Equation 12 holds, our bound on $V_{1}$ is derived as follows
+$V_{1}\approx\left(\frac{3}{2\cdot 100^{2}}+\frac{3}{100}+2\right)nV_{\mathsf{ms}}^{2}\approx 2nV_{\mathsf{ms}}^{2},$
+and we can simplify $\nu_{1}=\delta_{1}\delta^{\prime}_{1}s^{2}/p_{L-1}$.
+We now present the generic level $\ell$.
+We begin by illustrating the case $\ell=2$.
+As we shall see, the input to level 2 corresponds to that of a generic level $\ell\geq 2$, and it produces the same type of output, thus making the analysis carried out for the second level applicable to all subsequent levels.
+Level 2{2}.
+At the beginning of level $2$ modulo switch is performed, yielding a critical quantity of the form $\nu_{2}^{\mathsf{ms}}={\nu_{1}}/{p_{L-2}}+{\delta_{1}}s/{p_{L-2}}$  , with coefficient variance,
+$V_{2}^{\mathsf{ms}}=\frac{V_{1}}{p_{L-2}^{2}}+V_{\mathsf{ms}}\leq\left(\frac{1}{100}+1\right)V_{\mathsf{ms}}\,.$
+Now, the product of two critical quantities $\nu_{1}^{\mathrm{ms}}$ can be expanded into four terms: the first is the product of two $\nu_{1}/p_{L-2}$ terms, the second and third are the products of a $\nu_{1}/p_{L-2}$ term with a $\nu_{1}^{\mathrm{ms}}$ term and the fourth is the product of two $\nu_{1}^{\mathrm{ms}}$ terms.
+Applying our theorem, the variance after the multiplication is then given by
+$V_{2}=n\left(\frac{V_{1}^{2}}{p_{L-2}^{4}}F_{s}(2,2)+2\frac{V_{1}}{p_{L-2}^{2}}V_{\mathsf{ms}}F_{s}(2,1)+2V_{\mathsf{ms}}^{2}\right),$
+which can be bounded, for the conditions over $p_{L-2}$, as
+$V_{2}\leq\left(\frac{6}{100^{2}}+\frac{6}{100}+2\right)nV_{\mathsf{ms}}^{2}\approx 2nV_{\mathsf{ms}}^{2}\,.$
+Again, the critical quantity at the end of level 2 can be simplified as $\nu_{2}=\delta_{1}\delta^{\prime}_{1}s^{2}/p_{L-2}$.
+From this observation, we can deduce the variance estimates for each level $\ell$.
+Level $\ell$.
+For any level $\ell\geq 2$, the variance after modulus switching $V_{\ell}^{\mathsf{ms}}$ and the variance after multiplication $V_{\ell}$ can be approximated as
+$V_{\ell}^{\mathsf{ms}}\approx\,\frac{101}{100}V_{\mathsf{ms}}$
+$V_{\ell}\,\,\,\,\approx\,(2+\epsilon)nV_{\mathsf{ms}}^{2},\quad\epsilon=\frac{6}{100^{2}}+\frac{6}{100}$ (15)
+Table 1 presents a comparison of the estimated variances of the error coefficients obtained using our approach, denoted as our and the corresponding experimental values, denoted as exp.
+The estimates were obtained using the OpenFHE library, setting the parameters according to the required multiplicative depth, with $t=65537$.
+For the experimental values, 50000 samples were computed for $n=2^{13}$, 8000 for $n=2^{14}$ and 5000 for $n=2^{15}$.
+To enhance readability, variances are presented in terms of their base-2 logarithms.
+   Encryption     Modulo Switch    1 Multiplication    6 Multiplications
+   $n$    our    exp    our    exp    our    exp    our    exp
+   $2^{13}$    48.76    48.76    40.84    40.83    95.68    95.65    95.71    95.25
+   $2^{14}$    49.76    49.76    41.84    41.79    98.68    97.62    98.71    97.63
+   $2^{15}$    50.76    50.72    42.84    42.82    101.68    101.62    101.71    101.59
+Table 1: Encryption, modulo switch, and multiplication of fresh ciphertexts. The last two columns report the case of six consecutive multiplications as in the reference circuit.
+Remark 2
+At first glance, it might seem that, in order to correctly estimate the variance of the multiplication, it is sufficient to apply a factor of 2, arising from the correction of the product between two $s$-terms, thus making it unclear why the dependence on the public key should be considered.
+It is crucial to emphasize that these estimates are strictly related to a circuit model that we are specifically constructing to guarantee the Gaussianity of the error coefficients according to Equation 12.
+However, if one wanted to accurately study the variance in different circuits, where the term $V_{\mathsf{ms}}$ is not the only significant contribution, the need to account for these dependencies would become even more evident.
+This necessity reaches its maximum importance when considering circuits in which modulus switching is not applied.
+Since the use of modulus switching is mandatory for the practical implementation of BGV, in this work we omitted the results obtained from the analysis of such circuits.
+However, whether modulus switching is not used, or it is applied with primes that do not satisfy Equation 12, it remains crucial to take into account all dependencies arising from both $e$ and $s$ as well as from their powers.
+As an example of the applicability of our approach, if error coefficient distributions that are non-Gaussian, but still suitable for use, were considered, the variance analysis we propose would remain valid and could still provide accurate estimates that never underestimate the experimental values (see Theorem 4.2).
+7 Comparison with Previous Works
+When estimating the error in schemes such as BGV, existing approaches can broadly be divided into two main categories: worst-case and average-case.
+The former includes analyses based on the Euclidean norm, as in [6], the infinity norm, as in [22, 30], and the canonical norm, as in [15, 17, 25, 29, 34], with the latter providing the tightest bounds among the worst-case analyses.
+In contrast, average-case approaches model the noise coefficients as random variables and focus on their mean and variance to derive probabilistic bounds on the minimum ciphertext modulus required to guarantee correctness.
+Although the latter approaches appear more promising in reducing the ciphertext modulus size, research for a general method to estimate the variance remains incomplete.
+Most works, such as [35], treat the error coefficients as independent, which, as pointed out in [4, 16, 35], leads to imprecise bounds and often underestimates the modulus size required to prevent decryption failures.
+Consequently, the scheme may be exposed to potential vulnerabilities, including key-recovery attacks [10].
+Other types of methods, by contrast, avoid such underestimations or failures but are limited to specific settings, as in [18], where the authors analyze the error for HElib [28].
+The objective of our work is to develop an average case approach that takes into account the intrinsic dependencies introduced by the secret and public key, allowing to obtain variance estimates very close to experimental values, without ever underestimating them.
+At the same time, our aim is to propose a method independent of the specific library or circuit considered, providing the theoretical basis necessary to build ad hoc estimates depending on the circuit and library of interest.
+Although, to highlight the effectiveness of our approach, we have focused on circuits that exclusively include multiplications preceded by the modulus switch, our analyses and estimates remain valid for any type of homomorphic circuit.
+Our method is in fact intrinsically modular: each homomorphic operation is analyzed independently, producing variance bounds that can be composed to precisely estimate the overall noise growth of a given circuit, assuming only that the multiplications are preceded by a modulus switch – a completely reasonable hypothesis in the context of BGV typical usage.
+In this section, we demonstrate the efficacy of our average-case approach by comparing it to the state-of-the-art works.
+Specifically, we illustrate how our average-case analysis provides tighter and more practical bounds than traditional worst-case methods [34], and we further compare it with other average-case approaches [35, 18].
+We then show how our choice of ciphertext moduli achieves a substantial size reduction compared to those adopted in widely used libraries such as OpenFHE [1] and HElib [28].
+Canonical norm.
+For the comparison with the worst-case approach, we specifically refer to the estimates proposed in [34].
+   Homomorphic operation    Error bounds with canonical norm
+   Enc    $\nu_{\mathsf{clean}}||^{\text{can}}\leq Dt\sqrt{n\left(1/12+2nV_{e}V_{s}+V_{e}\right)}$
+[…]
+[…]
+Table 2: Canonical norm depending on the homomorphic operations.
+[…]
+In contrast, average-case approaches, such as the one proposed in this work, allow significantly tighter bounds by imposing the condition $\|\nu\|_{\infty}\leq D\sqrt{2V}$, where $V$ denotes the variance of each coefficient of $\nu$.
+According to Equation 2, this bound holds with probability at least $1-n(1-\operatorname{erf}(D))$, which, for $D=6$, exceeds $1-2^{-40}$.
+However, in practical scenarios, choosing $D=8$ is preferable, as it limits the failure probability to $2^{-77}$, when $n\leq 2^{15}$.
+Note that this bound better captures the actual distribution of the error and results in much tighter predictions after multiplications.
+In fact, by Section 6, after a multiplication that follows modulo switch, we have $\nu_{\text{mul}}||_{\infty}\leq 2D\sqrt{n}V_{\mathsf{ms}}$.
+Current Average-Case Approaches.
+Within average-case analysis frameworks, the main novelty of our approach lies in tracking the dependencies between the error coefficients, particularly those arising from the multiplication of two ciphertexts.
+Specifically, given the critical quantities $\nu$ and $\nu^{\prime}$ of two ciphertexts with coefficient variances $V$ and $V^{\prime}$, respectively, the current average-case methodology [35] yields the expression $\mathsf{Var}(\nu\nu^{\prime}|_{i})\approx nVV^{\prime}$.
+Therefore, assuming that two ciphertexts have just undergone modulus switching, the variance resulting from their product is, in accordance with the previously established considerations
+$\mathsf{Var}(\nu\nu^{\prime}|_{i})\approx nV_{\mathsf{ms}}^{2}\,.$
+In contrast, our method estimates this value according to Section 6 , resulting in
+$\mathsf{Var}(\nu\nu^{\prime}|_{i})\approx 2nV_{\mathsf{ms}}^{2}\,.$
+The factor 2, which arises from accounting for the dependencies among the critical quantities, allows for extremely accurate values, as shown in Table 1.
+From this comparison, it should be evident that accounting for the dependencies in the coefficients of the error polynomial is fundamental for obtaining accurate and correct estimates of the experimental variance.
+In particular, we believe that this consideration is precisely what overcomes the underestimation inherent in the approach presented in [35].
+It is crucial to point out that the factor of 2 is specific to the type of circuit we constructed, in which the variance is independent of the circuit level.
+However, when analyzing errors in circuits where this condition is not met, it is still possible to derive general upper bounds using Theorem 4.2, once again obtaining estimates that never underestimate the error.
+Nevertheless, for the reasons outlined above, including guaranteeing the Gaussianity of the error, we recommend choosing primes that allow the variance of the error after the modulus switch to be well approximated by $V_{\mathsf{ms}}$, as explained previously.
+Our results appear to be very close to those reported in [18]. In particular, it can be observed that a factor of 2 also appears in [18, Lemma 9].
+Current Libraries.
+Finally we compare the ciphertext modulus $q$ as estimated by our method against those adopted by two of the most widely used libraries: OpenFHE [1] and HElib [28].
+We recall that the ciphertext modulus must be selected to ensure that $\|\nu\|<\frac{q}{2}$.
+In our selection, we additionally require (12) to guarantee the Gaussian behavior of the error.
+OpenFHE.
+In OpenFHE [1], the generation of the moduli depends on the multiplicative depth selected as input, in a manner similar to the choice of moduli described in this paper.
+Moreover, the way multiplications are executed in a fixed-depth circuit of the library is consistent with the circuits analyzed in our work, where each multiplication between two ciphertexts is preceded by a modulus switching operation.
+For a circuit of multiplicative depth $M$, $M+2$ moduli are generated, to allow for an additional modulus switching after the final multiplication.
+Therefore, we generate the moduli in a way consistent with the library’s design, to allow a fair comparison between the modulus sizes used in OpenFHE and those proposed in this paper.
+In Figure 5, we compare the ciphertext modulus sizes $q$ for circuits with multiplicative depths 3 (Figure 5(a)) and 6 (Figure 5(b)).
+The values are obtained using our method (denoted as our) and those generated by OpenFHE (denoted as OpenFHE).
+Note that, for the generation of our parameters, we fixed $D=8$.
+   $n$    $2^{12}$    $2^{13}$    $2^{14}$    $2^{15}$
+   OpenFHE    147.3    151.8    156.3    161.6
+   our    121.0    124.5    128.0    131.5
+(a) Circuit of depth 33.
+   $n$    $2^{12}$    $2^{13}$    $2^{14}$    $2^{15}$
+   OpenFHE    249.3    256.8    264.3    272.6
+   our    210.3    216.8    223.3    229.8
+(b) Circuit of depth 66.
+Figure 5: Comparison of $\log_{2}(q)$ in the circuit shown in Figure 4 (setting $D=8$).
+From this comparison, it should be evident that our approach can lead to an effective reduction in the modulus size, resulting in corresponding improvements in the overall efficiency of the scheme.
+HElib Comparison.
+To provide a more comprehensive overview of the effectiveness of our work and to highlight its applicability across various contexts and libraries, we also present a comparison between our choice of moduli and those used in HElib [27], currently one of the most competitive FHE libraries, together with OpenFHE.
+We first point out that a strategy for optimizing HElib specific parameters has been analyzed in [18].
+However, a direct comparison with that approach lies beyond the scope of this work.
+Our focus is instead on emphasizing that the strength of our method lies in its library-independent nature, offering a general framework for accurately estimating noise in different types of circuits and determining the conditions that ciphertext moduli must meet to ensure both efficiency and security (as well as the Gaussianity of the noise distributions, see Section 5).
+For this reason, we intend to provide in this section just an idea of how our method can potentially improve current HElib parameters as well.
+Our method proposes to select parameters based on the multiplicative depth of the circuit, as in OpenFHE and in previous versions of HElib.
+However, in the new version of the library, the selection of parameters is based on the bit length of the largest modulus in the chain.
+Therefore, to compare the two approaches, we focus on comparing the ratio between the moduli of successive levels, as also done in [18].
+Due to the way ciphertext moduli are constructed in HElib, the ratio between two moduli of adjacent levels, i.e.
+a prime $p_{\ell}$, is necessarily larger than 36 bits.
+However, as observed in [18], it is typically much larger, often reaching sizes of around 54 bits, for a ring dimension $n\in\{2^{12},2^{13},2^{14},2^{15}\}$.
+Instead, with our method, we simply require that Equation 12 is satisfied, i.e., for 0<l<L10<l<L-1
+$p_{\ell}$ $>\sqrt{\frac{(2+\epsilon)nV_{\mathsf{ms}}}{\alpha}}\,.$
+Table 3 reports the bit-size of the ratio between adjacent moduli resulting from our approach, for ring dimensions in $\{2^{12},2^{13},2^{14},2^{15}\}$, assuming a secret key with coefficients sampled from a ternary uniform distribution, with hamming weight $h=\nicefrac{{n}}{{2}}$, plaintext modulus $t=65537$ and $\alpha=\nicefrac{{1}}{{100}}$.
+As shown in the table, our approach achieves a reduction of up to 6 bits in the ratio between consecutive ciphertext moduli.
+   $n$    $2^{12}$    $2^{13}$    $2^{14}$    $2^{15}$
+   $\log_{2}(p_{\ell})$    29.55    30.55    31.55    32.55
+Table 3: Ratio between adjacent ciphertext moduli for different ring dimensions $n$, according to our approach, represented using their bit size.
 8 Conclusions
-In this work, we focused on improving the current average-case approach for tracking the error and its growth under homomorphic operations.
-In particular, we demonstrated that the primary reason for the underestimation of noise in existing average-case methods lies in the fact that the coefficients of the error polynomial are treated as independent.
-However, to obtain accurate bounds that never underestimate the actual values observed in practice, it is crucial to account for the dependencies among these coefficients.
-Accurate noise estimation plays a key role in the proper selection of scheme parameters, especially the ciphertext modulus $q$, to ensure correctness, security, and improved efficiency.
-To evaluate the effectiveness of our analysis, we applied our results to specific homomorphic circuits, comparing our estimations with both experimental data and those obtained from the current average-case approach.
-The circuits were implemented using the MAGMA language.
-Our findings show that the proposed method effectively overcomes the limitations of existing average-case techniques, particularly their tendency to underestimate noise variance.
-Furthermore, in the context of initial parameter selection, the high accuracy of our estimations, consistently close to experimental results, represents a significant improvement over the worst-case methods that currently dominate state-of-the-art parameter generation.
-This leads to tighter bounds and thus enables more efficient configurations of the scheme, which is especially critical for promoting the practical adoption of homomorphic encryption, and in particular of the BGV scheme, in real-world applications.
-Future work will include extending our approach to other cryptographic schemes, such as CKKS, as well as studying scenarios in which ciphertexts are generated in a dependent manner.
-We are currently developing a comprehensive framework for modulus selection that is fully aligned with the new approach introduced in this paper.
+In this work, we introduce a new approach to average cases that can accurately estimate the error arising from homomorphic operations, especially multiplication.
+The main novelty of our method lies in the definition of error bounds that account for the dependencies between the critical quantities of multiplied ciphertexts, arising from both the public and private keys.
+We believe that this is precisely what enables our approach to overcome the typical underestimations observed in current average-case analyses, thereby providing accurate bounds on the error variance without ever underestimating it.
+Furthermore, this work aims to provide general guidelines for studying noise growth in circuits independently of the specific implementation of the homomorphic encryption library employed.
+Based on these estimates, we present a method to select ciphertext moduli appropriately in a generic circuit, demonstrating that the new solution facilitates a significant reduction in their size.
+This leads to improved efficiency compared to current techniques and compared to the moduli employed in major homomorphic encryption libraries, with no loss of security and proper scheme functioning.
+Finally, we provide theoretical and empirical evidence supporting the validity of average-case approaches in the study of homomorphic schemes such as BGV.
+Precisely, we show that a proper choice of moduli together with application of modulus switching on a systematic basis in BGV allow error distributions to be well approximated by Gaussian distributions.
+This confirms the validity of such approaches and highlights their practical relevance in improving efficiency, thereby contributing to the potential widespread adoption of these schemes in real-world applications.
 Acknowledgment
-The third author was partially supported by the Italian Ministry of University and Research in the framework of the Call for Proposals for scrolling of final rankings of the PRIN 2022 call - Protocol no.
-2022RFAZCJ and by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.
+The third and fourth authors were partially supported by the Italian Ministry of University and Research in the framework of the Call for Proposals for scrolling of final rankings of the PRIN 2022 call - Protocol no.
+2022RFAZCJ.
+The third author acknowledges project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.
 Appendix 0.A Proof of Lemma 1
@@
 From which it follows that
-$b^{\mathsf{mul}}_{\mu}({\iota})=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}b_{\mu_{1}}({\iota_{1}})b^{\prime}_{\mu_{2}}({\iota_{2}}).$
+$b^{\mathsf{mul}}_{\mu}({\iota})|_{i}=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\sum_{j}b_{\mu_{1}}({\iota_{1}})|_{j}\,b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i-j}.$
 From the independence of $b_{\mu_{1}}(\iota_{1})$ and $b^{\prime}_{\mu_{2}}(\iota_{2})$ $\forall\mu_{1},\mu_{2},\iota_{1},\iota_{2}$, and for the linearity of the expected value, one can deduce that
-$\mathbb{E}[b^{\mathsf{mul}}_{\mu}({\iota})]$ $=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\mathbb{E}[b_{\mu_{1}}({\iota_{1}})b^{\prime}_{\mu_{2}}({\iota_{2}})]$
-$=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\mathbb{E}[b_{\mu_{1}}({\iota_{1}})]\mathbb{E}[b^{\prime}_{\mu_{2}}({\iota_{2}})]=0,$
+$\mathbb{E}[b^{\mathsf{mul}}_{\mu}({\iota})|_{i}]$ $=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\sum_{j}\mathbb{E}[b_{\mu_{1}}({\iota_{1}})|_{j}\,b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i-j}]$
+$=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\sum_{j}\mathbb{E}[b_{\mu_{1}}({\iota_{1}})|_{j}]\mathbb{E}[b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i-j}]=0,$
 which easily proves property $b)$.
@@
 The critical quantity associated to $\bm{c^{\prime}}$ can be expressed as
-$\nu^{\prime}=\frac{q_{l^{\prime}}}{q_{l}}(\nu+\nu_{\mathsf{ms}})\quad\text{ where}\quad\nu_{\mathsf{ms}}=\delta_{0}+\delta_{1}s$
+$\nu^{\prime}=\frac{q_{l^{\prime}}}{q_{l}}(\nu+\nu_{\mathsf{ms}})\quad\text{ where}\quad\nu_{\mathsf{ms}}=\delta_{0}+\delta_{1}s\,.$
 It is possible to observe that the ciphertext components ${c}_{0},c_{1}$ can be thought as randomly distributed over $R_{q_{l}}$, ${c_{0}},c_{1}\leftarrow\mathcal{U}_{q_{l}}$, and therefore the $\delta_{i}$ can be treated as independent polynomials with coefficients chosen randomly over $I={(-\frac{tq_{l}}{2q_{l^{\prime}}},\frac{tq_{l}}{2q_{l^{\prime}}})}$, i.e.
@@
 $a_{\iota}s^{\iota}|_{i}=\sum_{\mu}(b_{\mu}(\iota)e^{\mu}s^{\iota})|_{i}=\sum_{\mu}\sum_{j=0}^{n-1}\xi(i,j)b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j}.$
-Recall that, given two random variables X and Y, the following properties hold:
-[…]
+Thus, given two random variables $X$ and $Y$, the following properties hold:
+$\text{a. \hskip 5.69046pt}\mathsf{Var}(XY)=(\mathsf{Var}(X)+\mathbb{E}[X]^{2})(\mathsf{Var}(Y)+\mathbb{E}[Y]^{2})+\mathsf{Cov}(X^{2},Y^{2})$
 $\hskip 68.28644pt-(\mathsf{Cov}(X,Y)+\mathbb{E}[X]\mathbb{E}[Y])^{2}$
-[…]
-where the second property can be generalized for $k$ random variables $\{X_{i}\}_{i=0}^{k}$ as follows
+$\text{b. \hskip 5.69046pt}\mathsf{Var}(X+Y)=\mathsf{Var}(X)+\mathsf{Var}(Y)+2\mathsf{Cov}(X,Y),$
+where the second property can be generalized for $k$ random variables $\{X_{i}\}_{i=0}^{k}$ as
 $\mathsf{Var}\left(\sum_{i=0}^{k}X_{i}\right)=\sum_{i=0}^{k}\mathsf{Var}(X_{i})+\sum_{i_{1}\neq i_{2}}\mathsf{Cov}(X_{i_{1}},X_{i_{2}}).$
@@
 $+\sum_{\mu_{1}\neq\mu_{2}\text{ or }j_{1}\neq j_{2}}\xi(i,j_{1})\xi(i,j_{2})\mathsf{Cov}(b_{\mu_{1}}(\iota)e^{\mu_{1}}|_{j_{1}}s^{\iota}|_{i-j_{1}},b_{\mu_{2}}(\iota)e^{\mu_{2}}|_{j_{2}}s^{\iota}|_{i-j_{2}}),$
-where the covariances vanishes based on property 1.
-Thus, the following equality holds
+where the covariances vanishes based on property 1. Thus, the following equality holds
 $\mathsf{Var}(a_{\iota}s^{\iota}|_{i})=\sum_{\mu}\sum_{j=0}^{n-1}\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j}).$ (16)
@@
   • $\mathsf{Cov}(b_{\mu}(\iota)e^{\mu}|_{j},s^{\iota}|_{i-j})=\mathsf{Cov}(b_{\mu}(\iota)e^{\mu}|_{j}^{2},s^{\iota}|_{i-j}^{2})=0$ as $b_{\mu}(\iota)e^{\mu}|_{j},s^{\iota}|_{i-j}$ are independent;
-This results in
-$\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j})=\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j})(\mathsf{Var}(s^{\iota}|_{i-j})+\mathbb{E}[s^{\iota}|_{i-j}]^{2}).$
-In addition, for a random variable $X$, it holds that
-$\mathsf{Var}(X)=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}.$
+This results in $\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j})=\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j})(\mathsf{Var}(s^{\iota}|_{i-j})+\mathbb{E}[s^{\iota}|_{i-j}]^{2})$.
+In addition, for a random variable $X$, it holds that $\mathsf{Var}(X)=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}$.
 Therefore,
@@
 $\mathsf{Var}(a_{\iota}s^{\iota}|_{i})=\sum_{\mu}\mathsf{Var}(b_{\mu}(\iota)|_{i})\sum_{k=0}^{n-1}\mathbb{E}[e^{\mu}|_{k}^{2}]\sum_{j=0}^{n-1}\mathbb{E}[s^{\iota}|_{j}^{2}].$
+The expression for the variance of the coefficients of the critical quantity simply follows from the observation that the covariance between $a_{\iota_{1}}s^{\iota_{1}}|_{i}$ and $a_{\iota_{2}}s^{\iota_{2}}|_{i}$ vanishes whenever $\iota_{1}\neq\iota_{2}$.
+Indeed, by bilinearity of the covariance operator, $\mathsf{Cov}(a_{\iota_{1}}s^{\iota_{1}}|_{i},a_{\iota_{2}}s^{\iota_{2}}|_{i})$ can be expressed as a sum of terms of the form
+$\mathsf{Cov}(b_{\mu_{1}}({\iota_{1}})|_{j}(e^{\mu_{1}}s^{\iota_{1}})|_{i-j},b_{\mu_{2}}({\iota_{2}})|_{j}(e^{\mu_{2}}s^{\iota_{2}})|_{i-j})\,$,
+each of which vanishes by Property 1. Therefore, we obtain
+$\mathsf{Var}(\nu|_{i})=\mathsf{Var}(\sum_{\iota\geq 0}a_{\iota}s^{\iota}|_{i})=\sum_{\iota\geq 0}\mathsf{Var}(a_{\iota}s^{\iota}|_{i})+\sum_{\iota_{1}\neq\iota_{2}}\mathsf{Cov}(a_{\iota_{1}}s^{\iota_{1}}|_{i},a_{\iota_{2}}s^{\iota_{2}}|_{i})\,$,
+thus concluding the proof.
+Appendix 0.C On the Negligibility of Key Switching
+In this work, we chose to omit and not explicitly address the contribution of key switching in our analysis of noise variance, despite its necessity for ensuring the practicality of evaluated circuits.
+This decision is justified by the fact that, in general, the parameters associated with key switching can be selected so that its impact on the overall noise remains negligible in comparison to that of other operations — especially the multiplication operation, which immediately precedes the key switching step.
+While our approach can easily accommodate this contribution, we opted to exclude it from our presentation for the sake of clarity and due to the aforementioned reason.
+Various key switching methods exist.
+To support the reasonableness of our choice, we provide a justification of its negligible impact for one of the main variants: GHS (Gentry Halevi Smart).
+Let $q$ be the modulus of the ciphertext to be relinearized, and let $Q$ be a modulus specifically chosen for this purpose, such that $q<Q$ and $q\mid Q$.
+The core idea of the GHS variant is to perform relinearization in the larger ring $\mathcal{R}_{Q}$, and then apply modulus switching to return to the smaller ring $\mathcal{R}_{q}$, reducing the error.
+Let $c=(c_{0},c_{1},c_{2})$ be the result of the multiplication of two ciphertexts, with $c_{0},c_{1},c_{2}\in\mathcal{R}_{q}$.
+The key-switching key for the GHS variant is defined as
+$({ek_{0}},{ek_{1}})=\left(\left(-a\cdot s+te+\frac{Q}{q}s^{2}\right),a\right)\mod Q\,.$
+The ciphertext resulting from relinearization in $R_{Q}$ is given by
+[…]
+Then, this new ciphertext is scaled back modulo $q$ using the modulus switching technique previously introduced, thereby obtaining
+[…]
+It is easy to verify that the critical quantity of the new relinearized ciphertext is given by
+$\nu_{\mathsf{ks}}=[\hat{c_{0}}+\hat{c_{1}}\cdot s]_{q}=\nu_{\mathsf{mul}}+\left[\frac{q}{Q}\left(tc_{2}\cdot e+\delta_{0}+\delta_{1}\cdot s\right)\right]_{q}\,.$
+A typical choice for the size of $Q$ is $Q\approx q^{2}$ [30]. In this way, the variance of the second component will be
+$V_{\frac{q}{Q}(tc_{2}\cdot e+\delta_{0}+\delta_{1}\cdot s)}=\frac{1}{q^{2}}\left({nt^{2}V_{e}V_{c_{2}}}+{V_{\delta_{0}}+nV_{s}V_{\delta_{1}}}\right)=\frac{t^{2}}{12}\left({nV_{e}}+{1+nV_{s}}\right)\,.$
+On the other hand, in accordance with the circuits proposed in the paper, we may assume that the ciphertext to be relinearized results from the product of two terms that have already undergone modulus switching.
+Consequently, the variance of the first component, namely $\nu_{\mathsf{mul}}$, is at least $nV_{\mathsf{ms}}^{2}F_{s}(1,1)$ where $V_{\mathsf{ms}}=\frac{t^{2}nV_{s}}{12}$, thereby making evident the negligible contribution of the key switching step.